### v1

# An unoriented skein exact triangle in unoriented link Floer homology

Abstract.
We define band maps in unoriented link Floer homology and show that they form an unoriented skein exact triangle. These band maps are similar to the band maps in equivariant Khovanov homology given by the Lee deformation.
As a key tool, we use a Heegaard Floer analogue of Bhat’s recent $2$-surgery exact triangle in instanton Floer homology, which may be of independent interest. Unoriented knot Floer homology corresponds to $I^{\sharp}$ of the knot in our $2$-surgery exact triangle.
1. Introduction
Heegaard Floer homology, defined by Ozsváth and Szabó [OS04c], is a powerful Floer theoretic invariant of closed, oriented three-manifolds. A closely related invariant for knots is knot Floer homology, defined by Ozsváth and Szabó [OS04a] and independently by Rasmussen [Ras03], which generalizes to oriented links. Link Floer homology categorifies the Alexander polynomial, and satisfies an oriented skein exact triangle [OS04a, Theorem 10.2] that categorifies the oriented skein relation. For unoriented links, there is a closely related invariant, unoriented link Floer homology, defined by Ozsváth, Stipsicz, and Szabó [OSS17b]. In this paper, we show that an algebraic variant of unoriented link Floer homology [^1] See Remark 1.9 for a comparison with the invariant defined in [OSS17b]., $\boldsymbol{HFL}^{\prime}{}^{-}(Y,L)$, satisfies an unoriented skein exact triangle over the field $\mathbb{F}=\mathbb{Z}/2$ with two elements.
Definition 1.1.
Three links $L_{a},L_{b},L_{c}\subset Y$ form an unoriented skein triple if the links are identical outside of a ball $B^{3}\subset Y$, in which they differ as in Figure 1.1.
Theorem 1.2.
Let $L_{a},L_{b},L_{c}\subset Y$ be an unoriented skein triple. Then there exists an $\mathbb{F}\llbracket U\rrbracket$-linear exact triangle
$\cdots\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{b})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{c})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\cdots.$
A key tool in proving Theorem 1.2 is Claim 5.3, a local version of Theorem 1.3, which may be of independent interest. Theorem 1.3 is a Heegaard Floer analogue of Bhat’s recent $2$-surgery exact triangle [Bha23] for $I^{\sharp}$, an instanton Floer theoretic invariant [Flo88] for three-manifolds and links defined by Kronheimer and Mrowka [KM11].
Theorem 1.3.
Let $K$ be a knot in $Y$. Then given any framing $\lambda$ of $K$, there exists an $\mathbb{F}\llbracket U\rrbracket$-linear exact triangle
$\cdots\to\boldsymbol{HFL^{\prime}}^{-}(Y,K)\to\boldsymbol{HF}^{-}(Y_{\lambda}(K))\to\boldsymbol{HF}^{-}(Y_{\lambda+2\mu}(K))\to\boldsymbol{HFL^{\prime}}^{-}(Y,K)\to\cdots,$
where $\mu$ is the meridian, $Y_{\lambda}(K)$, resp., $Y_{\lambda+2\mu}(K)$ is the $\lambda$, resp., $\lambda+2\mu$-surgery of $Y$ along $K$, and $\boldsymbol{HF}^{-}$ denotes the Heegaard Floer homology of the three-manifold.
These Heegaard Floer theoretic invariants come in various flavors, for which our exact triangles also hold (see Theorems 4.2 and 4.5). We propose the unreduced hat version, $\widehat{HFL^{\prime}}$ (Definition 1.6) as the counterpart of $I^{\sharp}$ in Heegaard Floer homology. Indeed, Theorem 1.2 is partly motivated by Kronheimer and Mrowka’s skein exact triangle for $I^{\sharp}$ [KM11]. Another closely related motivation is to define a spectral sequence from reduced Khovanov homology to knot Floer homology. See Subsection 1.4 for further discussion.
Figure 1.1. A local diagram for an unoriented skein triple
Remark 1.4.
We thank Fan Ye for communicating to the author that a similar analogue of $I^{\sharp}$ for knots in Heegaard Floer homology [^2] For a knot $K$, put two link basepoints on it and consider the chain complex over $\mathbb{Z}[H]/H^{2}$ where both basepoints have weight $2H$. has been suggested in his Miami talk [Ye23].
1.1. Unoriented link Floer homology
Let us define the unoriented link Floer homology groups that we study in this paper.
Definition 1.5.
Let $L$ be a $k$-component link in a closed, oriented three-manifold $Y$. To define the unoriented link Floer homology of $L$, we need to choose two basepoints $w_{i},z_{i}\in L_{i}$ for each component $L_{i}$ of $L$. Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{w},\boldsymbol{z})$ be an admissible, $2k$-pointed Heegaard diagram that represents this pointed link $L\subset Y$. Then the unoriented link Floer homology chain complex is freely generated by intersection points ${\bf x}\in\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}}$:
$\boldsymbol{CFL^{\prime}}^{-}(Y,L)=\bigoplus_{{\bf x}\in\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}}}\mathbb{F}\llbracket U_{1}^{1/2},\cdots,U_{k}^{1/2}\rrbracket{\bf x},$
and the differential is given by counting both basepoints $w_{i}$ and $z_{i}$ with weight $U_{i}^{1/2}$:
$\partial^{-}{\bf x}=\sum_{{\bf y}\in\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}}}\sum_{\phi\in D({\bf x},{\bf y}),\ \mu(\phi)=1}\#\mathcal{M}(\phi)\prod_{i}U_{i}^{\frac{1}{2}\left(n_{w_{i}}(\phi)+n_{z_{i}}(\phi)\right)}{\bf y}.$
The unoriented link Floer homology of the link $L\subset Y$ is $\boldsymbol{HFL^{\prime}}^{-}(Y,L)$, the homology of $\boldsymbol{CFL^{\prime}}^{-}(Y,L)$.
The action of the $U_{i}$’s on $\boldsymbol{CFL^{\prime}}^{-}(Y,L)$ are homotopic [^3] The same idea as [Zem15, Lemma 5.1] works., and so we also view $\boldsymbol{HFL^{\prime}}^{-}(Y,L)$ as an $\mathbb{F}\llbracket U\rrbracket$-module, where $U$ acts by multiplying by $U_{i}$.
We define the unreduced and reduced hat versions of unoriented link Floer homology. We will be lax about issues related to naturality in this section (in particular, basepoints) and postpone the discussion to Section 3.
Definition 1.6.
The (unreduced) hat version of unoriented link Floer homology, $\widehat{HFL^{\prime}}(Y,L)$, is the homology of the chain complex
$\widehat{CFL^{\prime}}(Y,L)=\boldsymbol{CFL^{\prime}}^{-}(Y,L)/U_{1}.$
Definition 1.7.
A marked link $L$ is a link together with a distinguished basepoint (which we call the marked point). The reduced (hat version of) unoriented link Floer homology of a marked link $L$, $\widetilde{HFL^{\prime}}(Y,L)$, is the homology of
$\widetilde{CFL^{\prime}}(Y,L)=\boldsymbol{CFL^{\prime}}^{-}(Y,L)/U_{i}^{1/2},$
where $L_{i}$ is the link component that the marked point is on.
Given a basepoint on a link $L$, the corresponding basepoint action is multiplication by $U_{i}^{1/2}$ if the point is on the link component $L_{i}$.
For knots, the reduced hat version, $\widetilde{HFK^{\prime}}$, is just knot Floer homology, $\widehat{HFK}$. We will study some examples shortly, in Subsubsections 1.2.1 and 1.2.2.
We end this subsection with some remarks that explain various aspects of the definitions.
Remark 1.8.
In the definition of the unreduced hat version, we could have quotiented out by any $U_{i}$, since multiplication by $U_{1}$ is homotopic to multiplication by $U_{i}$.
Remark 1.9.
All these unoriented link Floer chain complexes can be defined using a more general chain complex ${\mathcal{}CFL}^{-}$ defined over ${\mathcal{}R}=\mathbb{F}\llbracket Z_{1},\cdots,Z_{k},W_{1},\cdots,W_{k}\rrbracket$. For instance, we have
$\boldsymbol{CFL^{\prime}}^{-}={\mathcal{}CFL}^{-}\otimes_{{\mathcal{}R}}\mathbb{F}\llbracket U_{1}^{1/2},\cdots,U_{k}^{1/2},Z_{1},\cdots,Z_{k},W_{1},\cdots,W_{k}\rrbracket/(U_{i}^{1/2}=Z_{i}=W_{i}\ {\rm for\ all\ }i).$
In [OSS17b], they identify all the $U_{i}^{1/2}$’s together and work with polynomial rings instead of power series rings: the completion of their chain complex is
${\mathcal{}CFL}^{-}\otimes_{{\mathcal{}R}}\mathbb{F}\llbracket U^{1/2},Z_{1},\cdots,Z_{k},W_{1},\cdots,W_{k}\rrbracket/(U^{1/2}=Z_{i}=W_{i}\ {\rm for\ all\ }i).$
This is different from ours for links with more than one component, although they are the same for knots.
Remark 1.10.
For simplicity, let us focus on links in $S^{3}$. Let $L\subset S^{3}$ be a $k$-component oriented link. Link Floer homology $\widehat{HFL}(Y,L)$ has the Maslov gradings ${\rm gr}_{\boldsymbol{w}},{\rm gr}_{\boldsymbol{z}}\in\mathbb{Z}$ [^4] There are other conventions that differ from ours by an overall shift $\in\frac{1}{2}\mathbb{Z}$. and the Alexander grading [OS08a], which takes values in $\mathbb{H}=h+\mathbb{Z}^{k}\subset(\frac{1}{2}\mathbb{Z})^{k}$ for some $h$ [^5] If $L$ is a knot, then $\mathbb{H}=\mathbb{Z}$.. The collapsed Alexander grading, obtained by summing the coordinates of $\mathbb{H}$, is $({\rm gr}_{\boldsymbol{w}}-{\rm gr}_{\boldsymbol{z}})/2\in\mathbb{Z}$.
Hence, we can define $\chi_{{\rm gr}_{\boldsymbol{z}}}(\widehat{HFL}(S^{3},L))$, the Euler characteristic of $\widehat{HFL}(S^{3},L)$ with respect to ${\rm gr}_{\boldsymbol{z}}$, as an element of the free, rank $1$ module $\mathbb{Z}[\mathbb{H}]$ over the group ring $\mathbb{Z}[\mathbb{Z}^{k}]$. This Euler characteristic can be written in terms of the multi-variable Alexander polynomial.
Our unoriented link Floer homology group is singly graded, by the $\delta$-grading $({\rm gr}_{\boldsymbol{w}}+{\rm gr}_{\boldsymbol{z}})/2\in\mathbb{Z}$ (as in [OSS17b, Subsection 2.1]), since we identify $Z_{i}=W_{i}$ in the sense of Remark 1.9. However, the Alexander grading modulo $(2\mathbb{Z})^{k}$ still exists. Let $\mathbb{H}^{un}:=h+(\mathbb{Z}/2)^{\oplus k}\subset\left(\frac{1}{2}\mathbb{Z}/2\mathbb{Z}\right)^{\oplus k}$ be the corresponding $(\mathbb{Z}/2)^{\oplus k}$-torsor.
Hence, we can consider the Euler characteristic [^6] If we want to consider $L$ as an unoriented link, then the spaces $\mathbb{H}$ and $\mathbb{H}^{un}$ should be thought of as $H_{1}(Y\backslash L;\mathbb{Z})$- and $H_{1}(Y\backslash L;\mathbb{Z}/2)$-torsors living in $H_{1}(Y\backslash L;\frac{1}{2}\mathbb{Z})$ and $H_{1}(Y\backslash L;\frac{1}{2}\mathbb{Z}/2\mathbb{Z})$, respectively. Also, changing the orientation of a link component changes the $\delta$-grading by an overall shift, and hence changes the Euler characteristic by a factor of ±$\pm 1$. (Compare [OSS17b, Proposition 7.1].) Viewed this way, these Euler characteristics, up to sign, do not depend on the orientation of $L$. of the unreduced and reduced hat versions (let us mark the $j$th component of $L$) with respect to the $\delta$-grading as an element of the free, rank $1$ module $\mathbb{Z}[\mathbb{H}^{un}]$ over the group ring $\mathbb{Z}[(\mathbb{Z}/2)^{\oplus k}]$. Then, we have
$\chi_{\delta}(\widehat{HFL^{\prime}}(S^{3},L))=(1-m_{j})\chi_{\delta}(\widetilde{HFL^{\prime}}(S^{3},L))=\prod_{i=1}^{k}(1-m_{i})\Phi(\chi_{{\rm gr}_{\boldsymbol{z}}}(\widehat{HFL}(S^{3},L))),$
where $m_{i}$ is the $i$th unit vector of $\mathbb{Z}^{k}$, $(\mathbb{Z}/2)^{\oplus k}$; and $\Phi:\mathbb{Z}[\mathbb{H}]\to\mathbb{Z}[\mathbb{H}^{un}]$ is given by $\prod_{i}m_{i}^{k_{i}}\mapsto(-1)^{\sum_{i}k_{i}}\prod m_{i}^{k_{i}}.$ (We thank Jacob Rasmussen for this remark.)
Remark 1.11.
We assigned the weights $U_{i}^{1/2}$ to the basepoints and called multiplication by $U_{i}^{1/2}$ the basepoint action, to match the convention of equivariant Khovanov homology [Kho00, Lee05, LS22]: the homology groups have a $U$-action, and the basepoint action squared is the $U$-action. Similarly, the unreduced and reduced hat versions are defined analogously to Khovanov homology, which we now recall.
The equivariant Khovanov homology given by the Lee deformation uses the Frobenius algebra ${\mathcal{}A}=\mathbb{F}[x,U]/(x^{2}=U)$ over $\mathbb{F}[U]$ and the multiplication and comultiplication maps are $\mathbb{F}[U]$-linear maps given by the following.
[…]
$\Delta:{\mathcal{}A}\to{\mathcal{}A}\otimes_{\mathbb{F}[U]}{\mathcal{}A}:1\mapsto 1\otimes x+x\otimes 1,\ x\mapsto x\otimes x+U1\otimes 1$
Given a link diagram, one considers the cube of resolutions, and uses this Frobenius algebra to build a chain complex. In particular, the vertices of the cube of resolutions chain complex are the equivariant Khovanov homology of unlinks, which is
$\mathbb{F}[x_{1},\cdots,x_{k},U]/(x_{1}^{2}=\cdots=x_{k}^{2}=U)$
for the $k$-component unlink. The homology of this chain complex is equivariant Khovanov homology, which corresponds to the minus version [^7] Since we work over a field of characteristic $2$, the homology is not as interesting.. To get unreduced Khovanov homology, or the hat version, we quotient out the chain complex by $U$. If the link has a marked point, then the chain complex can be viewed as a chain complex over $\mathbb{F}[x,U]/(x^{2}=U)$, where the action of $x$ is given by the marked point in each resolution. The reduced Khovanov homology, or the reduced hat version, is given by quotienting out the chain complex by $x$.
Remark 1.12.
Unlike $I^{\sharp}$ and $I^{\natural}$, we need two link basepoints on each link component to define unoriented link Floer homology. In link Floer homology, different choices of basepoints give rise to isomorphic homology groups, but the isomorphism is in general noncanonical and interesting [Sar15]. However, moving the two unmarked link basepoints of a link component around the link component induces the identity on unoriented link Floer homology, although it is in general not null-homotopic. We define the hat versions more precisely in Definition 3.4 and show this in Proposition 3.10.
Remark 1.13.
We can define the minus version of unoriented link Floer homology over a polynomial ring instead of a power series ring; this involves assuming strong $\mathfrak{s}$-admissibility, which depends on the ${\rm Spin}^{c}$-structure $\mathfrak{s}$. The unoriented skein exact triangle, Theorem 1.16, holds in this version as well (Remark 4.4), although the $2$-surgery exact triangle, Theorem 1.3, does not (Remark 4.6). We mainly work with power series rings and suppress ${\rm Spin}^{c}$-structures to simplify the exposition.
1.2. An unoriented skein exact triangle
We will continue to be lax about basepoints in this subsection.
Definition 1.14.
Three (marked) links $L_{a},L_{b},L_{c}\subset Y$ form an unoriented skein triple if the links and the marked points (if the links are marked) are identical outside of a ball $B^{3}\subset Y$, in which they differ as in Figure 1.1.
Definition 1.15.
A band on a link $L\subset Y$ is an ambient $2$-dimensional $1$-handle, i.e. it is an embedding $\iota:[0,1]\times[0,1]\hookrightarrow Y$ such that $\iota^{-1}(L)=[0,1]\times\{0,1\}$. A link $L^{\prime}$ is obtained by a band move on a link $L$ if it is given by surgering $L$ along a band on $L$.
The following are the three types of bands:
  • a band is a non-orientable band if it does not change the number of connected components;
  • a band is a split band if it increases the number of connected components;
  • a band is a merge band if it decreases the number of connected components.
Unoriented skein triples are related by band moves, as in Figure 1.2, and the cyclic order of the three bands is always non-orientable, split, and merge.
Figure 1.2. A local diagram for an unoriented skein triple together with the band maps
Theorem 1.16.
Let $L_{a},L_{b},L_{c}\subset Y$ be an unoriented skein triple. Then there exist exact triangles in the various versions
$\cdots\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{b})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{c})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\cdots$
$\cdots\to\widehat{HFL^{\prime}}(Y,L_{a})\to\widehat{HFL^{\prime}}(Y,L_{b})\to\widehat{HFL^{\prime}}(Y,L_{c})\to\widehat{HFL^{\prime}}(Y,L_{a})\to\cdots$
$\cdots\to\widetilde{HFL^{\prime}}(Y,L_{a})\to\widetilde{HFL^{\prime}}(Y,L_{b})\to\widetilde{HFL^{\prime}}(Y,L_{c})\to\widetilde{HFL^{\prime}}(Y,L_{a})\to\cdots,$
where the maps can be interpreted as band maps. The maps are $\mathbb{F}\llbracket U\rrbracket$-linear, and also are equivariant with respect to suitable basepoint actions.
If $Y=S^{3}$, then the above band maps are homogeneous with respect to the $\delta$-grading [^8] In general, they preserve relative homological gradings; see Subsubsection 3.3.4..
We show that band maps for planar links coincide with Khovanov homology in Theorem 9.3.
1.2.1. Unlinks and the Hopf link
Recall that link Floer homology $\widehat{HFL}$ cannot have a skein exact triangle without any modifications, as there should be an exact triangle involving the unknot, the Hopf link, and the unknot [^9] In Manolescu’s unoriented skein exact triangle [Man07], the unknots have two extra basepoints each; this doubles the rank of $\widehat{HFL}({\rm unknot})$., but $\widehat{HFL}({\rm unknot})$ has rank $1$ and $\widehat{HFL}({\rm Hopf\ link})$ has rank $4$.
In contrast, the unreduced hat version of unoriented link Floer homology of the unlink $UL_{k}$ with $k$ components has rank $2^{k}$, and that of the Hopf link $H$ has rank $4$. The Hopf link and two unknots indeed form an exact triangle, and in fact all these homology groups are the same as Khovanov homology, as vector spaces.
1.2.2. Trefoil knots
The unoriented knot Floer homology of a trefoil in $S^{3}$ is
$\mathbb{F}\llbracket U^{1/2}\rrbracket\oplus\mathbb{F}\llbracket U^{1/2}\rrbracket/(U^{1/2}),$
and so the unreduced hat version has rank $4$ and the reduced hat version has rank $3$. The rank of the reduced version is the same as the ranks of both reduced Khovanov and reduced instanton Floer homology over any field, but the rank of the unreduced version is the same as the ranks of unreduced Khovanov and unreduced instanton Floer homology over $\mathbb{Q}$, but not over $\mathbb{F}=\mathbb{Z}/2$. This is as expected: similarly, it is conjectured that the $\widehat{HF}$ and $I^{\sharp}$ of a three-manifold are isomorphic over $\mathbb{Q}$ ([KM10, Conjecture 7.24], [LY22, Conjecture 1.1]), but they are not isomorphic over $\mathbb{F}$ ([Bha23, Theorem 1.5], [LY24]).
1.2.3. Comparing with Manolescu’s maps [Man07]
Manolescu’s unoriented skein exact triangle and ours both involve pointed links and maps between them. Let us discuss them for one specific example: a non-orientable band $B:K_{a}\to K_{b}$ between two unknots in $S^{3}$. In this case, the number of basepoints on the knots and the rank of the maps are different: Manolescu’s map has half rank, and our map is zero (as in Khovanov homology).
Recall that if we have an admissible Heegaard diagram with three attaching curves $\boldsymbol{\alpha},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b}$, then we can define a holomorphic triangle counting map
$\mu_{2}:CF(\boldsymbol{\alpha},\boldsymbol{\beta}_{a})\otimes CF(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})\to CF(\boldsymbol{\alpha},\boldsymbol{\beta}_{b}).$
The band maps are $\mu_{2}(-\otimes\Theta_{B})$ for some cycle $\Theta_{B}\in CF(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})$.
Manolescu’s unoriented skein exact triangle requires the two unknots to each have at least four basepoints. See Figure 1.3101010The crossing of $K_{b}$ is drawn correctly; see Remarks 2.12 and 2.43 for orientation conventions.: we consider the simplest case where there are exactly four. First, the attaching curves $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}_{i}$ describe $K_{i}$ for $i\in\{a,b\}$, i.e. $\widehat{HFK}(S^{3},K_{i})=\widehat{HF}(\boldsymbol{\alpha},\boldsymbol{\beta}_{i})$ [^11] By $\widehat{HF}(\boldsymbol{\alpha},\boldsymbol{\beta}_{i})$, we mean the homology of the chain complex freely generated over $\mathbb{F}$ by intersection points $\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}_{i}}$ and where the differential counts holomorphic bigons that do not cross any basepoints.. Under this idenfication, the cycle $\Theta_{B}=\tau_{0}+\tau_{1}\in\widehat{CF}(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})$ induces the band map $\widehat{HFK}(S^{3},K_{a})\to\widehat{HFK}(S^{3},K_{b})$. It turns out that all the maps in Manolescu’s exact triangle come from similar looking diagrams.
In contrast, our exact triangle requires the two knots to have exactly two basepoints, and they need to be on the same side of the knot with respect to the band, as in Figure 1.4 (we could have chosen $w_{1},z_{1}$ instead). See Figure 1.4: similarly, $\boldsymbol{HFK^{\prime}}^{-}(S^{3},K_{i})=\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{i})$ [^12] By $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{i})$, we mean the homology of the chain complex freely generated over $\mathbb{F}\llbracket U^{1/2}\rrbracket$ by intersection points $\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}_{i}}$ and where the differential counts every holomorphic bigon, where both basepoints have weight $U^{1/2}$, as in Definition 1.5., and under this idenfication, the cycle $\tau_{1}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})$ induces the band map $\boldsymbol{HFK^{\prime}}^{-}(S^{3},K_{a})\to\boldsymbol{HFK^{\prime}}^{-}(S^{3},K_{b})$ (and also in the unreduced and reduced hat versions). It turns out that all the non-orientable band maps in our exact triangle come from similar looking diagrams; however, the diagrams for the split and merge band maps are different, and these band maps are more complicated to describe.
Figure 1.3. A non-orientable band between two unknots for Manolescu’s exact triangle, and a Heegaard diagram for it
Figure 1.4. A non-orientable band between two unknots for our exact triangle, and a Heegaard diagram for it
1.3. A word on the proof
We discuss some key ideas involved in the definition of the split and merge band maps and the proof of Theorem 1.16.
The split and merge band maps involve links with different numbers of link components, and it is not obvious how to define them since we want exactly two basepoints on each link component. (Compare [OS09].) We achieve this by adding a basepoint in the three-manifold when we consider the links with one fewer component (this does not change the quasi-isomorphism type of the chain complex [OS08a, Proposition 6.5]). This basepoint becomes two link basepoints on the link with more components.
See Figure 1.5: if we consider only $\boldsymbol{\beta}_{a}$ or $\boldsymbol{\beta}_{b}$, then we treat the basepoints $w_{1}$ and $z_{1}$ as one basepoint (with weight $U_{1})$ on the three-manifold. When we consider $\boldsymbol{\beta}_{c}$, this basepoint becomes two basepoints (with weight $U_{1}^{1/2}$ each) on the extra link component.
As usual, we prove the exact triangle by essentially using the triangle detection lemma [OS05, Lemma 4.2] and reducing it to a local computation. We have to work in ${\rm Sym}^{2}(\mathbb{T}^{2})$, i.e. we have to compute the number of certain kinds of holomorphic disks, triangles, and quadrilaterals in ${\rm Sym}^{2}(\mathbb{T}^{2})$, which makes the local computation challenging. Instead of directly carrying out the local computation in ${\rm Sym}^{2}(\mathbb{T}^{2})$, we use a trick that reduces it to two local computations in $\mathbb{T}^{2}$, which are combinatorial and well understood. This trick utilizes Claim 5.3, a local version of Theorem 1.3, which is a Heegaard Floer analogue of Bhat’s $2$-surgery exact triangle [Bha23]. (See the discussion right before Theorem 4.5 and Section 5.)
Figure 1.5. We reduce the proof of Theorem 1.16 to a local computation, Theorem 4.3.
Remark 1.17.
We prove Theorem 1.3 directly, by doing a model computation on the torus in the spirit of [OS04b, OS05]. We thank Ian Zemke for communicating to the author that Theorem 1.3 can also be proven using his bordered $\boldsymbol{HF}^{-}$ theory [Zem21a, MO10, OS08b].
Remark 1.18.
The instanton invariants $I^{\sharp}(Y,L)$ come from the orbifold whose underlying three-manifold is $Y$ and $L$ is the $\mathbb{Z}/2$-orbifold points, and the maps in Bhat’s $2$-surgery exact triangle [Bha23] come from orbifold cobordism maps.
We show that $\boldsymbol{CFL^{\prime}}^{-}(Y,L)$ splits into ${\rm Spin}^{c}(Y(L))$-summands where ${\rm Spin}^{c}(Y(L))$ is an $H_{1}^{orb}(Y(L))$-torsor, but we do not discuss further connections with orbifolds in this paper: we do not interpret the maps in Theorem 1.3 as orbifold cobordism maps. However, our maps come from triple Heegaard diagrams, and so it is possible to interpret them as the composition of a link cobordism map (in the sense of [Zem19]) together with a purely algebraic map.
1.4. Motivation
Link Floer homology and instanton Floer homology are closely related to Khovanov homology [Kho00], which is defined combinatorially from an unoriented cube of resolutions. Rasmussen [Ras05] conjectured that the rank of reduced Khovanov homology is always greater than or equal to the rank of knot Floer homology; Dowlin [Dow24] proved this conjecture over $\mathbb{Q}$.
One challenge in relating Khovanov homology to knot Floer homology lies in the difficulty of finding a suitable generalization of knot Floer homology for links that satisfies an unoriented skein exact triangle [^13] Dowlin’s spectral sequence uses an oriented cube of resolutions which involves singular knots.. Manolescu’s unoriented skein exact triangle [Man07] in link Floer homology requires additional basepoints on the links. As a result, the $E^{2}$ page of the corresponding cube of resolutions is not a link invariant, which was shown by Baldwin and Levine [BL12]. Also, Baldwin, Levine, and Sarkar [BLS17] suggested that one should take the Koszul resolution with respect to the basepoint actions on the Khovanov side.
In instanton Floer homology, Kronheimer and Mrowka [KM11] defined a spectral sequence from unreduced Khovanov homology to $I^{\sharp}$ by defining and iterating an unoriented skein exact triangle for $I^{\sharp}$. They also defined a spectral sequence from reduced Khovanov homology to $I^{\natural}$, a reduced version of $I^{\sharp}$, which is conjecturally isomorphic to $\widehat{HFK}$ for knots over $\mathbb{Q}$ ([KM10, Conjecture 7.24], [LY22, Conjecture 1.1]). Hence, a Heegaard Floer analogue of $I^{\sharp}$ and $I^{\natural}$ for (marked) links would be a good candidate for a suitable generalization of knot Floer homology from the perspective of Khovanov homology.
Our motivation is to develop a theory in Heegaard Floer homology that is analogous to $I^{\sharp}$ and $I^{\natural}$, which we hope will include a spectral sequence from Khovanov homology. We propose the unreduced and reduced hat versions of unoriented link Floer homology (Definitions 1.6 and 1.7) as the counterparts of $I^{\sharp}$ and $I^{\natural}$, and show Theorems 1.3 and 1.16 which are counterparts of theorems in instanton Floer homology.
It is interesting to compare our proposal with the existing conjectures that relate Heegaard Floer homology and $I^{\#},I^{\natural}$. First, for knots, the counterpart of $I^{\natural}$ that we propose is indeed $\widehat{HFK}$. Also, for a three-manifold $Y$, $I^{\sharp}(Y)$ is conjecturally isomorphic to $\widehat{HF}(Y)$ over $\mathbb{Q}$, and a version of our $2$-surgery exact triangle involves $\widehat{HF}(Y_{\lambda}(K))$, $\widehat{HF}(Y_{\lambda+2\mu}(K))$, and $\widehat{HFK^{\prime}}(Y,K)$ (over $\mathbb{F}$), whereas Bhat’s $2$-surgery exact triangle [Bha23] involves $I^{\sharp}(Y_{\lambda}(K))$, $I^{\sharp}(Y_{\lambda+2\mu}(K))$, and $I^{\sharp}(Y,K)$.
In Theorem 9.3, we show that for planar links, the band maps in the various versions of unoriented link Floer homology agree with the band maps in (equivariant) Khovanov homology, and hence also with the band maps in $I^{\sharp}$ and $I^{\natural}$. Hence, if one could iterate our unoriented skein exact triangle, then one would get a spectral sequence from equivariant Khovanov homology to the minus version of unoriented link Floer homology, and also a spectral sequence from reduced Khovanov homology to knot Floer homology.
1.5. Future directions
It is natural to ask whether one can define unoriented link cobordism maps in unoriented link Floer homology, just like $I^{\sharp}$. Although we do not discuss this in this paper, it turns out that showing that band maps commute is a key step for iterating our unoriented skein exact triangle and getting spectral sequences from Khovanov homology to unoriented link Floer homology. We hope to achieve this in a later work.
It would also be interesting to be able to compute these band maps (or more ambitiously, the conjectured spectral sequence), either using ideas from bordered Heegaard Floer homology [LOT18] or grid homology [OSS15]. One difficulty in interpreting our maps in grid homology comes from that we only consider links with exactly two basepoints on each component. We do not know whether there are similar band maps in the case where there are more basepoints, such that they form an unoriented skein exact triangle.
Also, we do not deal with signs in this paper and work over a field of characteristic $2$. In contrast, Dowlin [Dow24] has to divide by $2$. It would be interesting to see whether the theorems in this paper hold over $\mathbb{Z}$.
There are two different $2$-surgery exact triangles in Heegaard Floer homology: there is Theorem 1.3, and also Ozsváth and Szabó’s integer surgery exact triangle [OS04b, Section 9.5], which has two copies of $\boldsymbol{HF}^{-}(Y)$ instead of $\boldsymbol{HFL^{\prime}}^{-}(Y,K)$ (and the maps are different). In [Nah], we show analogous statements for any positive rational surgery slope: we get (at least) $p$ different $p/q$-surgery exact triangles, one of which generalizes Ozsváth and Szabó’s $n$ or $1/n$-surgery exact triangle. The others involve algebraic modifications of knot Floer homology; more precisely, these are variants (if necessary) of $2i/p$-modified knot Floer homology [OSS17a] for $i=0,1,\cdots,p-1$. Note that $1$-modified knot Floer homology is unoriented knot Floer homology.
1.6. Organization
In Section 2, we start by setting notations and recalling some definitions in Heegaard Floer theory in Subsection 2.1. In order to prove Theorems 1.3 and 1.16, we consider local systems that a priori involve negative powers of $U$. We define the type of local systems we consider in Subsection 2.2, and discuss weak admissibility and positivity in Subsection 2.3, which ensure that the differential and (higher) composition maps are well-defined. We use the language of $A_{\infty}$-categories to simplify the homological algebra; we recall this in Subsection 2.4. We set up notations for standard translates in Subsection 2.5 and recall theorems on stabilizations in Subsection 2.6. In Subsection 2.7, we clarify our conventions for link Floer homology, and we define ${\rm Spin}^{c}$-structures for unoriented links and strong admissibility in Subsection 2.8. We set up conventions for gradings in Subsection 2.9, especially when nontrivial local systems are present. We introduce a simplifying assumption in Subsection 2.10, in which case checking whether gradings exist is simpler.
In Section 3, we precisely define the objects we consider, define the band maps, and show that they are well-defined. Using these notions, we state the main theorem, the unoriented skein exact triangle (and the $2$-surgery exact triangle), in Section 4. In Section 5, we briefly discuss the key steps of the proof, and we carry out the local computations in Sections 6 and 7. Section 7 is also the local computation for the $2$-surgery exact triangle. We use these computations in Section 8 and finish off the proof.
Finally, we compute the band maps for planar links in Section 9.
1.7. Acknowledgements
We thank Peter Ozsváth for his continuous support, explaining a lot of the arguments in this paper, and helpful discussions. We also thank Ian Zemke for his continuous support, teaching the author a lot of previous works, especially [Zem23], and helpful discussions. We thank Deeparaj Bhat for sharing his work on the $2$-surgery exact triangle back in March 2023. We thank John Baldwin, Deeparaj Bhat, Fraser Binns, Evan Huang, Yonghwan Kim, Jae Hee Lee, Adam Levine, Jiakai Li, Robert Lipshitz, Marco Marengon, Sucharit Sarkar, Zoltán Szabó, Alison Tatsuoka, Jacob Rasmussen, Joshua Wang, and Fan Ye for helpful discussions. We also thank Robert Lipshitz, Peter Ozsváth, Jacob Rasmussen, Joshua Wang, and Ian Zemke for helpful comments on earlier drafts of this paper.
2. Preliminaries
2.1. Heegaard Floer homology and local systems
We set up notations and notions that we use throughout this paper. We use local systems that a priori involve negative powers of $U$, as in [Zem23]. Although it is not necesary to consider nontrivial local systems to define the unoriented link Floer homology groups and the band maps, we will use them to prove the exact triangles.
We do not need any new analytic input, and thus we will refrain from discussing the analytic foundations and refer the readers to [OS04c, OS04b, Lip06, LOT16, HHSZ20].
Definition 2.1 ([OS08a, Definition 3.1]).
Given a closed, oriented, genus $g$ surface $\Sigma$ together with a set of basepoints $\boldsymbol{p}=\{p_{i}\}_{i=1}^{l}$ on $\Sigma$, an attaching curve is a set $\boldsymbol{\alpha}$ of pairwise disjoint, simple closed curves on $\Sigma\backslash\boldsymbol{p}$ whose images span a $g$-dimensional subspace of $H_{1}(\Sigma)$. A multi-Heegaard diagram [^14] This definition will be slightly modified in Subsection 2.2. is a collection $(\Sigma,\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{m},\boldsymbol{p})$ of such a pointed surface together with attaching curves. We assume that attaching curves intersect transversely, and that there are no triple intersections.
Notation 2.2.
We write $\alpha$ for one alpha circle. The boldsymbol $\boldsymbol{\alpha}$ means a set of alpha circles, and write $\boldsymbol{\alpha}=\{\alpha^{1},\cdots,\alpha^{g}\}$. If the curves $\alpha^{i}$ lie in $\Sigma$, then, by abuse of notation, $\boldsymbol{\alpha}$ also means the image of $\mathbb{T}_{\boldsymbol{\alpha}}=\alpha^{1}\times\cdots\times\alpha^{g}$ in ${\rm{\rm Sym}^{g}(\Sigma)}$ and $\alpha^{1}\cup\cdots\cup\alpha^{g}\subset\Sigma$.
Definition 2.3.
Given a multi-Heegaard diagram $(\Sigma,\boldsymbol{\alpha}_{0},\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{m},\boldsymbol{p})$, an elementary two-chain is a connected component of $\Sigma\backslash\left(\boldsymbol{\alpha}_{0}\cup\cdots\cup\boldsymbol{\alpha}_{m}\right)$, and a two-chain is a formal sum ${\mathcal{}D}$ of elementary two-chains. A two-chain is nonnegative if its local multiplicities are nonnegative. Denote the $\boldsymbol{\alpha}_{i}$-boundary of ${\mathcal{}D}$ as $\partial_{\boldsymbol{\alpha}_{i}}{\mathcal{}D}$ (which is a one-chain). A cornerless two-chain is a two-chain ${\mathcal{}D}$ for which $\partial_{\boldsymbol{\alpha}_{i}}{\mathcal{}D}$ is a cycle for all $i$. A cornerless two-chain is periodic if it does not contain any basepoints.
Definition 2.4.
[…]
$\partial(\partial_{\boldsymbol{\alpha}_{i_{j}}}{\mathcal{}D})={\bf x}_{j+1}-{\bf x}_{j},\ \partial_{\boldsymbol{\alpha}_{i}}{\mathcal{}D}=0$
for $j=0,\cdots,k$ and $i\neq i_{0},\cdots,i_{k}$ [^16] Not all two-chains lift to a domain.. Let $D({\bf x}_{0},\cdots,{\bf x}_{k})$ be the set of domains with vertices ${\bf x}_{0},\cdots,{\bf x}_{k}$.
An $\boldsymbol{\alpha}_{i_{0}}\cdots\boldsymbol{\alpha}_{i_{k}}$-domain is a domain with vertices ${\bf x}_{0},\cdots,{\bf x}_{k}$ for some intersection points ${\bf x}_{j}\in\boldsymbol{\alpha}_{i_{j}}\cap\boldsymbol{\alpha}_{i_{j+1}}$, and a domain with $k+1$ vertices is an $\boldsymbol{\alpha}_{i_{0}}\cdots\boldsymbol{\alpha}_{i_{k}}$-domain for some $i_{0},\cdots,i_{k}$.
By abuse of notation, we also denote domains as ${\mathcal{}D}$. Also, by abuse of notion, we say a two-chain ${\mathcal{}D}$ is in $D({\bf x}_{0},\cdots,{\bf x}_{k})$ if there exists a domain in $D({\bf x}_{0},\cdots,{\bf x}_{k})$ whose two-chain is ${\mathcal{}D}$, and we may identify ${\mathcal{}D}$ with the corresponding domain.
Remark 2.5.
Note that the two-chain of a domain uniquely determines the domain if the domain has at least three vertices (some of which might be identical). This may not be the case otherwise: for instance, given any $\boldsymbol{\alpha},\boldsymbol{\beta}$-cornerless two-chain for some $\boldsymbol{\alpha},\boldsymbol{\beta}$, for any ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$, there exists a domain in $D({\bf x},{\bf x})$ whose underlying two-chain is the given two-chain [^17] A cornerless two-chain can be lifted to a domain if and only if it has exactly two vertices (which are identical).. (These are not the only examples.)
The Maslov index $\mu({\mathcal{}D})$ of a domain ${\mathcal{}D}$ can be computed (and/or defined) combinatorially, using the formulas in [Lip06, Section 4] and [Sar11]. Note that $\mu({\mathcal{}D})$, in general, cannot be defined just from the underlying two-chain of ${\mathcal{}D}$ [^18] Similarly, the monodromy (Definition 2.10) also depends on the vertices of ${\mathcal{}D}$..
It is shown in [Sar11, Theorems 3.2, 3.3] that $\mu({\mathcal{}D})$ is cyclically symmetric in the vertices of ${\mathcal{}D}$, and that $\mu$ is additive, i.e. $\mu({\mathcal{}D}_{1})+\mu({\mathcal{}D}_{2})=\mu({\mathcal{}D}_{1}+{\mathcal{}D}_{2})$, with respect to the following composition of domains: given ${\mathcal{}D}_{1}\in D({\bf x}_{a},\cdots,{\bf x}_{a+b-1},{\bf y})$ and ${\mathcal{}D}_{2}\in D({\bf x}_{0},\cdots,{\bf x}_{a-1},{\bf y},{\bf x}_{a+b},\cdots,{\bf x}_{k})$ ($a\geq 0,b\geq 1$), their composition is ${\mathcal{}D}_{1}+{\mathcal{}D}_{2}\in D({\bf x}_{0},\cdots,{\bf x}_{k})$.
To define the Heegaard Floer chain complex, we need to choose a coefficient ring and assign weights to basepoints. We also equip attaching curves with local systems.
Definition 2.6.
If $R$ is a power series ring $R=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$, denote
$R^{\infty}=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket[X_{1}^{-1},\cdots,X_{n}^{-1}].$
Note that this depends on the identification $R=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$. For an $R=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$-module $E$, write $E^{\infty}:=E\otimes_{R}R^{\infty}.$
A coefficient ring is a power series ring $R=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$, its quotient, or $R^{\infty}$.
Definition 2.7.
If $\boldsymbol{p}$ is the set of all basepoints and $R$ is a coefficient ring, then a weight function is a function $w:\boldsymbol{p}\to R$. Given a weight function $w$, define the weight of a two-chain ${\mathcal{}D}$ as
$w({\mathcal{}D}):=\prod_{p\in\boldsymbol{p}}w(p)^{n_{z}({\mathcal{}D})}.$
Definition 2.8.
A local system over $R$ on an attaching curve $\boldsymbol{\alpha}$ is a pair $(E,\Phi)$ where $E$ is a free $R$-module, and $\Phi$ is the monodromy: it is a groupoid representation
$\Phi:\Pi_{1}(\boldsymbol{\alpha})\to{\rm Hom}_{R}(E,E),$
where $\Pi_{1}(\boldsymbol{\alpha})$ is the fundamental groupoid of $\boldsymbol{\alpha}=\mathbb{T}_{\boldsymbol{\alpha}}$, and ${\rm Hom}_{R}(E,E)$ is a groupoid with one object whose automorphism group is ${\rm Hom}_{R}(E,E)$. Write $\boldsymbol{\alpha}^{(E,\Phi)}$, or simply $\boldsymbol{\alpha}^{E}$, to signify that $\boldsymbol{\alpha}$ is equipped with the local system $(E,\Phi)$.
The local systems in this paper will all be specified by an oriented arc $G$ on the Heegaard surface that satisfies certain conditions. We define this in Subsection 2.2 and also modify the definition of a Heegaard diagram to a collection $(\Sigma,\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{m},\boldsymbol{p})$ or $(\Sigma,\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{m},\boldsymbol{p},G)$.
Definition 2.9.
A Heegaard datum is a tuple that consists of a multi-Heegaard diagram $(\Sigma,\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{m},\boldsymbol{p})$, a coefficient ring $R$, a weight function $w:\boldsymbol{p}\to R$, and local systems $(E_{i},\Phi_{i})$ over $\boldsymbol{\alpha}_{i}$ for $i=0,\cdots,m$.
Definition 2.10.
Given a Heegaard datum with Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$, coefficient ring $R$, weight function $w:\boldsymbol{p}\to R$, and local systems $(E_{\boldsymbol{\alpha}},\Phi_{\boldsymbol{\alpha}}),(E_{\boldsymbol{\beta}},\Phi_{\boldsymbol{\beta}})$ over $\boldsymbol{\alpha},\boldsymbol{\beta}$, respectively, define the group $CF(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ as a direct sum of ${\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})$’s:
$CF_{R,w}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})=\bigoplus_{{\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}}{\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}}){\bf x},$
and define the differential $\partial$ to be an $R$-linear map such that
(2.1) $\partial(e{\bf x})=\sum_{{\mathcal{}D}\in D({\bf x},{\bf y}),\ \mu({\mathcal{}D})=1}\#{\mathcal{}M}({\mathcal{}D})w({\mathcal{}D})\rho({\mathcal{}D})(e){\bf y},$
where $e\in{\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})$ and $\rho({\mathcal{}D}):{\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})\to{\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})$ is the monodromy of ${\mathcal{}D}$:
$\rho({\mathcal{}D})=\Phi_{\boldsymbol{\beta}}(\partial_{\boldsymbol{\beta}}{\mathcal{}D})\circ e\circ\Phi_{\boldsymbol{\alpha}}(\partial_{\boldsymbol{\alpha}}{\mathcal{}D}).$
We refer to elements of $CF_{R,w}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ as morphisms from $\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}}$ to $\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}$. In most cases, we omit the subscripts $R$ and $w$. Also, we often decorate the symbol $CF$: for instance, $\boldsymbol{CF}^{-}$ means that we are working over a power series ring, and $\widehat{CF}$ means that we are working over some quotient of the power series ring. We will mostly focus on the $\boldsymbol{CF}^{-}$ case.
We get (higher) composition maps $\mu_{n}$ [^19] The differential $\partial=\mu_{1}$ is also a composition map. when we have many attaching curves. If $\boldsymbol{\alpha}_{0}^{(E_{0},\Phi_{0})},\cdots,\boldsymbol{\alpha}_{d}^{(E_{d},\Phi_{d})}$ are attaching curves with local systems, then the composition map is given by
(2.2) $\mu_{d}:CF(\boldsymbol{\alpha}_{0}^{E_{0}},\boldsymbol{\alpha}_{1}^{E_{1}})\otimes\cdots\otimes CF(\boldsymbol{\alpha}_{d-1}^{E_{d-1}},\boldsymbol{\alpha}_{d}^{E_{d}})\to CF(\boldsymbol{\alpha}_{0}^{E_{0}},\boldsymbol{\alpha}_{d}^{E_{d}}):\\ e_{1}{\bf x}_{1}\otimes\cdots\otimes e_{d}{\bf x}_{d}\mapsto\sum_{{\mathcal{}D}\in D({\bf x}_{1},\cdots,{\bf x}_{d},{\bf y}),\ \mu({\mathcal{}D})=2-d}\#{\mathcal{}M}({\mathcal{}D})w({\mathcal{}D})\rho({\mathcal{}D})(e_{1}\otimes\cdots\otimes e_{d}){\bf y},$
where $\rho({\mathcal{}D})$ is the monodromy defined as follows: given a collection of paths $p_{j}:{\bf x}_{j}\to{\bf x}_{j+1}$ in $\boldsymbol{\alpha}_{j}$ ($j=0,\cdots,d$) for intersection points ${\bf x}_{j}\in\boldsymbol{\alpha}_{j-1}\cap\boldsymbol{\alpha}_{j}$ ($j=0,\cdots,d$, $\boldsymbol{\alpha}_{-1}:=\boldsymbol{\alpha}_{d}$), define the monodromy of $p_{0},\cdots,p_{d}$ as
$\rho(p_{0},\cdots,p_{d}):{\rm Hom}_{R}(E_{0},E_{1})\otimes\cdots\otimes{\rm Hom}_{R}(E_{d-1},E_{d})\to{\rm Hom}_{R}(E_{0},E_{d}):\\ e_{1}\otimes\cdots\otimes e_{d}\mapsto\Phi_{d}(p_{d})\circ e_{d}\circ\Phi_{d-1}(p_{d-1})\circ e_{d-1}\circ\cdots\circ\Phi_{1}(p_{1})\circ e_{1}\circ\Phi_{0}(p_{0}).$
Given an $\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{d}$-domain ${\mathcal{}D}$, its monodromy is $\rho({\mathcal{}D})=\rho(\partial_{\boldsymbol{\alpha}_{0}}({\mathcal{}D}),\cdots,\partial_{\boldsymbol{\alpha}_{d}}({\mathcal{}D}))$.
Remark 2.11.
We need the Heegaard datum to satisfy an admissibility criterion (Subsubsection 2.3.1) for the sum in the definition of $\mu_{n}$ to be well-defined. Also, the groupoid representations we consider in this paper will a priori have negative powers of $U$: indeed, they will be maps into ${\rm Hom}_{R}(E_{\boldsymbol{\alpha}}^{\infty},E_{\boldsymbol{\beta}}^{\infty})\simeq{\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})\otimes_{R}R^{\infty}$. We show in Subsection 2.3.2 that this is not a problem for the Heegaard data that we consider in this paper (see Subsection 2.2).
We need to choose a family of almost complex structure (on ${\rm Sym}^{d}(\Sigma)$, or on $\Sigma\times D_{n}$, where $D_{n}$ is a disk with $n$ punctures, if we work in the cylindrical reformulation) to define the Heegaard Floer chain complex and the composition maps. In this paper, we will have to work in various different (families of) almost complex structures (for instance, see Section 5). Using the arguments of [Sei08, Section 10c] and [LOT16, Section 3.3], it is possible to show that a generic one-parameter family that interpolates two such (families of) almost complex structures induces an $A_{\infty}$-functor between the $A_{\infty}$-categories given by the two choices, and that this $A_{\infty}$-functor induces isomorphisms on the homology groups. (See Remark 2.25 for a definition of the $A_{\infty}$-categories we consider.)
Remark 2.12.
We use the convention that almost complex structures on surfaces rotate the tangent spaces in a counterclockwise direction. This is the same as [Man07] but is different from [BL12, BLS17].
2.2. Weights and local systems for unoriented link Floer homology
From now on, all the Heegaard data will be of the following form, unless otherwise specified. We recall the conventions for the link we get from such Heegaard diagrams in Subsection 2.7.
The local systems will be specified by an oriented arc $G\subset\Sigma$ on the Heegaard surface. Hence, instead of including local systems as a part of a Heegaard datum, we simply incorporate $G$ into the Heegaard diagram: modify the definition of a multi-Heegaard diagram to be a collection of a pointed surface, attaching curves, and possibly also an oriented arc $G\subset\Sigma$ (this will always be drawn dashed and grey). If it includes $G$, then the coefficient ring $R$ has a distinguished variable $U$; $R=\mathbb{F}\llbracket X_{1},X_{2},\cdots,X_{k}\rrbracket$ or $\mathbb{F}\llbracket U,X_{1},X_{2},\cdots,X_{k}\rrbracket$ (sometimes we consider a quotient of it or $R^{\infty}$).
The negative boundary $\partial_{-}G$ of $G$ is a basepoint. For any attaching curve $\boldsymbol{\alpha}$, there is at most one circle $\alpha^{i}\in\boldsymbol{\alpha}$ that intersects $G$, and the two endpoints of $G$ lie in the same connected component of $\Sigma\backslash\boldsymbol{\alpha}$. Furthermore, we assume that there are at most two circles that intersect $G$, and that they are standard, small translates [^20] These two circles intersect each other at two points, and intersect other circles in the “same way”: see Definition 2.32. of each other, that intersect each other at two points, right above and below $G$. See Figure 2.1 for a schematic near $G$.
If an attaching curve $\boldsymbol{\alpha}$ does not intersect $G$, then it has the trivial local system. If some $\alpha^{i}\in\boldsymbol{\alpha}$ intersects $G$, then we consider the local system $(E,\phi)$ where $E$ is the rank $2$ free $R$-module $E=e_{0}R\oplus e_{1}R$, and the monodromy sends a path $p$ in $\boldsymbol{\alpha}$ to $\phi^{\#(G\cap p)}$, where
$\phi=e_{1}e_{0}^{\ast}+Ue_{0}e_{1}^{\ast}\in{\rm Hom}_{R}(E^{\infty},E^{\infty}).$
The intersection $G\cap p$ is positive if $p$ goes from top to bottom in Figure 2.1.
Furthermore, $X_{i}=U_{i}$ or $U_{i}^{1/2}$ for each $i$. If $X_{i}=U_{i}^{1/2}$, then there are exactly two basepoints with weight $U_{i}^{1/2}$; we call these basepoints link basepoints. If $X_{i}=U_{i}$, then there is exactly one basepoint with weight $U_{i}$; we call these free basepoints [^21] The basepoint $\partial_{-}G$ is neither a link basepoint nor a free basepoint.. We also require that for each attaching curve $\boldsymbol{\alpha}$ and connected component ${\mathcal{}D}$ of $\Sigma\backslash\boldsymbol{\alpha}$, ${\mathcal{}D}$ has weight $U$ or $U_{i}$ for some $i$.
Figure 2.1. A local diagram for the Heegaard diagram near $G$
Definition 2.13.
We chose a basis for the nontrivial local system $E$. Choose the trivial basis $\{1\}$ for the trivial local system. These induce $R$-bases on the $R$-Hom spaces. An $R$-generator (or simply a generator) of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is of the form $f{\bf x}$, where $f$ is an $R$-basis element of ${\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})$ and ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$.
Definition 2.14.
If ${\mathcal{}D}$ is a two-chain, define the total multiplicity of ${\mathcal{}D}$ as
$P({\mathcal{}D}):=n_{link}({\mathcal{}D})+2n_{free}({\mathcal{}D}),$
where $link$ is the set of link basepoints together with $\partial_{+}G,\partial_{-}G$ (if $G$ exists), and $free$ is the set of all free basepoints.
Example 2.15.
Let us consider the first two genus $1$ Heegaard diagrams in Figure 2.2, where we work over $\mathbb{F}\llbracket U\rrbracket$, and $U$ is the distinguished variable.
Let us consider the first diagram. The attaching curve $\alpha$ has a nontrivial local system, but $\beta$’s local system is trivial. The chain complex $\boldsymbol{CF}^{-}(\alpha^{E_{\alpha}},\beta)$ is freely generated by $e_{0}a,e_{1}a,e_{0}b,e_{1}b,e_{0}c,e_{1}c$ as an $\mathbb{F}\llbracket U\rrbracket$-module. The differential acts on the generators by
$e_{0}a\mapsto e_{0}b,\ e_{1}a\mapsto e_{1}b,\ e_{0}a\mapsto e_{1}c,\ e_{1}a\mapsto Ue_{0}c,$
where the first two are given by the two-chain $D_{1}$ and the last two are given by $D_{2}$. Note that if we identify $e_{0}x$ with $x$ and $e_{1}x$ with $U^{1/2}x$ for $x=a,b,c$, then this chain complex is isomorphic to the chain complex $\boldsymbol{CF}^{-}(\alpha,\beta)$ given by the third diagram, where we work over $\mathbb{F}\llbracket U^{1/2}\rrbracket$, and where we assign the weight $U^{1/2}$ to both $z$ and $w$.
Let us consider the second diagram. Both attaching curves $\alpha$ and $\beta$ have a nontrivial local system. The module $\boldsymbol{CF}^{-}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$ is freely generated by $e_{i}e_{j}^{\ast}\theta^{+},e_{i}e_{j}^{\ast}\theta^{-}$ ($i,j=0,1$) as an $\mathbb{F}\llbracket U\rrbracket$-module. The differential is as follows:
$\begin{gathered}e_{0}e_{0}^{\ast}\theta^{+}\mapsto e_{1}e_{1}^{\ast}\theta^{-}+e_{0}e_{0}^{\ast}\theta^{-},\ e_{1}e_{1}^{\ast}\theta^{+}\mapsto e_{0}e_{0}^{\ast}\theta^{-}+e_{1}e_{1}^{\ast}\theta^{-},\\ e_{1}e_{0}^{\ast}\theta^{+}\mapsto Ue_{0}e_{1}^{\ast}\theta^{-}+e_{1}e_{0}^{\ast}\theta^{-},\ e_{0}e_{1}^{\ast}\theta^{+}\mapsto U^{-1}e_{1}e_{0}^{\ast}\theta^{-}+e_{0}e_{1}^{\ast}\theta^{-}.\end{gathered}$
In each case, the first summand is the contribution of $D_{1}$; the second summand is that of $D_{2}$. Note that the differential is well defined on $\boldsymbol{CF}^{\infty}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$, but not well defined on $\boldsymbol{CF}^{-}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$ as we get a negative power of $U$. Our solution (see Subsubsection 2.3.2) will be to consider the free submodule $\boldsymbol{CF}_{fil}^{-}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})\leq\boldsymbol{CF}^{-}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$ generated by $Ue_{0}e_{1}^{\ast}\theta^{+}$ together with all $e_{i}e_{j}^{\ast}\theta^{+},e_{i}e_{j}^{\ast}\theta^{-}$ for $i,j=0,1$ except $e_{0}e_{1}^{\ast}\theta^{+}$ (i.e. replace $e_{0}e_{1}^{\ast}\theta^{+}$ by $Ue_{0}e_{1}^{\ast}\theta^{+}$). Then, $\boldsymbol{CF}_{fil}^{-}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$ will be a sub chain complex of $\boldsymbol{CF}^{\infty}(\alpha^{E_{\alpha}},\beta^{E_{\beta}})$.
Figure 2.2. Some genus 1 examples
Remark 2.16.
One can define the composition maps $\mu_{n}$ without mentioning local systems, if there is at most one attaching curve with a nontrivial local system. Let us demonstrate this for $\mu_{1}$ (we have observed a special case in the first example of Example 2.15). Assume that exactly one of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ has a nontrivial local system, and let the coefficient ring be $S=\mathbb{F}\llbracket U^{1/2},X_{1},X_{2},\cdots,X_{k}\rrbracket$. Then $\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is $S$-linearly isomorphic to the chain complex $\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$, where $R=\mathbb{F}\llbracket U^{1/2},X_{1},X_{2},\cdots,X_{k}\rrbracket$, and we add $w:=\partial_{+}G$ to the set of basepoints, assign the weight $U^{1/2}$ to $w$ and $z$, and assign the same weights as before to all the basepoints except $w$ and $z$. The $S$-linear chain isomorphism is given by mapping $e_{0}{\bf x}\mapsto{\bf x}$, $e_{1}{\bf x}\mapsto U^{1/2}{\bf x}$ if $E_{\boldsymbol{\beta}}$ is nontrivial; and $e_{1}^{\ast}{\bf x}\mapsto{\bf x}$, $e_{0}^{\ast}{\bf x}\mapsto U^{1/2}{\bf x}$ is $E_{\boldsymbol{\alpha}}$ is nontrivial.
We will see an instance of this for $\mu_{2}$ in Subsubsection 3.3.2: compare Definition 3.21 and Remark 3.22.
2.3. Weak admissibility and positivity
We mainly work in the minus version. Recall that in Subsection 2.2, we have specialized to specific kinds of Heegaard data, which involved local systems that a priori involve negative powers of $U$. Under these assumptions, we first deal with admissibility, which ensures that the sums in Equations 2.1 and 2.2 are well-defined in $\boldsymbol{CF}^{\infty}$.
We then deal with positivity, i.e. we show that our maps only involve nonnegative powers of $U$. More precisely, we define a $\boldsymbol{CF}^{-}$-type sub chain complex of $\boldsymbol{CF}^{\infty}$, which we call the space of filtered maps $\boldsymbol{CF}_{fil}^{-}$ as in [Zem23], on which the other composition maps ($\mu_{n}$ for $n\geq 2$) are defined as well. By $\boldsymbol{CF}^{-}$-type, we mean that $\boldsymbol{CF}_{fil}^{-}$ is a (free) sub $R$-module of $\boldsymbol{CF}^{-}$, and that $R^{\infty}\boldsymbol{CF}_{fil}^{-}=\boldsymbol{CF}^{\infty}$. In our case, positivity is ensured because we have at most two attaching curves with nontrivial local systems [^22] When we prove the exact triangle, we need to consider two, and fortunately only two.. If at most one of $E_{\boldsymbol{\alpha}}$ or $E_{\boldsymbol{\beta}}$ is a trivial local system, then we will have $\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})=\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$.
We first record a useful lemma.
Definition 2.17.
Define a grading
${\rm gr}_{U}:{\rm Hom}_{R}(A,B)\backslash\{0\}\to\frac{1}{2}\mathbb{Z}$
for $A,B\in\{E^{\infty},R^{\infty}\}$ as follows.
First, for $r\in R^{\infty}\backslash\{0\}$, let ${\rm gr}_{U}(r)$ be the largest integer $n$ such that $r\in U^{n}R[X_{1}^{-1},\cdots,X_{k}^{-1}]$, and define ${\rm gr}_{U}:{\rm Hom}_{R}(R^{\infty},R^{\infty})\backslash\{0\}\to\mathbb{Z}$ by identifying ${\rm Hom}_{R}(R^{\infty},R^{\infty})\simeq R^{\infty}$.
Then, define ${\rm gr}_{U}$ for the basis elements $e_{0},e_{1},e_{0}^{\ast},e_{1}^{\ast},e_{0}e_{0}^{\ast},\cdots$ such that it is multiplicative and ${\rm gr}_{U}e_{0}=0$, ${\rm gr}_{U}e_{1}=1/2$ (for instance, ${\rm gr}_{U}e_{i}^{\ast}=-1/2$). For $\sum r_{i}f_{i}\in{\rm Hom}_{R}(A,B)$ where $f_{i}$ are distinct basis elements and $r_{i}\in R^{\infty}\backslash\{0\}$, let
${\rm gr}_{U}\left(\sum r_{i}f_{i}\right):=\min_{i}({\rm gr}_{U}(r_{i}){\rm gr}_{U}(f_{i})).$
Figure 2.3. A local diagram for Lemmas 2.18 and 2.23
Lemma 2.18.
Let $\boldsymbol{\alpha}_{0}^{E_{0}},\cdots,\boldsymbol{\alpha}_{d}^{E_{d}}$ be attaching curves with local systems. Consider generators $f_{i}{\bf x}_{i}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i-1}^{E_{i-1}},\boldsymbol{\alpha}_{i}^{E_{i}})$ for $i=1,\cdots,d$ and ${\bf y}\in\boldsymbol{\alpha}_{0}\cap\boldsymbol{\alpha}_{d}$. Assume that there is a domain ${\mathcal{}D}\in D({\bf x}_{1},\cdots,{\bf x}_{d},{\bf y})$, and let
$g=w({\mathcal{}D})\rho({\mathcal{}D})(f_{1}\otimes\cdots\otimes f_{d}).$
If $g\neq 0$, then
${\rm gr}_{U}g=\frac{1}{2}(n_{w}({\mathcal{}D})+n_{z}({\mathcal{}D}))+\sum{\rm gr}_{U}f_{i},$
where $w=\partial_{+}G$ and $z=\partial_{-}G$.
Proof.
This follows from inspecting Figure 2.3. Note that for this claim, we do not need that only at most two of $\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{d}$ have nontrivial local systems. ∎
2.3.1. Weak admissibility
There are two kinds of admissibility conditions: strong admissibility and weak admissibility. Strong admissibility lets us work with polynomial coefficient rings by ensuring that for each $m$, there are only finitely many domains with Maslov index $m$, if we restrict to a specific ${\rm Spin}^{c}$-structure. Weak admissibility lets us work with power series rings $R=\mathbb{F}\llbracket Y_{1},\cdots,Y_{k}\rrbracket$ by ensuring that there are only finitely many domains that contribute when the coefficient ring is $R/(Y_{1}^{N},\cdots,Y_{k}^{N})$, for each $N$ [^23] We have not discussed positivity yet, so to be precise, we can only talk about $\boldsymbol{CF}^{\infty}$; in this case, weak admissibility ensures a finite count for each $N$, if the coefficient module is $R^{\infty}/(Y_{1}^{N}R+\cdots+Y_{k}^{N}R)$..
We discuss strong admissibility when there are two attaching curves in Subsubsection 2.8.3; for more than two attaching curves, we only study a restrictive case in Subsubsection 2.10.1. We can work with either weak or strong admissibility for the purpose of this paper if we only consider two attaching curves, but when there are more than two, we will mainly work with weak admissibility, since in some cases, we have to consider the sum of infinitely many polygon maps, as in [OS04b, Section 9].
Definition 2.19.
Let $\boldsymbol{q}$ be the set of basepoints together with $\partial_{+}G$. The Heegaard diagram is weakly admissible if every cornerless two-chain ${\mathcal{}D}$ such that $n_{\boldsymbol{q}}({\mathcal{}D})=0$ has both positive and negative local multiplicities.
Lemma 2.20.
If the Heegaard diagram of a Heegaard datum is weakly admissible, then the sums in Equations 2.1 and 2.2 are well-defined for $\boldsymbol{CF}^{\infty}$.
Proof.
Fix generators $f_{i}{\bf x}_{i}\in\boldsymbol{CF}^{\infty}(\boldsymbol{\alpha}_{i-1}^{E_{i-1}},\boldsymbol{\alpha}_{i}^{E_{i}})$ for $i=1,\cdots,d$ and ${\bf y}\in\boldsymbol{\alpha}_{0}\cap\boldsymbol{\alpha}_{d}$. For each $M\in\mathbb{Z}$, we would like to show that there are only finitely many nonnegative domains ${\mathcal{}D}\in D({\bf x}_{1},\cdots,{\bf x}_{d},{\bf y})$ such that
$g\notin U^{M}{\rm Hom}_{R}(E_{0},E_{d})+X_{1}^{M}{\rm Hom}_{R}(E_{0},E_{d})+\cdots+X_{k}^{M}{\rm Hom}_{R}(E_{0},E_{d}),$
where $g=w({\mathcal{}D})\rho({\mathcal{}D})(f_{1}\otimes\cdots\otimes f_{d})$ [^24] Here, identify ${\rm Hom}_{R}(E_{i}^{\infty},E_{j}^{\infty})\simeq{\rm Hom}_{R}(E_{i},E_{j})\otimes_{R}R^{\infty}$..
The proof of [OS04c, Lemma 4.13] shows that for each $N$, there are only finitely many nonnegative $\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{d}$-domains ${\mathcal{}D}$ such that for all $q\in\boldsymbol{q}$, we have $n_{q}({\mathcal{}D})\leq N$. Hence the above follows, from Lemma 2.18 and that $g\in U^{\left\lfloor{\rm gr}_{U}g\right\rfloor}{\rm Hom}_{R}(E_{0},E_{d})$. ∎
We can now define $\boldsymbol{CF}^{\infty}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ and show that it is a chain complex. Similarly, one can check that the higher $A_{\infty}$-relations hold: these are simpler since there are no contributions from boundary degenerations.
Lemma 2.21.
Let us consider a weakly admissible Heegaard datum with attaching curves with local systems $\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}$. The map $\partial$ on $\boldsymbol{CF}^{\infty}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ satisfies $\partial^{2}=0$.
Proof.
The proof follows the usual proof of $\partial^{2}=0$, which studies the ends of the moduli spaces ${\mathcal{}M}({\mathcal{}D})$ for domains in ${\mathcal{}D}({\bf x},{\bf y})$ with Maslov index $2$. The ends consist of broken flowlines and boundary degenerations, and so $\partial^{2}$ equals the contribution from the boundary degenerations. See [OS08a, Theorem 5.5], [Zem17, Lemma 2.1], [HHSZ20, Proposition 7.9]. (Our case is algebraically slightly more complicated as we have nontrivial local systems.)
Let the Heegaard surface be $\Sigma$, and let the coefficient ring be $R^{\infty}$ where $R=\mathbb{F}\llbracket U,X_{1},\cdots,X_{k}\rrbracket$ (as in Subsection 2.2, in particular $X_{i}=U_{i}$ or $U_{i}^{1/2}$), where $U$ is the distinguished variable. Let ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$, and let us consider domains ${\mathcal{}D}\in D({\bf x},{\bf x})$. We claim that
$\sum_{{\mathcal{}D}}w({\mathcal{}D})\rho({\mathcal{}D})=\left(U+\sum_{i}U_{i}\right){\rm Id}_{{\rm Hom}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})},$
where the two-chain of ${\mathcal{}D}$ ranges through the connected components of $\Sigma\backslash\boldsymbol{\alpha}$, or the connected components of $\Sigma\backslash\boldsymbol{\beta}$. This is almost exactly like the case where all the local systems are trivial, but with the added complication that $\partial{\mathcal{}D}$ might intersect $G$. However, this is not possible as the two endpoints of $G$ lie in the same connected component of $\Sigma\backslash\boldsymbol{\alpha}$ (and $\Sigma\backslash\boldsymbol{\beta}$).
Let #$\#\widetilde{N}({\mathcal{}D},{\bf x})$ be the number of boundary degenerations with two-chain ${\mathcal{}D}$ and vertex ${\bf x}$. We have
$\partial^{2}(e{\bf x})=\sum_{{\mathcal{}D}}\#\widetilde{N}({\mathcal{}D},{\bf x})w({\mathcal{}D})\rho({\mathcal{}D})(e){\bf x},$
and this is zero since #$\#\widetilde{N}({\mathcal{}D},{\bf x})$ does not depend on ${\mathcal{}D}$ (and ${\bf x}$). ∎
2.3.2. Positivity
We define the space of filtered maps.
Definition 2.22.
Define the space of filtered maps, $\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})\leq\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$, as follows.
  • If either $\boldsymbol{\alpha}$ or $\boldsymbol{\beta}$ has the trivial local system, then every element of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is filtered.
  • If both $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ have nontrivial local systems, then $\boldsymbol{y}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is filtered if for each summand $f{\bf x}$ of $\boldsymbol{y}$ such that $\theta^{+}\in{\bf x}$, we have that ${\rm gr}_{U}f\geq 0$.
In other words, we only rule out $e_{0}e_{1}^{\ast}{\bf x}$ (but allow $Ue_{0}e_{1}^{\ast}{\bf x}$) for ${\bf x}$ such that $\theta^{+}\in{\bf x}$.
The following is the main lemma that ensures positivity. (Compare [Zem23, Lemmas A.3, A.5])
Lemma 2.23.
Let $\boldsymbol{\alpha}_{0}^{E_{0}},\cdots,\boldsymbol{\alpha}_{d}^{E_{d}}$ be attaching curves with local systems. Assume that at most two of them have nontrivial local systems. Consider generators $f_{i}{\bf x}_{i}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i-1}^{E_{i-1}},\boldsymbol{\alpha}_{i}^{E_{i}})$ for $i=1,\cdots,d$ and ${\bf y}\in\boldsymbol{\alpha}_{0}\cap\boldsymbol{\alpha}_{d}$. If there exists a nonnegative domain ${\mathcal{}D}\in D({\bf x}_{1},\cdots,{\bf x}_{d},{\bf y})$, let
$g=w({\mathcal{}D})\rho({\mathcal{}D})(f_{1}\otimes\cdots\otimes f_{d}).$
If $f_{1}{\bf x}_{1},\cdots,f_{d}{\bf x}_{d}$ are filtered, then $g{\bf y}$ is filtered.
Proof.
There is nothing to show if all the local systems are trivial. Let us assume that $\boldsymbol{\alpha}_{a}$ and $\boldsymbol{\alpha}_{b}$ ($a\leq b$) have nontrivial local systems, and that the rest have trivial local systems. Also, assume that $g\neq 0$. Let $w=\partial_{+}G$, $z=\partial_{-}G$ as in Figure 2.3. We will use the following repeatedly: if $n_{w}({\mathcal{}D})=n_{z}({\mathcal{}D})=0$ and the boundary $\partial_{\boldsymbol{\alpha}_{a}}{\mathcal{}D}$ or $\partial_{\boldsymbol{\alpha}_{b}}{\mathcal{}D}$ intersects $G$, then $n_{A}({\mathcal{}D})=n_{C}({\mathcal{}D})=0$ and $n_{B}({\mathcal{}D})=1$. Hence, $a=0$, $b=d=1$, $\theta^{+}\in{\bf x}_{1}$, and $\theta^{-}\in{\bf y}$. Note that in this case, ${\rm gr}_{U}g={\rm gr}_{U}f_{1}\geq 0$, and so $g{\bf y}$ is filtered.
We divide into a few cases. If $b\neq d$, then we we are done if ${\rm gr}_{U}g\geq-1/2$. If ${\rm gr}_{U}g<-1/2$, then it must be the case that ${\rm gr}_{U}f_{a+1}={\rm gr}_{U}f_{b+1}=-1/2$, ${\rm gr}_{U}f_{i}=0$ for all $i\neq a+1,b+1$, and $n_{w}({\mathcal{}D})=n_{z}({\mathcal{}D})=0$. Hence, $f_{b}=e_{0}$ or $e_{0}e_{1}^{\ast}$, and $f_{b+1}=e_{1}^{\ast}$. Since $g{\bf y}\neq 0$, the $\boldsymbol{\alpha}_{b}$-boundary $\partial_{\boldsymbol{\alpha}_{b}}{\mathcal{}D}$ intersects $G$.
If $b=d$, then we are done if ${\rm gr}_{U}g\geq 0$. Hence, assume that ${\rm gr}_{U}g<0$. Then it must be the case that ${\rm gr}_{U}f_{a+1}=-1/2$ (and so $f_{a+1}=e_{1}^{\ast}$ or $e_{0}e_{1}^{\ast}$), and ${\rm gr}_{U}f_{i}=0$ for all $i\neq a+1$, and $n_{w}({\mathcal{}D})=n_{z}({\mathcal{}D})=0$. Hence, we can assume that $\partial_{\boldsymbol{\alpha}_{a}}{\mathcal{}D}$ and $\partial_{\boldsymbol{\alpha}_{b}}{\mathcal{}D}$ do not intersect $G$.
  • If $a\neq 0$, then $f_{a}=e_{0}$, and so $\partial_{\boldsymbol{\alpha}_{a}}{\mathcal{}D}$ intersects $G$.
  • If $a=0$ and $\theta^{+}\in{\bf y}$, then $n_{z}({\mathcal{}D})+n_{w}({\mathcal{}D})=n_{B}({\mathcal{}D})+n_{C}({\mathcal{}D})+1\geq 1$.
  • If $a=0$ and $\theta^{-}\in{\bf y}$, then since $\partial_{\boldsymbol{\alpha}_{a}}{\mathcal{}D}$ and $\partial_{\boldsymbol{\alpha}_{b}}{\mathcal{}D}$ do not intersect $G$, $g=e_{0}e_{1}^{\ast}$. Hence $g{\bf y}$ is filtered.
As everything is well-defined, we can now show that $\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is a chain complex, and similarly that the higher $A_{\infty}$-relations hold.
Corollary 2.24.
Let us consider a weakly admissible Heegaard datum with attaching curves with local systems $\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}$. Then $\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ together with $\partial$ is a chain complex.
Proof.
Lemma 2.21 shows that $(\boldsymbol{CF}^{\infty}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}),\partial)$ is a chain complex, and Lemma 2.23 shows that $\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is a sub chain complex of $\boldsymbol{CF}^{\infty}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$. ∎
By abuse of notation, we mean $\boldsymbol{CF}_{fil}^{-}$ whenever we write $\boldsymbol{CF}^{-}$ from now on.
2.4. Twisted complexes
We use the language of twisted complexes in order to nicely package the necessary algebra. We mostly follow [Zem23, Subsection 2.3] and [Sei08, Sections 1b, 3k, 3l, and 3m]. Also compare [LOT16].
2.4.1. $A_{\infty}$-categories and twisted complexes
We work with non-unital $A_{\infty}$-categories over $\mathbb{F}$. Our conventions are slightly different from [Sei08], due to the nature of Heegaard Floer homology: we use the homological convention instead of the cohomological convention, and $A_{\infty}$-categories are not necessarily graded. However, the $A_{\infty}$-categories we consider will be graded in some special cases: we say that an $A_{\infty}$-category is graded if the Hom spaces ${\rm Hom}(\alpha,\beta)$ are $\mathbb{Z}$-graded and the composition maps $\mu_{d}$ have degree $d-2$. When everything is graded, we use the convention that for a homogeneous $f\in{\rm Hom}(\alpha,\beta)$, the corresponding element in ${\rm Hom}(\alpha[k],\beta[l])$ has degree $\deg f+k-l$. In other words, $[k]$ shifts the grading down by $k$.
Remark 2.25.
If one wants to justify the use of the language of $A_{\infty}$-categories, the following is one unsatisfying way that works for this paper. Let us say that we have a weakly admissible Heegaard datum where the attaching curves together with the local systems are $\boldsymbol{\alpha}_{0}^{E_{0}},\cdots,\boldsymbol{\alpha}_{m}^{E_{m}}$, such that the local systems $E_{i}$ are trivial for all but at most two $i$’s. Also assume that we have made the analytic choices necessary to define the composition maps in Heegaard Floer theory. Then, the $A_{\infty}$-category we work with has objects $\boldsymbol{\alpha}_{0}^{E_{0}},\cdots,\boldsymbol{\alpha}_{m}^{E_{m}}$, and the Hom spaces and composition maps $\mu_{d}$ are nontrivial only if the attaching curves are in the correct order, i.e.
${\rm Hom}(\boldsymbol{\alpha}_{i}^{E_{i}},\boldsymbol{\alpha}_{j}^{E_{j}})=\begin{cases}\boldsymbol{CF}_{fil}^{-}(\boldsymbol{\alpha}_{i}^{E_{i}},\boldsymbol{\alpha}_{j}^{E_{j}})&{\rm if}\ i<j\\ 0&{\rm if}\ i\geq j\end{cases}.$
When we write $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$, it will be implicit that $\boldsymbol{\alpha}$ comes before $\boldsymbol{\beta}$.
Let ${\mathcal{}A}$ be an $A_{\infty}$-category. Then, the additive enlargement $\Sigma{\mathcal{}A}$ is defined as follows. Objects are formal direct sums
$\bigoplus_{i\in I}V_{i}\otimes\alpha_{i}$
where $I$ is a finite set, $\{\alpha_{i}\}$ is a family of objects of ${\mathcal{}A}$, and $\{V_{i}\}$ is a family of finite-dimensional vector spaces. Morphisms are defined as a combination of morphisms between the vector spaces and morphisms in ${\mathcal{}A}$:
${\rm Hom}_{\Sigma{\mathcal{}A}}\left(\bigoplus_{i\in I}V_{i}\otimes\alpha_{i},\bigoplus_{j\in J}W_{j}\otimes\beta_{j}\right)=\bigoplus_{i,j}{\rm Hom}_{\mathbb{F}}(V_{i},W_{j})\otimes{\rm Hom}_{{\mathcal{}A}}(\alpha_{i},\beta_{j}).$
Similarly, compositions $\mu_{d}$ are defined by combining the ordinary composition of maps between vector spaces and morphisms in ${\mathcal{}A}$. The additive enlargment $\Sigma{\mathcal{}A}$ forms an $A_{\infty}$-category.
A twisted complex in ${\mathcal{}A}$ consists of an object $\underline{\alpha}\in{\rm Ob}(\Sigma{\mathcal{}A})$ together with a differential $\delta_{\underline{\alpha}}\in{\rm Hom}_{\Sigma{\mathcal{}A}}\left(\underline{\alpha},\underline{\alpha}\right)$, such that
$\sum_{n\geq 1}\mu_{n}^{\Sigma{\mathcal{}A}}(\delta_{\underline{\alpha}},\cdots,\delta_{\underline{\alpha}})=0.$
Seidel assumes that the differential $\delta_{\underline{\alpha}}$ is strictly lower triangular with respect to a filtration on $\underline{\alpha}$ to ensure that the above sum is finite.
Twisted complexes in ${\mathcal{}A}$ also form an $A_{\infty}$-category ${\rm Tw}{\mathcal{}A}$. Morphisms between twisted complexes are the same as before,
${\rm Hom}_{{\rm Tw}{\mathcal{}A}}\left((\underline{\alpha},\delta_{\underline{\alpha}}),(\underline{\beta},\delta_{\underline{\beta}})\right)={\rm Hom}_{\Sigma{\mathcal{}A}}(\underline{\alpha},\underline{\beta}),$
but the compositions are different:
$\mu_{d}^{{\rm Tw}{\mathcal{}A}}(\underline{\varphi_{1}},\cdots,\underline{\varphi_{d}})=\sum_{i_{0},\dots,i_{d}\geq 0}\mu_{d+i_{0}+\cdots+i_{d}}^{\Sigma{\mathcal{}A}}(\overbrace{\delta_{\underline{\alpha_{0}}},\cdots,\delta_{\underline{\alpha_{0}}}}^{i_{0}},\underline{\varphi_{1}},\overbrace{\delta_{\underline{\alpha_{1}}},\cdots,\delta_{\underline{\alpha_{1}}}}^{i_{1}},\underline{\varphi_{2}},\cdots,\underline{\varphi_{d}},\overbrace{\delta_{\underline{\alpha_{d}}},\cdots,\delta_{\underline{\alpha_{d}}}}^{i_{d}}).$
Remark 2.26.
If ${\mathcal{}A}$ is a graded $A_{\infty}$-category, then its graded additive enlargement $\Sigma{\mathcal{}A}$ is defined exactly the same, except that the finite-dimensional vector spaces $V_{i}$ are also $\mathbb{Z}$-graded. It forms a graded $A_{\infty}$-category. Similarly, graded twisted complexes are defined exactly the same, except that we require that the differential has degree $-1$. Morphisms between twisted complexes and their compositions are defined exactly the same. The $A_{\infty}$-category of graded twisted complexes, ${\rm Tw}{\mathcal{}A}$, is graded.
Example 2.27.
  • Given an object $\alpha$ of ${\mathcal{}A}$, $\alpha$ together with the zero differential $\delta_{\alpha}=0$ form a twisted complex in ${\mathcal{}A}$.
  • Given objects $\beta_{1},\beta_{2}$ of ${\mathcal{}A}$ and a morphism $\psi:\beta_{1}\to\beta_{2}$ such that $\mu_{1}^{{\mathcal{}A}}(\psi)=0$, the mapping cone of $\psi$ is the twisted complex $(\underline{\beta},\delta_{\underline{\beta}})$ given by $\underline{\beta}=\beta_{1}\oplus\beta_{2}$ and $\delta_{\underline{\beta}}=\psi$. We write this as
[…]
Notation 2.28.
As above, we will usually underline twisted complexes and maps between them. An exception is when the twisted complex comes from a single object $\alpha$, as in the first example of Example 2.27. Also, if $\alpha_{0}$ and $\beta_{0}$ are summands of the twisted complexes $\underline{\alpha}$ and $\underline{\beta}$, respectively, and $\theta:\alpha_{0}\to\beta_{0}$ is a map, then we denote the map $\underline{\alpha}\to\underline{\beta}$ whose $\alpha_{0}\to\beta_{0}$ component is $\theta$ and all the other components are 00 as $\underline{\theta}:\underline{\alpha}\to\underline{\beta}$.
2.4.2. $A_{\infty}$-functors and twisted complexes
We will define twisted complexes of attaching curves in one almost complex structure, and deform the almost complex structure. To do this, we have to understand how the $A_{\infty}$-functor induced by deforming the almost complex structure extends to the category of twisted complexes.
Recall that an $A_{\infty}$-functor ${\mathcal{}F}:{\mathcal{}A}\to{\mathcal{}B}$ between two $A_{\infty}$-categories is a collection of a map ${\rm Ob}{\mathcal{}F}:{\rm Ob}{\mathcal{}A}\to{\rm Ob}{\mathcal{}B}$ and multilinear maps ${\mathcal{}F}_{d}:{\rm Hom}_{{\mathcal{}A}}(\alpha_{0},\alpha_{1})\otimes\cdots\otimes{\rm Hom}_{{\mathcal{}A}}(\alpha_{d-1},\alpha_{d})\to{\rm Hom}_{{\mathcal{}B}}({\mathcal{}F}(\alpha_{0}),{\mathcal{}F}(\alpha_{d}))$ for $d\geq 1$ that satisfy the appropriate $A_{\infty}$-relations. If the $A_{\infty}$-categories ${\mathcal{}A}$ and ${\mathcal{}B}$ are graded, then ${\mathcal{}F}$ is graded if ${\mathcal{}F}_{d}$ has degree $d-1$ for all $d\geq 1$.
Given an $A_{\infty}$-functor ${\mathcal{}F}:{\mathcal{}A}\to{\mathcal{}B}$ between $A_{\infty}$-categories, the induced $A_{\infty}$-functor $\Sigma{\mathcal{}F}:\Sigma{\mathcal{}A}\to\Sigma{\mathcal{}B}$ is defined as follows: on the objects,
$\left(\Sigma{\mathcal{}F}\right)\left(\bigoplus_{i\in I}V_{i}\otimes\alpha_{i}\right)=\bigoplus_{i\in I}V_{i}\otimes{\mathcal{}F}(\alpha_{i}).$
On the morphisms, $\left(\Sigma{\mathcal{}F}\right)_{d}(\underline{\varphi_{1}},\cdots,\underline{\varphi_{d}})$ is defined analogously to $\mu_{d}^{\Sigma{\mathcal{}A}}$, where we combine ordinary composition of maps between vector spaces and ${\mathcal{}F}_{d}$.
The induced $A_{\infty}$-functor ${\rm Tw}{\mathcal{}F}:{\rm Tw}{\mathcal{}A}\to{\rm Tw}{\mathcal{}B}$ is defined as follows:
$\left({\rm Tw}{\mathcal{}F}\right)\left(\underline{\alpha},\delta_{\underline{\alpha}}\right)$ $=\left(\left(\Sigma{\mathcal{}F}\right)\left(\underline{\alpha}\right),\sum_{n\geq 1}\left(\Sigma{\mathcal{}F}\right)_{n}\left(\delta_{\underline{\alpha}},\cdots,\delta_{\underline{\alpha}}\right)\right),$
$\left({\rm Tw}F\right)_{d}(\underline{\varphi_{1}},\cdots,\underline{\varphi_{d}})$ $=\sum_{i_{0},\dots,i_{d}\geq 0}\left(\Sigma{\mathcal{}F}\right)_{d+i_{0}+\cdots+i_{d}}(\overbrace{\delta_{\underline{\alpha_{0}}},\cdots,\delta_{\underline{\alpha_{0}}}}^{i_{0}},\underline{\varphi_{1}},$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \overbrace{\delta_{\underline{\alpha_{1}}},\cdots,\delta_{\underline{\alpha_{1}}}}^{i_{1}},\underline{\varphi_{2}},\cdots,\underline{\varphi_{d}},\overbrace{\delta_{\underline{\alpha_{d}}},\cdots,\delta_{\underline{\alpha_{d}}}}^{i_{d}}).$
If everything is graded, then ${\rm Tw}{\mathcal{}F}$ is graded as well.
Remark 2.29.
As noted above, we will define twisted complexes of attaching curves and maps between them in one almost complex structure, and will continue talking about them in a different almost complex structure. The latter should be interpreted as the image under ${\rm Tw}{\mathcal{}F}$, where ${\mathcal{}F}$ is the $A_{\infty}$-functor given by deforming the almost complex structure. Note that ${\mathcal{}F}$ depends on the choice of the deformation. Fortunately, it will turn out that all the twisted complexes $\underline{\boldsymbol{\beta}}$ that we consider in this paper are “invariant” under ${\rm Tw}{\mathcal{}F}$. Furthermore, although the maps between the twisted complexes that we consider might depend on the deformation, it will not matter since we will not care about exactly what these maps are.
2.4.3. Exact triangles and twisted complexes
[…]
$\cdots\to H(C_{1})\xrightarrow{f_{\ast}}H(C_{2})\to H(C_{3})\to H(C_{1})\to\cdots.$
In this paper, we care about what the maps in this exact triangle are. Compare with the triangle detection lemma [OS05, Lemma 4.2].
Lemma 2.30.
Let the following be a twisted complex
[…]
If the homology of ${\rm Hom}(\alpha,\underline{\beta})$ is 00, then the following sequence of homology groups is exact.
$H({\rm Hom}(\alpha,\beta_{1}))\xrightarrow{\mu_{2}(-,f)}H({\rm Hom}(\alpha,\beta_{2}))\xrightarrow{\mu_{2}(-,g)}H({\rm Hom}(\alpha,\beta_{3}))$
Proof.
One can show this using a diagram chasing argument and the definition of ${\rm Hom}(\alpha,\underline{\beta})$. ∎
Remark 2.31.
[…]
2.5. Standard translates
Our main “local computation,” Theorem 4.3, morally says that two specific maps $\underline{f}:\underline{\boldsymbol{\beta}_{1}}\to\underline{\boldsymbol{\beta}_{2}}$ and $\underline{g}:\underline{\boldsymbol{\beta}_{2}}\to\underline{\boldsymbol{\beta}_{1}}$ are homotopy inverses of each other. However, the $A_{\infty}$-categories we work in are not unital: we do not even define $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\alpha})$. Instead of trying to define units and homotopy inverses, we use standard translates and use some ad-hoc arguments. The definitions in this subsection should be treated as useful shorthands rather than actual definitions.
Figure 2.4. Standard translates of $\beta$
Definition 2.32.
Let $\boldsymbol{\beta}=\{\beta^{1},\cdots,\beta^{n}\}$ be an attaching curve. A standard translate $\boldsymbol{\beta}^{\prime}=\{\beta^{\prime}{}^{1},\cdots,\beta^{\prime}{}^{n}\}$ of $\boldsymbol{\beta}$ is given by slightly translating each circle $\beta^{i}$, such that #$\#(\beta^{i}\cap\beta^{\prime}{}^{j})=2\delta_{ij}$. If we consider other attaching curves $\boldsymbol{\alpha}$, then we assume that the translate is sufficiently small in the sense that $\boldsymbol{\alpha}$ meets $\boldsymbol{\beta}$ and $\boldsymbol{\beta}^{\prime}$ in the same way (see Figure 2.4). Denote the top grading generator of $\boldsymbol{CF}^{-}(\boldsymbol{\beta},\boldsymbol{\beta}^{\prime})$ as $\Theta_{\boldsymbol{\beta}}^{+}$, or simply $\Theta^{+}$. If $\boldsymbol{\beta}$ intersects the oriented arc $G$ (i.e. has a nontrivial local system), then the translate is small enough so that $G$ also meets $\boldsymbol{\beta}$ and $\boldsymbol{\beta}^{\prime}$ in the same way. We will sometimes denote the generator ${\rm Id}_{E}\Theta_{\boldsymbol{\beta}}^{+}$ of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}^{E},\boldsymbol{\beta}^{\prime}{}^{E})$ as $\Theta_{\boldsymbol{\beta}}^{+}$ or $\Theta^{+}$.
Given a map $f:\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}}\to\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}$, define the corresponding maps $f^{\prime}:a\to b$ for $a\in\{\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\alpha}^{\prime}{}^{E_{\boldsymbol{\alpha}^{\prime}}}\}$ and $b\in\{\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}},\boldsymbol{\beta}^{\prime}{}^{E_{\boldsymbol{\beta}^{\prime}}}\}$ by taking the nearest intersection points. Using these, define standard translates $\underline{\boldsymbol{\alpha}^{\prime}}$ of twisted complexes $\underline{\boldsymbol{\alpha}}$ of attaching curves (see Remark 2.33 for a caveat), and given a map $\underline{f}:\underline{\boldsymbol{\alpha}}\to\underline{\boldsymbol{\beta}}$, define the corresponding maps $\underline{f^{\prime}}:a\to b$ for $a\in\{\underline{\boldsymbol{\alpha}},\underline{\boldsymbol{\alpha}^{\prime}}\}$ and $b\in\{\underline{\boldsymbol{\beta}},\underline{\boldsymbol{\beta}^{\prime}}\}$ as well.
Similarly, if $\underline{\boldsymbol{\beta}}$ is a twisted complex, then denote the map $\underline{\boldsymbol{\beta}}\to\underline{\boldsymbol{\beta}^{\prime}}$ given by $\Theta_{\boldsymbol{\beta}}^{+}:\boldsymbol{\beta}\to\boldsymbol{\beta}^{\prime}$ for each summand $\boldsymbol{\beta}$ as $\Theta_{\underline{\boldsymbol{\beta}}}^{+}$, or simply $\Theta^{+}\in\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}},\underline{\boldsymbol{\beta}^{\prime}})$.
Remark 2.33.
We do not claim that the induced map $f^{\prime}$, resp., $\underline{f^{\prime}}$ is always a cycle if $f$, resp., $\underline{f}$ is a cycle, that $\underline{\boldsymbol{\alpha}^{\prime}}$ is always a twisted complex, and that $\Theta_{\underline{\boldsymbol{\beta}}}^{+}$ is a cycle. However, they will be for all the cases that we consider in this paper.
Instead of working with homotopy equivalences (and homotopy inverses), we use the following and work with quasi-isomorphisms in the sense of Definition 2.37  [^25] By Yoneda, they should be the same; but our $A_{\infty}$-categories do not even have ${\rm Hom}(\alpha,\alpha)$..
Lemma 2.34.
Let us consider a weakly admissible Heegaard datum. Consider a twisted complex $\underline{\boldsymbol{\beta}}$ of attaching curves with local systems and its small translate $\underline{\boldsymbol{\beta}^{\prime}}$, with the same filtration. If $\underline{\boldsymbol{\beta}^{\prime}}$ is a twisted complex, and if $\underline{\Theta}:\underline{\boldsymbol{\beta}}\to\underline{\boldsymbol{\beta}^{\prime}}$ is a cycle that respects the filtration (i.e. lower triangular), and for each summand $\boldsymbol{\beta}$ of $\underline{\boldsymbol{\beta}}$, the component of $\underline{\Theta}$ that is from $\boldsymbol{\beta}$ to $\boldsymbol{\beta}^{\prime}$ is homotopic to $\Theta_{\boldsymbol{\beta}}^{+}:\boldsymbol{\beta}\to\boldsymbol{\beta}^{\prime}$, then for any other attaching curve $\boldsymbol{\alpha}$, the induced map
$\mu_{2}(-,\underline{\Theta}):\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}})\to\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}^{\prime}})$
is a quasi-isomorphism.
Proof.
The filtrations on $\underline{\boldsymbol{\beta}},\underline{\boldsymbol{\beta}^{\prime}}$ induce a filtration on the chain complexes $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}})$ and $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}^{\prime}})$. With respect to these filtrations, $\mu_{2}(-,\underline{\Theta})$ is lower triangular, and the induced maps on the successive quotients are quasi-isomorphisms. ∎
Remark 2.35.
Note that if $\Theta_{\underline{\boldsymbol{\beta}}}^{+}:\underline{\boldsymbol{\beta}}\to\underline{\boldsymbol{\beta}^{\prime}}$ is a cycle, then it satisfies the condition of Lemma 2.34. Also, if the $A_{\infty}$-functor induced by changing the almost complex structure preserves $\underline{\boldsymbol{\beta}}$ and $\underline{\boldsymbol{\beta}^{\prime}}$, then the image of $\Theta_{\underline{\boldsymbol{\beta}}}^{+}$ satisfies the condition of Lemma 2.34.
Lemma 2.36 (Derived Nakayama).
Let $R$ be a power series ring $\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$ over a field $\mathbb{F}$, and let $C,C^{\prime}$ be two finite, free chain complexes over $R$. Then, a $R$-linear chain map $f:C\to C^{\prime}$ is a quasi-isomorphism if
$f\otimes_{R}\mathbb{F}:C\otimes_{R}\mathbb{F}\to C^{\prime}\otimes_{R}\mathbb{F}$
is a quasi-isomoprhism.
Proof.
Let $R_{k}=R/(X_{1},\cdots,X_{k})$, and consider the mapping cones $C_{k}$ of $f\otimes_{R}R_{k}$. We show by induction that the homology $H(C_{0})$ is trivial if $H(C_{n})$ is trivial.
Assume that $H(C_{k+1})$ is trivial. Consider the long exact sequence induced by the short exact sequence
$0\to C_{k}\xrightarrow{\cdot X_{k}}C_{k}\to C_{k+1}\to 0$
Since $H(C_{k+1})$ is trivial, the map $\cdot X_{k}:H(C_{k})\to H(C_{k})$ is an isomorphism, and so $H(C_{k})$ is trivial by Nakayama’s Lemma. ∎
Definition 2.37.
Two cycles $\underline{f}:\underline{\boldsymbol{\beta}_{1}}\to\underline{\boldsymbol{\beta}_{2}}$ and $\underline{g}:\underline{\boldsymbol{\beta}_{2}}\to\underline{\boldsymbol{\beta}_{1}}$ are morally quasi-inverses if the compositions $\mu_{2}(\underline{f},\underline{g^{\prime}}):\underline{\boldsymbol{\beta}_{1}}\to\underline{\boldsymbol{\beta}_{1}^{\prime}}$ and $\mu_{2}(\underline{g^{\prime}},\underline{f}):\underline{\boldsymbol{\beta}_{2}^{\prime}}\to\underline{\boldsymbol{\beta}_{2}}$ satisfy the conditions of Lemma 2.34 modulo $(X_{1},\cdots,X_{n})$, where the coefficient ring is of the form $R=\mathbb{F}\llbracket X_{1},\cdots,X_{n}\rrbracket$. If such $\underline{f},\underline{g}$ exist, then we say that $\underline{\boldsymbol{\beta}_{1}}$ and $\underline{\boldsymbol{\beta}_{2}}$ are morally quasi-isomorphic.
Similarly, $\underline{f}$ and $\underline{g}$ are morally homotopy inverses if the compositions $\mu_{2}(\underline{f},\underline{g^{\prime}})$ and $\mu_{2}(\underline{g^{\prime}},\underline{f})$ are $\Theta^{+}$. If such $\underline{f},\underline{g}$ exist, then we say that $\underline{\boldsymbol{\beta}_{1}}$ and $\underline{\boldsymbol{\beta}_{2}}$ are morally homotopy equivalent.
Note that if two maps are morally homotopy inverses, then they are morally quasi-inverses.
Remark 2.38.
Note that we do not claim that being morally quasi-isomorphic or morally homotopy equivalent is an equivalence relation, nor that it does not depend on the choice of the standard translate. In our case, we will have twisted complexes $\underline{\boldsymbol{\beta}_{i}}$ for $i=1,2,3,4$, and we will show that $\underline{\boldsymbol{\beta}_{i}}$ and $\underline{\boldsymbol{\beta}_{i+1}}$ are morally quasi-isomorphic. We will deduce that $\underline{\boldsymbol{\beta}_{1}}$ and $\underline{\boldsymbol{\beta}_{4}}$ are morally homotopy equivalent by using Lemma 2.34 and an ad-hoc argument. (There are other ways to finish off.)
2.6. Stabilizations
If the multi-Heegaard diagram is the connected sum of two Heegaard diagrams where the curves are sufficiently simple on one side, we can often reduce the computations to a computation on the other side, by stretching the neck, as in [OS04b, Theorem 9.4]. A subtlety is that we have infinitely many domains that might contribute. One could quotient out the coefficient ring by $(U_{1}^{N},\cdots,U_{l}^{N})$ for large $N$’s, in which case only finitely many domains contribute; another way is to degenerate the neck and work in a pinched Heegaard surface as in [HHSZ20, Sections 8, 9, and 10] and [HHSZ24, Section 6]. Also compare [Zem23, Subsection 6.1]. In this subsection, we consider the latter and work in a pinched Heegaard surface, but if there are only finitely many domains that could contribute, then all the statements hold for a sufficiently long neck.
Consider a Heegaard datum with Heegaard surface $\Sigma$. Let us first discuss genus stabilization, free stabilization, and free link stabilization.
Figure 2.5. Different kinds of stabilizations that we consider. $p$ is the point we connected sum along, and $z$ is a basepoint.
Genus stabilization is where the Heegaard surface $S(\Sigma)$ is given by connected summing $\Sigma$ with $\mathbb{T}^{2}$, and the attaching curves are $S(\boldsymbol{\beta})$, which are given by the attaching curve $\boldsymbol{\beta}$ on $\Sigma$ together with some fixed, essential curve, in a standard way (i.e. the fixed curves are pairwise standard translates of each other) as in Figure 2.5. The coefficient ring and weight functions are unchanged. For a generator $f{\bf x}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$, we define
$S(f{\bf x}):=f{\bf x}\times\theta^{+}\in\boldsymbol{CF}^{-}(S(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}}),S(\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}))$
where $\theta^{+}$ is the top intersection point in the connected summed $\mathbb{T}^{2}$.
Free stabilization is similar, but we connected sum with a pointed $S^{2}$ as in Figure 2.5. The coefficient ring is $R\llbracket V\rrbracket$, where $V$ is a new variable, and the weight function is the same except that we assign the weight $V$ to the new basepoint. We called this connected summed part the free stabilization region.
Free link stabilization is a slight variant of free stabilization. We connected sum with a doubly pointed $S^{2}$, the coefficient ring is $R\llbracket V^{1/2}\rrbracket$ , and we assign the weight $V^{1/2}$ to both the two new basepoints.
A Heegaard datum ${\mathcal{}H}^{stab}$ is a stabilization of a Heegaard datum ${\mathcal{}H}$ if ${\mathcal{}H}^{stab}$ is obtained from ${\mathcal{}H}$ by a sequence of genus stabilizations, free stabilizations, and free link stabilizations. To apply [HHSZ24, Proposition 6.5] as stated, we will assume that subsequent genus, free, and free link stabilizations are connected summed at the stabilized region, i.e. a stabilization of a Heegaard datum with Heegaard surface $\Sigma$ is given by connected summing with a pointed surface $\Sigma^{\prime}$ at one point [^26] This is not a serious contraint, as one can always assume this after some handleslides. [^27] See [Zem21b, Proposition 6.1] for generalized connected sums, where we “connected sum” at more than one point.. We denote the composition of the $S$ maps as $S$.
Proposition 2.39 ([HHSZ24, Proposition 6.5]).
If the connected summing region is pinched, then
$\mu_{n}(S(\boldsymbol{y}_{1}),\cdots,S(\boldsymbol{y}_{n}))=S(\mu_{n}(\boldsymbol{y}_{1},\cdots,\boldsymbol{y}_{n})).$
Another case we need is when we have a triple Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{p})$ and we connect sum with $(\mathbb{T}^{2},\alpha_{0},\beta_{0},\gamma_{0})$ where $\alpha_{0}$ is the meridian and $\beta_{0}$ and $\gamma_{0}$ are (standard translates of) the longitude, and get $(\Sigma\#\mathbb{T}^{2},T(\boldsymbol{\alpha}),T(\boldsymbol{\beta}),T(\boldsymbol{\gamma}),\boldsymbol{p})$. Note that if we only consider the $\beta$ and $\gamma$ attaching curves, then this is a genus stabilization. For a generator $f{\bf x}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ or $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\gamma}^{E_{\boldsymbol{\gamma}}})$, let $T(f{\bf x}):=f{\bf x}\times\theta$ where $\theta$ is the unique intersection point in the connected summed $\mathbb{T}^{2}$. For $\boldsymbol{y}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}},\boldsymbol{\gamma}^{E_{\boldsymbol{\gamma}}})$, let $T(\boldsymbol{y}):=S(\boldsymbol{y})$. Then, [HHSZ24, Proposition 6.5] also implies the following proposition.
Proposition 2.40 ([HHSZ24, Proposition 6.5]).
If the connected summing region is pinched, then
$\mu_{2}(T(\boldsymbol{y}_{1}),T(\boldsymbol{y}_{2}))=T(\mu_{2}(\boldsymbol{y}_{1},\boldsymbol{y}_{2})).$
We understand the chain complexes of stabilized diagrams. We will need a statement for free (link) stabilizations. Note that for free link stabilizations, since $z$ and $w$ are in the same elementary two-chain, if $\boldsymbol{CF}^{-}({\mathcal{}H}^{freelink})$ and $\boldsymbol{CF}^{-}({\mathcal{}H}^{free})$ are the chain complexes of free link stabilization and free stabilization, respectively, then we simply have $\boldsymbol{CF}^{-}({\mathcal{}H}^{freelink})=\boldsymbol{CF}^{-}({\mathcal{}H}^{free})\oplus V^{1/2}\boldsymbol{CF}^{-}({\mathcal{}H}^{free})$ as chain complexes.
Proposition 2.41 ([OS08a, Proposition 6.5], [Zem15, Proposition 6.5, Lemma 14.4]).
Let us consider a weakly admissible Heegaard datum ${\mathcal{}H}$ with Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ and coefficient ring $R$. Let ${\mathcal{}H}^{stab}$ be a free stabilization of ${\mathcal{}H}$ at the point $p\in\Sigma$. Then, if the connected summing region is pinched, then the differential for ${\mathcal{}H}^{stab}$ is of the form
$\begin{pmatrix}\partial_{{\mathcal{}H}}\otimes_{R}{\rm Id}_{R\llbracket V\rrbracket}&F\\ 0&\partial_{{\mathcal{}H}}\otimes_{R}{\rm Id}_{R\llbracket V\rrbracket}\end{pmatrix}$
with respect to the identification $\boldsymbol{CF}^{-}({\mathcal{}H}^{stab})\simeq\boldsymbol{CF}^{-}({\mathcal{}H})\llbracket V\rrbracket\oplus\boldsymbol{CF}^{-}({\mathcal{}H})\llbracket V\rrbracket$ where the two summands consist of the generators that have $\theta^{+}$, resp., $\theta^{-}$.
Furthermore, if $W$ is the weight of the connected component of $\Sigma\backslash\boldsymbol{\alpha}$ that contains the point $p$, then the map
$F:\boldsymbol{CF}^{-}({\mathcal{}H})\{\theta^{-}\}\llbracket V\rrbracket\to\boldsymbol{CF}^{-}({\mathcal{}H})\{\theta^{+}\}\llbracket V\rrbracket$
is homotopic to $V-W$ over $R\llbracket V\rrbracket$.
In particular, if we assume the condition described in Subsection 2.2, then $W=U_{i}$ for some $i$, and so the free-stabilization map $S:\boldsymbol{CF}^{-}({\mathcal{}H})\to\boldsymbol{CF}^{-}({\mathcal{}H}^{stab})$ is a quasi-isomorphism.
2.7. Links and Heegaard diagrams
We assume that the reader is familiar with the definition of link Floer homology [OS08a]. There are various additional data that we can equip to a link; we clarify our conventions.
Definition 2.42.
A minimally pointed link is a link $L\subset Y$ together with a set of (link) basepoints $\boldsymbol{u}\subset L$ such that for each connected component $L_{i}$ of $L$, we have $|L_{i}\cap\boldsymbol{u}|=2$.
A suture datum on a minimally pointed link $L\subset Y$ is a choice of one of the two connected components $L_{i}\backslash\boldsymbol{u}$ (which corresponds to the component inside the $\alpha$-handlebody) for each connected component $L_{i}$.
A $w|z$-partition of a minimally pointed link $(L,\boldsymbol{u})$ is a partition $\boldsymbol{u}=\boldsymbol{w}\sqcup\boldsymbol{z}$ of the link basepoints into $w$-basepoints $\boldsymbol{w}$ and $z$-basepoints $\boldsymbol{z}$ such that each link component has exactly one $w$-basepoint and one $z$-basepoint.
Let us consider the Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ of a Heegaard datum that we considered in Subsection 2.2. In this subsection and Subsubsection 2.8.1, we assume that there is no oriented arc $G$, for simplicity. Equivalently, let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ be a Heegaard diagram such that for each connected component ${\mathcal{}D}_{\boldsymbol{\alpha}}$ of $\Sigma\backslash\boldsymbol{\alpha}$, we have $|{\mathcal{}D}_{\boldsymbol{\alpha}}\cap\boldsymbol{p}|=1$ or $2$, and there exists a connected component ${\mathcal{}D}_{\boldsymbol{\beta}}$ of $\Sigma\backslash\boldsymbol{\beta}$ such that ${\mathcal{}D}_{\boldsymbol{\alpha}}\cap\boldsymbol{p}={\mathcal{}D}_{\boldsymbol{\beta}}\cap\boldsymbol{p}$.
This Heegaard diagram specifies a sutured manifold [Juh06, JTZ21] whose boundary components are either a sphere with one suture or a torus with two sutures. Equivalently, this specifies a minimally pointed link $(L,\boldsymbol{u})$ inside a pointed three-manifold $(Y,\boldsymbol{v})$ equipped with a suture datum. (We have $\boldsymbol{p}=\boldsymbol{v}\sqcup\boldsymbol{u}$).
Remark 2.43.
We use the convention that the orientation of the Heegaard surface is the same as the boundary of the $\alpha$-handlebody and is different from the boundary of the $\beta$-handlebody. Also see Remark 2.12. Note that [Man07] uses a different convention (see [BLS17, Section 4.1]); one should mirror the links to translate to our convention.
The sutured manifold corresponding to a suture datum is given as follows. The underlying three-manifold is $Y\backslash N(L\cup\boldsymbol{v})$ where $N$ means tubular neighborhood. The boundary component $\partial N(L_{i})$ has two meridional sutures on $\partial N(L_{i})$, one for each basepoint in $L_{i}\cap\boldsymbol{u}$; they are oriented in a way that $R_{-}$ [^28] Our convention is that $R_{-}$ corresponds to the $\alpha$-handlebody part. is the part that lies over the chosen component of $L_{i}\backslash\boldsymbol{u}$. Also, each connected component of $\partial N(\boldsymbol{v})$ has one suture.
Given a minimally pointed link $(L,\boldsymbol{u})$, there are three different kinds of choices we can make: a $w|z$-partition, an orientation on $L$, and a suture datum. Two of these determine the third: we use the convention that the link goes from $z$-basepoints to $w$-basepoints in the chosen components of $L\backslash\boldsymbol{u}$ (i.e. the $\alpha$-handlebody). In particular, given a suture datum, a $w|z$-partition is equivalent to an orientation on $L$. We say that the Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ together with such choice $\boldsymbol{z},\boldsymbol{w}\subset\boldsymbol{p}$ represents the oriented link $\overrightarrow{L}\subset Y$ (equipped with the $w|z$-partition, or a sutured datum).
2.8. ${\rm Spin}^{c}$-structures for unoriented links
In this subsection, we discuss ${\rm Spin}^{c}$-structures for unoriented links equipped with a suture datum [^29] It is in fact possible to define ${\rm Spin}^{c}$-structures without the choice of a suture datum. in a three-manifold, and discuss relative gradings of the corresponding Heegaard Floer chain complex.
A Heegaard diagram with more than two attaching curves defines a four-manifold $X$ with boundary, and in our case, the link basepoints define an embedded surface $S$. It should be possible to define the composition map $\mu_{n}^{\mathfrak{s}}$ corresponding to a relative ${\rm Spin}^{c}$-structure $\mathfrak{s}$ on $(X,S)$, such that $\mu_{n}=\sum_{\mathfrak{s}}\mu_{n}^{\mathfrak{s}}$. We will not discuss this in this paper in full generality. However, we will discuss two crude versions of this idea: one is in the form of a $\mathbb{Z}/2$-splitting, which we introduce in Subsubsection 2.9.2; the other case is when the Heegaard diagram is sufficiently simple, which we discuss in Subsection 2.10.
2.8.1. The case $G=\emptyset$
Let $L$ be an unoriented, minimally pointed link equipped with a suture datum (Definition 2.42) in a pointed three-manifold $(Y,\boldsymbol{v})$. Let us consider a Heegaard datum whose Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ represents $L$ in $(Y,\boldsymbol{v})$. In particular, assume that there is no oriented arc $G$. Let $Y_{L}^{sut}$ be the sutured manifold given by $L$.
There are two different spaces of relative ${\rm Spin}^{c}$-structures that we can consider. We can consider ${\rm Spin}^{c}(Y_{L}^{sut})$, the space of ${\rm Spin}^{c}$-structures on the sutured manifold $Y_{L}^{sut}$, as in [Juh06, Section 4], which comes with a map $\mathfrak{s}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to{\rm Spin}^{c}(Y_{L}^{sut})$ that specifies the ${\rm Spin}^{c}$-structure of intersection points. We can also consider $\underline{{\rm Spin}^{c}}(Y,L)$, the space of relative ${\rm Spin}^{c}$-structures on $Y\backslash N(L)$, as in [OS08a, Section 3], which comes with a map $c_{1}:\underline{{\rm Spin}^{c}}(Y,L)\to H^{2}(Y,L)$. Both spaces are an $H^{2}(Y,L)$-torsor.
Given an orientation $\mathfrak{o}$ on $L$, [OS08a, Section 3.6] gives an isomorphism $F_{\mathfrak{o}}:{\rm Spin}^{c}(Y_{L}^{sut})\to\underline{{\rm Spin}^{c}}(Y,L)$. Hence, we can also define a map $\mathfrak{s}_{\mathfrak{o}}:=F_{\mathfrak{o}}\circ\mathfrak{s}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to\underline{{\rm Spin}^{c}}(Y,L)$ [^30] In [OS08a], this is denoted as $\mathfrak{s}_{\boldsymbol{w},\boldsymbol{z}}$ where $\boldsymbol{w},\boldsymbol{z}$ are the $w$-, $z$-basepoints, respectively, which are determined by $\mathfrak{o}$ and the Heegaard diagram.. If $\mathfrak{o}^{\prime}$ is another orientation on $L$, then $\mathfrak{s}_{\mathfrak{o}^{\prime}}({\bf x})=\mathfrak{s}_{\mathfrak{o}}({\bf x})+PD(\mu)$ [OS08a, Lemma 3.12] for some $\mu\in M$ [^31] The homology class $\mu$ is the sum of the meridians of components of $L$ on which $\mathfrak{o}$ and $\mathfrak{o}^{\prime}$ are different; we do not need this fact., where $M$ is the $\mathbb{Z}$-linear subspace of $H_{1}(Y\backslash L)$ spanned by the meridians of components of $L$. Hence, given $\mathfrak{s}\in{\rm Spin}^{c}(Y_{L}^{sut})$, we can define its Chern class $c_{1}(\mathfrak{s})\in H^{2}(Y)$ as the image of $c_{1}(F_{\mathfrak{o}}(\mathfrak{s}))\in H^{2}(Y,L)$ under the map $H^{2}(Y,L)\to H^{2}(Y)$, which does not depend on the orientation $\mathfrak{o}$.
We do not need the entirety of ${\rm Spin}^{c}(Y_{L}^{sut})$: we consider a quotient. Let us identify $H^{2}(Y,L)\simeq H_{1}(Y\backslash L)$.
Definition 2.44.
Define the space of ${\rm Spin}^{c}$-structures as ${\rm Spin}^{c}(Y(L)):={\rm Spin}^{c}(Y_{L}^{sut})/2M$, which is an $H_{1}^{orb}(Y(L)):=H_{1}(Y\backslash L)/2M$-torsor. The above map $\mathfrak{s}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to{\rm Spin}^{c}(Y_{L}^{sut})$ descends to a map $\mathfrak{s}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to{\rm Spin}^{c}(Y(L))$.
For $\mathfrak{s}\in{\rm Spin}^{c}(Y(L))$, define its Chern class $c_{1}(\mathfrak{s})\in H^{2}(Y)$ [^32] We can define it as an element in $H_{1}^{orb}(Y(L))$, but this suffices since we only care about the pairing of $c_{1}(\mathfrak{s})$ with elements in $H_{2}(Y)$. as above. A ${\rm Spin}^{c}$-structure $\mathfrak{s}$ is torsion if $c_{1}(\mathfrak{s})$ is torsion.
Recall that the coefficient ring $R$ is of the form $\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{m}^{k_{m}}\rrbracket$ for $k_{i}\in\{1,1/2\}$. Recall that $k_{i}=1/2$ if and only if it corresponds to a link component $L_{i}$ of $L$; let $\mu_{i}\in H_{1}^{orb}(Y(L))$ be the meridian of $L_{i}$ in this case. If $k_{i}=1$, let $\mu_{i}=0$. Let
$\mathfrak{s}\left(U_{1}^{n_{1}}\cdots U_{m}^{n_{m}}{\bf x}\right):=\mathfrak{s}({\bf x})+\sum_{i}2n_{i}\mu_{i}\in{\rm Spin}^{c}(Y(L)).$
To talk about ${\rm Spin}^{c}(Y(L))$-summands, we should view $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$ as an $R^{\prime}:=\mathbb{F}\llbracket U_{1},\cdots,U_{m}\rrbracket$-module.
Definition 2.45.
Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ be a weakly admissible Heegaard diagram as above. An $R^{\prime}$-generator of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$ is an element of the form $U_{1}^{n_{1}}\cdots U_{m}^{n_{m}}{\bf x}$ where $n_{i}\in\{0,1/2\}$, and ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$. Note that these form an $R^{\prime}$-basis of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$.
If $\mathfrak{s}\in{\rm Spin}^{c}(Y(L))$, define $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta};\mathfrak{s})$ as the $R^{\prime}$-sub chain complex [^33] Indeed, this is a sub chain complex by [Juh06, Lemma 4.7]. of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})$ generated by the $R^{\prime}$-generators $\boldsymbol{y}$ such that $\mathfrak{s}(\boldsymbol{y})=\mathfrak{s}$.
Recall from [OS04c, Section 2.6], [Zem15, Section 4.2] that we can define $\mathfrak{s}_{\boldsymbol{r}}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to{\rm Spin}^{c}(Y)$ if $\boldsymbol{r}\subset\Sigma$ is a set of points such that each connected component of $\Sigma\backslash\boldsymbol{\alpha}$ and $\Sigma\backslash\boldsymbol{\beta}$ contains exactly one point in $\boldsymbol{r}$. The Chern class of $\mathfrak{s}({\bf x})$ can be computed in terms of the Chern class of $\mathfrak{s}_{\boldsymbol{r}}({\bf x})$: indeed, choose an orientation on $L$ and let $\overrightarrow{L}$ be the corresponding oriented link, and let $\boldsymbol{w},\boldsymbol{z}$ be the corresponding $w|z$-partition. Let $\boldsymbol{v}$ be the free basepoints. Then, [OS08a, Lemma 3.13] (also see [Zem19, Lemma 3.3]) implies
(2.3) $c_{1}(\mathfrak{s}({\bf x}))=c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\bf x}))-PD(\overrightarrow{L})=c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{z}}({\bf x}))+PD(\overrightarrow{L}).$
Lemma 2.46.
Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ be a Heegaard diagram as above, and let ${\mathcal{}D}\in D({\bf x},{\bf x})$ be a cornerless domain. Then, we have
$\mu({\mathcal{}D})=\left\langle c_{1}(\mathfrak{s}({\bf x})),H({\mathcal{}D})\right\rangle+P({\mathcal{}D}),$
where $P({\mathcal{}D})$ is the total multiplicity of ${\mathcal{}D}$ (Definition 2.14).
Proof.
This is a direct consequence of the Maslov index formula for cornerless domains in Heegaard diagrams for three-manifolds ([OS04b, Proposition 7.5], [Lip06, Corollary 4.12], [Zem15, Equation 4.9]) and Equation 2.3. (Compare [Fan17, Section 5.2].) ∎
Corollary 2.47.
Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ be a weakly admissible Heegaard diagram as above, and let $\mathfrak{s}\in{\rm Spin}^{c}(Y(L))$. Then, $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta};\mathfrak{s})$ has a relative $\mathbb{Z}/\mathfrak{d}(c_{1}(\mathfrak{s}))$-grading, where $\mathfrak{d}(c_{1}(\mathfrak{s}))$ is the divisibility of $c_{1}(\mathfrak{s})$.
Proof.
Immediate from Lemma 2.46. ∎
Example 2.48.
Consider the weakly admissible Heegaard diagram $(\mathbb{T}^{2},\alpha,\beta,\{w,z\})$ given by the left side of Figure 2.6. Work over $\mathbb{F}\llbracket U^{1/2}\rrbracket$, and assign the weight $U^{1/2}$ to both $w$ and $z$. The group $H_{1}^{orb}(Y(L))\simeq\mathbb{Z}\oplus\mathbb{Z}/2$; the meridian $\mu$ represents $(0,1)$. The ${\rm Spin}^{c}$-structures $\mathfrak{s}(\sigma)$ and $\mathfrak{s}(\tau)$ are both torsion, and they differ by $\mu$. We have
$\boldsymbol{CF}^{-}(\alpha,\beta;\mathfrak{s}(\sigma))=\sigma\mathbb{F}\llbracket U\rrbracket\oplus U^{1/2}\tau\mathbb{F}\llbracket U\rrbracket,\ \boldsymbol{CF}^{-}(\alpha,\beta;\mathfrak{s}(\tau))=\tau\mathbb{F}\llbracket U\rrbracket\oplus U^{1/2}\sigma\mathbb{F}\llbracket U\rrbracket.$
The differential is identically 00.
Now, consider the right side of Figure 2.6. The group $H_{1}^{orb}(Y(L))\simeq H_{1}(Y)\simeq\mathbb{Z}$, and $\mathfrak{s}(\sigma)=\mathfrak{s}(\tau)$. Their Chern class $c_{1}(\mathfrak{s}(\sigma))$ is a generator of $H_{1}(Y)$; in particular, it has divisibility $1$. Indeed, there are two domains in $D(\tau,\sigma)$ with weights $1,U^{1/2}$, respectively.
Figure 2.6. Some simple doubly pointed Heegaard diagrams
2.8.2. The case $G\neq\emptyset$
Let us consider the case where $G$ may exist. In this case, we can formally work with the Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p}\sqcup\{\partial_{+}G\})$; define the corresponding $L\subset Y$ and $\mathfrak{s}:\boldsymbol{\alpha}\cap\boldsymbol{\beta}\to{\rm Spin}^{c}(Y(L))$ accordingly.
If at most one of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ intersects $G$, then Remark 2.16 justifies this. However, if both $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ intersect $G$, then it turns out that the link $L$ is not the link we want: consider the projection map $\pi:N(\Sigma)\to\Sigma$, and assume that $L$ is vertical in $N(\Sigma)$, i.e. that $L\cap N(\Sigma)$ is the inverse image of the link basepoints on $\Sigma$. Then $\pi^{-1}(G)$ is a split band on $L$; the link we should consider is the link obtained by surgering $L$ along $\pi^{-1}(G)$. We will not try to fix this; for our purposes, it is sufficient to consider Chern class summands as follows.
Definition 2.49.
Let us consider a weakly admissible Heegaard diagram as above that may include an oriented arc $G$. Let $\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}}$, $\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}}$ be the attaching curves with local systems, and let $c\in H^{2}(Y)$. Define $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}};c)$ as the $R$-sub chain complex of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ generated by the intersection points ${\bf x}$ such that $c_{1}(\mathfrak{s}({\bf x}))=c$.
2.8.3. Strong $\mathfrak{s}$-admissibility and $CF^{-}$
Finally, let us comment on how to define $CF^{-}$.
Definition 2.50.
The Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ is $\mathfrak{s}$-strongly admissible if all cornerless two-chains ${\mathcal{}D}$ such that
$\left\langle c_{1}(\mathfrak{s}),H({\mathcal{}D})\right\rangle+P({\mathcal{}D})=0$
have both positive and negative local multiplicities, where $H({\mathcal{}D})\in H_{2}(Y)$ is the homology class that ${\mathcal{}D}$ represents.
Note that a weakly admissible Heegaard diagram is $\mathfrak{s}$-strongly admissible for all torsion $\mathfrak{s}$.
If $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{p})$ is $\mathfrak{s}$-strongly admissible, then we can work over the polynomial ring $\mathbb{F}[U_{1}^{k_{1}},\cdots,U_{m}^{k_{m}}]$ instead and define $CF^{-}(\boldsymbol{\alpha},\boldsymbol{\beta};\mathfrak{s})$. If $G\neq\emptyset$, similarly define $CF^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}};c)$.
Remark 2.51.
Equation 2.3 says that our strong admissibility condition considers the average of the ${\rm Spin}^{c}$-structures obtained by the different choices of the $z$-basepoints. This has the advantage that the left hand side of Figure 2.6 is strongly admissible with respect to both torsion ${\rm Spin}^{c}$-structures. In contrast, it is not even weakly admissible with respect to $z$ (or $w$).
2.9. Gradings
Gradings in $A_{\infty}$-categories, if they exist, will turn out to be very useful. For instance, we will use gradings to show that some generators do not arise as summands of certain composition maps.
2.9.1. A homological $\mathbb{Z}$-grading
Sometimes, the $A_{\infty}$-categories we consider can be given a (homological) $\mathbb{Z}$-grading, in the sense of Subsection 2.4. Compare [MO08, Section 3], [BL12, Section 5.2]. Note that our notions differ: their periodic domains are cornerless two-chains in this paper.
Definition 2.52.
A Heegaard datum is (homologically) $\mathbb{Z}$-gradable if whenever ${\mathcal{}D}$ and ${\mathcal{}D}^{\prime}$ are two domains with the same vertices, then
(2.4) $P({\mathcal{}D})-\mu({\mathcal{}D})=P({\mathcal{}D}^{\prime})-\mu({\mathcal{}D}^{\prime}).$
Lemma 2.18 implies the following proposition.
Proposition 2.53.
If a weakly admissible Heegaard datum is homologically $\mathbb{Z}$-gradable, then the associated $A_{\infty}$-category ${\mathcal{}A}$ in the sense of Remark 2.25 can be $\mathbb{Z}$-graded, such that:
  • each $R$-generator $f{\bf x}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is homogeneous;
  • $U_{i}^{k}$ has grading $-2k$, i.e. if $\boldsymbol{y}$ is homogeneous, then ${\rm gr}_{{\mathcal{}A}}(U_{i}^{k}\boldsymbol{y})={\rm gr}_{{\mathcal{}A}}(\boldsymbol{y})-2k$ for $k\in\mathbb{Z}$ or $1/2\mathbb{Z}$;
  • if $f{\bf x}$ and $g{\bf x}$ are generators, then ${\rm gr}_{{\mathcal{}A}}(f{\bf x})-{\rm gr}_{{\mathcal{}A}}(g{\bf x})=-2({\rm gr}_{U}(f)-{\rm gr}_{U}(g))$.
Similarly, we say that a Heegaard datum if it is homologically $\mathbb{Z}/n$-gradable if Equation 2.4 holds modulo $n$. If it is homologically $\mathbb{Z}/n$-gradable, then ${\mathcal{}A}$ can be $\mathbb{Z}/n$-graded.
Note that if the attaching curves are $\boldsymbol{\alpha}_{0}^{E_{0}},\cdots,\boldsymbol{\alpha}_{m}^{E_{m}}$, then we are free to shift the gradings of each $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i}^{E_{i}},\boldsymbol{\alpha}_{i+1}^{E_{i+1}})$ by a constant $C_{i}$.
Definition 2.54.
We use the convention that if $\boldsymbol{\alpha}_{j}$ is a small, standard translate of $\boldsymbol{\alpha}_{i}$, then $\Theta^{+}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i}^{E_{i}},\boldsymbol{\alpha}_{j}^{E_{j}})$ has grading 00343434If the local systems are trivial, then $\Theta^{+}$ can be viewed as an element of $\boldsymbol{CF}^{-}(\#^{N}S^{1}\times S^{2})$. Our convention is different from the absolute $\mathbb{Q}$-grading of $\Theta^{+}$, which is $N/2$..
2.9.2. An Alexander $\mathbb{Z}/2$-splitting
The Heegaard data that we consider will often not be $\mathbb{Z}$-gradable. However, the corresponding $A_{\infty}$-category can be given a nice $\mathbb{Z}/2$-splitting in the sense of Propositon 2.57. For links in $S^{3}$, this is the relative, collapsed Alexander $\mathbb{Z}/2$-grading, and in general, it is closely related to the ${\rm Spin}^{c}$-splitting discussed in Subsection 2.8. Note that now, we have to deal with the case where $G$ is present.
Definition 2.55.
A Heegaard datum is (Alexander) $\mathbb{Z}/2$-splittable if whenever ${\mathcal{}D}$ and ${\mathcal{}D}^{\prime}$ are two domains with the same vertices, we have $P({\mathcal{}D})=P({\mathcal{}D}^{\prime})$ modulo $2$.
To talk about the corresponding splitting in the $A_{\infty}$-category, we consider $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ not as an $R$-module, but as an $R^{\prime}:=\mathbb{F}\llbracket U_{1},\cdots,U_{m}\rrbracket$-module, as in Subsection 2.8. If $G$ exists, then one of the $U_{i}$’s is the distinguished variable $U$.
Definition 2.56.
An $R^{\prime}$-generator of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$ is of the form $fU_{1}^{n_{1}}\cdots U_{k}^{n_{k}}{\bf x}$ where $f$ is an $R$-basis element of ${\rm Hom}_{R}(E_{\boldsymbol{\alpha}},E_{\boldsymbol{\beta}})$, $n_{i}\in\{0,1/2\}$, and ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$. We write $U^{{\bf n}}=U_{1}^{n_{1}}\cdots U_{k}^{n_{k}}$ for ${\bf n}=(n_{1},\cdots,n_{k})$. Note that these form an $R^{\prime}$-basis of $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})$.
Let $f_{i}U^{{\bf n}_{i}}{\bf x}_{i}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i-1}^{E_{i-1}},\boldsymbol{\alpha}_{i}^{E_{i}})$ for $i=1,\cdots,d$ and $\boldsymbol{y}_{d+1}=f_{d+1}U^{{\bf n}_{d+1}}{\bf x}_{d+1}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{0}^{E_{0}},\boldsymbol{\alpha}_{d}^{E_{d}})$ be $R^{\prime}$-generators. Let $D(f_{1}U^{{\bf n}_{1}}{\bf x}_{1},\cdots,f_{d+1}U^{{\bf n}_{d+1}}{\bf x}_{d+1})$ be the set of domains ${\mathcal{}D}\in D({\bf x}_{1},\cdots,{\bf x}_{d+1})$ such that $w({\mathcal{}D})\rho({\mathcal{}D})(f_{1}U^{{\bf n}_{1}}\otimes\cdots\otimes f_{d}U^{{\bf n}_{d}})=rf_{d+1}U^{{\bf n}_{d+1}}$ for some nonzero $r\in R^{\prime}$.
A similar idea as above gives the following proposition.
Proposition 2.57.
If a weakly admissible Heegaard datum is $\mathbb{Z}/2$-splittable, then the associated $A_{\infty}$-category ${\mathcal{}A}$ has a $\mathbb{Z}/2$-splitting in the following sense:
  • The chain complexes $\boldsymbol{CF}^{-}(\boldsymbol{\alpha}^{E_{\boldsymbol{\alpha}}},\boldsymbol{\beta}^{E_{\boldsymbol{\beta}}})=C_{0}\oplus C_{1}$ are $\mathbb{Z}/2$-graded, and satisfies the three conditions in Proposition 2.53. In particular, the summands $C_{0}$ and $C_{1}$ are $R^{\prime}$-modules. We denote this grading as ${\rm gr}_{A}^{\mathbb{Z}/2}(\boldsymbol{y})$ for a homogeneous element $\boldsymbol{y}$, and call it the Alexander $\mathbb{Z}/2$-grading.
  • Let $\boldsymbol{y}_{i}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i-1}^{E_{i-1}},\boldsymbol{\alpha}_{i}^{E_{i}})$ for $i=1,\cdots,d$ and $\boldsymbol{y}_{d+1}\in\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{0}^{E_{0}},\boldsymbol{\alpha}_{d}^{E_{d}})$ be $R^{\prime}$-generators. If there exists a domain ${\mathcal{}D}\in D(\boldsymbol{y}_{1},\cdots,\boldsymbol{y}_{d+1})$, then
    $\sum_{i=1}^{d+1}{\rm gr}_{A}^{\mathbb{Z}/2}(\boldsymbol{y}_{i})=0.$
    In particular, the composition maps $\mu_{d}$ have degree 00.
As before, we use the convention that $\Theta^{+}$ has grading 00.
Example 2.58.
Recall Example 2.48. The left hand side of Figure 2.6 is $\mathbb{Z}/2$-splittable, and the corresponding $\mathbb{Z}/2$-splitting on $\boldsymbol{CF}^{-}(\alpha,\beta)$ is $\boldsymbol{CF}^{-}(\alpha,\beta;\mathfrak{s}(\sigma))\oplus\boldsymbol{CF}^{-}(\alpha,\beta;\mathfrak{s}(\tau))$.
Remark 2.59.
Note that this $\mathbb{Z}/2$-splitting does not see the Maslov index of domains: it only sees whether domains exist. The second condition of Proposition 2.53 implies that whenever we have a $\mathbb{Z}/2$-splitting, then the $A_{\infty}$-functor ${\mathcal{}F}$ induced by changing the almost complex structure preserves the splitting, i.e. ${\mathcal{}F}_{d}$ has degree 00 for all $d$.
Remark 2.60.
We can consider twisted complexes in this setting as well, in which case we add the condition that the differential $\delta$ is homogeneous of degree 00. The composition maps $\mu_{d}^{{\rm Tw}}$ and the functor induced by changing the almost complex structure are homogeneous of degree 00.
2.10. Homologous attaching curves
Definition 2.61.
Two attaching curves $\boldsymbol{\alpha},\boldsymbol{\beta}$ Heegaard surface $\Sigma$ are homologous if they span the same subspace of $H_{1}(\Sigma)$.
Note that whether two curves are homologous do not depend on any basepoints on $\Sigma$.
In this subsection, let us consider a Heegaard diagram $(\Sigma,\boldsymbol{\alpha}_{0},\cdots,\boldsymbol{\alpha}_{c},\boldsymbol{\beta}_{0},\cdots,\boldsymbol{\beta}_{d},\boldsymbol{p})$ (and possibly also $G$) such that the $\boldsymbol{\alpha}_{i}$’s are pairwise homologous and also the $\boldsymbol{\beta}_{i}$’s are pairwise homologous [^35] If all the attaching curves are homologous, then we also are implicitly making a choice of $c$; i.e. we choose which ones are the $\boldsymbol{\alpha}_{i}$’s and which are the $\boldsymbol{\beta}_{j}$’s..
Under this assumption, it is easier to check whether the Heegaard datum is $\mathbb{Z}$-gradable or $\mathbb{Z}/2$-splittable. Similar statements to the results in this subsection hold when the attaching curves are cyclically permuted.
Lemma 2.62.
If ${\mathcal{}D},{\mathcal{}D}^{\prime}\in D({\bf x}_{0},\cdots,{\bf x}_{c+d+1})$ are domains with the same vertices, then there exist cornerless domains ${\mathcal{}P}_{j}\in D({\bf x}_{j},{\bf x}_{j})$ for $j\in J$ where $J=\{0,\cdots,c+d+1\}\backslash\{c+d+1\}$ or $\{0,\cdots,c+d+1\}\backslash\{c\}$ [^36] In fact, this lemma holds for $J=\{0,\cdots,c+d+1\}\backslash\{k\}$ for any $k$., such that
${\mathcal{}D}={\mathcal{}D}^{\prime}+\sum_{j\in J}{\mathcal{}P}_{j}.$
Proof.
Let us consider the case where $J=\{0,\cdots,c+d+1\}\backslash\{c+d+1\}$; the other case follows similarly. Let ${\mathcal{}P}$ be the difference ${\mathcal{}D}-{\mathcal{}D}^{\prime}$ of the underlying two-chains. Recursively define ${\mathcal{}P}_{c+j}$ for $j=d,\cdots,1$ such that the $\boldsymbol{\beta}_{i}$-boundary of the two-chain ${\mathcal{}P}-\sum_{l=c+j}^{c+d}{\mathcal{}P}_{l}$ is empty for $i\geq j$. Similarly define ${\mathcal{}P}_{j}$ for $j=0,\cdots,c-1$ such that (all the $\boldsymbol{\beta}_{k}$-boundaries and) the $\boldsymbol{\alpha}_{i}$-boundary of the two-chain ${\mathcal{}P}-\sum_{l=c+1}^{c+d}{\mathcal{}P}_{l}-\sum_{l=0}^{j}{\mathcal{}P}_{l}$ is empty for $i\leq j$. Let ${\mathcal{}P}_{c}$ be the $\boldsymbol{\alpha}_{c}\boldsymbol{\beta}_{0}$-cornerless domain with underlying two-chain ${\mathcal{}P}-\sum_{j\in J\backslash\{c\}}{\mathcal{}P}_{j}$. ∎
Corollary 2.63.
If all cornerless domains [^37] In particular, they have two vertices. ${\mathcal{}D}$ satisfy $P({\mathcal{}D})=\mu({\mathcal{}D})$, then the Heegaard datum is $\mathbb{Z}$-gradable. If they satisfy $P({\mathcal{}D})=0$ modulo $2$ instead, then the Heegaard datum is $\mathbb{Z}/2$-splittable.
Proof.
This follows from Lemma 2.62 and the additivity of $P$ and $\mu$. ∎
Lemma 2.62 is equivalent to that a cohomology class $x\in H^{2}(X)$ on the corresponding four-manifold $X=X_{\boldsymbol{\alpha}_{0}\cdots\boldsymbol{\alpha}_{c}\boldsymbol{\beta}_{0}\cdots\boldsymbol{\beta}_{d}}$ [OS04c, Subsection 8.1] is determined by its restriction to the boundary (minus one boundary component).
Lemma 2.64.
Let ${\mathcal{}D}\in D({\bf x}_{0},\cdots,{\bf x}_{c+d+1})$ be a domain. If the ${\rm Spin}^{c}$-structures $\mathfrak{s}({\bf x}_{j})$ are torsion for $j\in\{0,\cdots,c+d+1\}\backslash\{c,c+d+1\}$, then $c_{1}(\mathfrak{s}({\bf x}_{c}))=c_{1}(\mathfrak{s}({\bf x}_{c+d+1}))$.
Proof.
Let $\boldsymbol{v}$ be the set of free basepoints. If $G$ exists, for the purpose of this Lemma, consider $\partial_{+}G,\partial_{-}G$ as link basepoints and replace $\boldsymbol{p}$ with $\boldsymbol{p}\sqcup\{\partial_{+}G\}$, and let us choose which link basepoints are $z$-basepoints; write $\boldsymbol{p}=\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}$.
We can define the $\boldsymbol{v}\sqcup\boldsymbol{w}$-${\rm Spin}^{c}$ structure of ${\mathcal{}D}$ and its Chern class $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\mathcal{}D}))\in H^{2}(X)$. The key property we use is that $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\mathcal{}D}))$ restricts to $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\bf x}_{i}))\in H^{2}(Y_{\boldsymbol{\delta\epsilon}})$ where ${\bf x}_{i}\in\boldsymbol{\delta}\cap\boldsymbol{\epsilon}$. Let $\overrightarrow{L_{\boldsymbol{\delta\epsilon}}}\subset Y_{\boldsymbol{\delta\epsilon}}$ be the corresponding oriented link; then we have $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\bf x}_{i}))=c_{1}(\mathfrak{s}({\bf x}_{i}))+PD(\overrightarrow{L_{\boldsymbol{\delta\epsilon}}})$ by Equation 2.3.
Recall the identification
$H^{2}(X)\simeq{\rm coker}\left(H_{1}(\Sigma)\to\bigoplus_{\boldsymbol{\gamma}}H_{1}(U_{\boldsymbol{\gamma}})\right),\ H^{2}(Y_{\boldsymbol{\delta\epsilon}})\simeq{\rm coker}\left(H_{1}(\Sigma)\to H_{1}(U_{\boldsymbol{\delta}})\oplus H_{1}(U_{\boldsymbol{\epsilon}})\right),$
where $U_{\boldsymbol{\gamma}}$ be the $\boldsymbol{\gamma}$-handlebody. For each pair of link basepoints, choose an oriented arc on $\Sigma$ from the $w$-basepoint to the $z$-basepoint, and let $\overrightarrow{L_{\boldsymbol{\gamma}}}\in H_{1}(U_{\boldsymbol{\gamma}})$ be obtained by adding the above oriented arc to an oriented arc on $\Sigma\backslash\boldsymbol{\gamma}$ from the $z$-basepoint to the $w$-basepoint. Then the oriented link $\overrightarrow{L_{\boldsymbol{\delta\epsilon}}}\in H_{1}(Y_{\boldsymbol{\delta\epsilon}})\simeq H^{2}(Y_{\boldsymbol{\delta\epsilon}})$ is represented by the class $(\overrightarrow{L_{\boldsymbol{\delta}}},\overrightarrow{L_{\boldsymbol{\epsilon}}})$ under the above identification. Using the above identification together with that the restriction map $H^{2}(X)\to H^{2}(Y_{\boldsymbol{\delta\epsilon}})$ is induced by the projection map $\bigoplus_{\boldsymbol{\gamma}}H_{1}(U_{\boldsymbol{\gamma}})\to H_{1}(U_{\boldsymbol{\delta}})\oplus H_{1}(U_{\boldsymbol{\epsilon}})$, one can write $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\bf x}_{c+d+1}))$ in terms of $c_{1}(\mathfrak{s}_{\boldsymbol{v}\sqcup\boldsymbol{w}}({\bf x}_{c}))$ and the $\overrightarrow{L_{\boldsymbol{\gamma}}}$’s. Combining everything, we get $c_{1}(\mathfrak{s}({\bf x}_{c}))=c_{1}(\mathfrak{s}({\bf x}_{c+d+1}))$. ∎
The above proof also shows that if the torsion hypothesis is removed, then there is a relation between the $c_{1}(\mathfrak{s}({\bf x}_{j}))$’s.
Proposition 2.65.
Let us consider a weakly admissible Heegaard datum whose underlying Heegaard diagram is as in the beginning of this subsection. Let $Y$ be the three-manifold obtained by attaching curves $\boldsymbol{\alpha}_{0}$ and $\boldsymbol{\beta}_{0}$ [^38] The three-manifold $Y_{\boldsymbol{\alpha}_{i}\boldsymbol{\beta}_{j}}$ does not depend on $i,j$., and choose $c\in H^{2}(Y)$. Consider the $A_{\infty}$-category ${\mathcal{}A}$ whose objects are the attaching curves (equipped with local systems), and
${\rm Hom}_{{\mathcal{}A}}(\boldsymbol{\alpha}_{a}^{E_{\boldsymbol{\alpha}_{a}}},\boldsymbol{\beta}_{b}^{E_{\boldsymbol{\beta}_{b}}})=\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{a}^{E_{\boldsymbol{\alpha}_{a}}},\boldsymbol{\beta}_{b}^{E_{\boldsymbol{\beta}_{b}}};c)$
${\rm Hom}_{{\mathcal{}A}}(\boldsymbol{\alpha}_{i}^{E_{\boldsymbol{\alpha}_{i}}},\boldsymbol{\alpha}_{j}^{E_{\boldsymbol{\alpha}_{j}}})=\boldsymbol{CF}^{-}(\boldsymbol{\alpha}_{i}^{E_{\boldsymbol{\alpha}_{i}}},\boldsymbol{\alpha}_{j}^{E_{\boldsymbol{\alpha}_{j}}};0),\ {\rm Hom}_{{\mathcal{}A}}(\boldsymbol{\beta}_{i}^{E_{\boldsymbol{\beta}_{i}}},\boldsymbol{\beta}_{j}^{E_{\boldsymbol{\beta}_{j}}})=\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{i}^{E_{\boldsymbol{\beta}_{i}}},\boldsymbol{\beta}_{j}^{E_{\boldsymbol{\beta}_{j}}};0),$
for all $a,b$ and $i<j$; if $i\geq j$, then define ${\rm Hom}_{{\mathcal{}A}}(\boldsymbol{\alpha}_{i}^{E_{\boldsymbol{\alpha}_{i}}},\boldsymbol{\alpha}_{j}^{E_{\boldsymbol{\alpha}_{j}}})={\rm Hom}_{{\mathcal{}A}}(\boldsymbol{\beta}_{i}^{E_{\boldsymbol{\beta}_{i}}},\boldsymbol{\beta}_{j}^{E_{\boldsymbol{\beta}_{j}}})=0$.
Then, ${\mathcal{}A}$ is homologically $\mathbb{Z}/\mathfrak{d}(c)$-gradable.
Proof.
The composition maps are well-defined and ${\mathcal{}A}$ is an $A_{\infty}$-category by Lemma 2.64, and it is homologically $\mathbb{Z}/\mathfrak{d}(c)$-gradable by Lemmas 2.46 and 2.62. ∎
2.10.1. Strong $c$-admissibility
We can define strong $c$-admissibility in the context of this subsection. Compare [OS04c, Definition 8.8 and Section 8.4.2] and [Zem15, Section 4.8].
Given a cornerless two-chain ${\mathcal{}D}$, by identifying the $U_{\boldsymbol{\alpha}_{i}}$’s together and the $U_{\boldsymbol{\beta}_{j}}$’s together, we can define the homology class $H({\mathcal{}D})$ of ${\mathcal{}D}$ as an element in $H_{2}(Y)$.
Definition 2.66.
Let the Heegaard diagram and the three-manifold $Y$ be as in Proposition 2.65. The Heegaard diagram is $c$-strongly admissible for $c\in H^{2}(Y)$ if all cornerless two-chains ${\mathcal{}D}$ such that
$\left\langle c,H({\mathcal{}D})\right\rangle+P({\mathcal{}D})=0$
have both positive and negative local multiplicities.
If ${\mathcal{}D},{\mathcal{}D}^{\prime}$ are two domains with the same vertices; if there exist some $i,j$ for which ${\mathcal{}D},{\mathcal{}D}^{\prime}$ have an $\boldsymbol{\alpha}_{i}\boldsymbol{\beta}_{j}$-vertex ${\bf x}$, assume $c_{1}(\mathfrak{s}({\bf x}))=c$ [^39] The Chern class does not depend on $i,j,{\bf x}$ by Lemma 2.64.. Consider the cornerless two-chain ${\mathcal{}D}-{\mathcal{}D}^{\prime}$: Lemmas 2.62 and 2.46, together with the additivity of $\mu$ and $P$, imply that
$\mu({\mathcal{}D})-\mu({\mathcal{}D}^{\prime})=\left\langle c,H({\mathcal{}D}-{\mathcal{}D}^{\prime})\right\rangle+P({\mathcal{}D})-P({\mathcal{}D}^{\prime}).$
Hence, Proposition 2.65 works for $CF^{-}$ as well, if we further assume that the Heegaard diagram is $c$-strongly admissible.
3. Band maps in unoriented link Floer homology
The goal of this section is to precisely define the objects we consider, to define the band maps, and to show that they are well-defined. Compare [Zem19, Section 6] and [Fan17]. Note that the objects we consider are different, and hence the band maps are different as well. The band maps we consider are different from [Fan17] even for non-orientable bands: Fan chooses “the other generator,” i.e. the generator that corresponds to $\sigma$ in the left hand side of Figure 3.3.
We assume that the reader is familiar with the definition of link Floer homology [OS08a].
3.1. Bands and balled links
In this subsection, we introduce the notion of a balled link and define its unoriented link Floer homology. We first explain the motivation behind considering balled links. Recall from Definition 1.5 that we needed exactly two link basepoints in each link component to define unoriented link Floer homology. Hence, the merge and split band maps involve links with different number of basepoints. As in the discussion of Subsection 1.3, to go around this issue, we free-stabilize our ambient manifold when we consider the link with fewer basepoints.
Figure 3.1 is a diagram of a kind of split band that we consider. The link $L_{a}$ has two link basepoints $z$ and $w$, but there is also a free basepoint $v$ in the underlying three-manifold. Band surgery along the specified band creates a new link component (the left hand side part of $L_{b}$). We added the free basepoint $v$ so that after we do this band surgery, we can make the basepoint $v$ into two link basepoints $v,v^{\prime}$ of the left hand side component (compare the Heegaard diagram). However, to do this, we have to isotope the left hand side component so that it goes through $v$ (and $v^{\prime}$). Hence, the link $L_{a}$, the basepoints $v,w,z$, and the band do not uniquely determine $L_{b}$, let alone the isotopy.
We introduce baseballs and consider balled links to specify $L_{b}$ and the isotopy up to choices that do not affect the unoriented link Floer homology and maps between them: we specify a baseball $B_{v}\ni v$ that intersects $L_{a}$ in an interval, and isotope the link inside the baseball $B_{v}$. (We will also let $v^{\prime}\in B_{v}$.)
Considering baseballs and balled links also has the advantage that the following can be encoded in a nice way: in this paper, we only consider $\beta$-bands in the sense of [Zem19, Section 6.1], which in effect means that the two link basepoints in each link component are “very close to each other” with respect to the bands, and that we treat the two basepoints as one basepoint. This is encoded as that each link has one link baseball (one should think of baseballs as being very small and very round), which contains these two basepoints. In other words, the link baseballs mark the $\alpha$-handlebody part of the link.
Figure 3.1. A split band and a Heegaard diagram for it
Definition 3.1.
A balled link $L$ in a three-manifold $Y$ is an (possibly empty) unoriented link $L\subset Y$ together with a finite, ordered sequence of baseballs $B_{b}$, such that they are pairwise disjoint and for each baseball $B_{b}$, it is either disjoint from $L$, or there exists a small neighborhood $U\supset B_{b}$ such that $(Y\cap U,L\cap U,B_{b})$ is diffeomorphic to $(\mathbb{R}^{3},\mathbb{R}\times\{(0,0)\},\{{\bf x}:\left|{\bf x}\right|\leq 1\})$. It is possible that the underlying link is empty, but we require the balled link $L$ to have at least one baseball.
Each baseball is labelled as a link baseball or a free baseball, and each baseball is either marked or unmarked. For each link component, there is exactly one link baseball that it intersects, and every link baseball has to intersect $L$. It is possible that none of the baseballs are marked.
To define the unoriented link Floer chain complex of a balled link $L\subset Y$, we define what it means for a Heegaard datum to represent the balled link $L\subset Y$. To define this, it is convenient to first define an auxiliary datum for $L\subset Y$.
Definition 3.2.
Let $L\subset Y$ be a balled link with baseballs $B_{1},\cdots,B_{n}$. An auxiliary datum for $L\subset Y$ is a tuple $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ of an oriented link $\overrightarrow{L}\subset Y$, link basepoints $\boldsymbol{w}\sqcup\boldsymbol{z}$, and free basepoints $\boldsymbol{v}$ that satisfies the following.
  • The sets $\boldsymbol{w},\boldsymbol{z}$ consist of (link) basepoints $w_{i},z_{i}\in L\cap P_{i}$, respectively, for each link baseball $P_{i}\in\{B_{1},\cdots,B_{n}\}$.
  • The set $\boldsymbol{v}$ consists of basepoints $v_{j}$ in the interior of $Q_{j}$ for each free baseball $Q_{j}\in\{B_{1},\cdots,B_{n}\}$.
  • The link $\overrightarrow{L}$ (ignoring its orientation for now) is obtained from $L$ by isotoping $L$ in the interior of each free baseball $Q_{j}$ so that $\overrightarrow{L}$ does not intersect the free basepoints $v_{j}$.
  • The components of $\overrightarrow{L}\backslash(\boldsymbol{w}\sqcup\boldsymbol{z})$ that are contained in the link baseballs go from a $z$-basepoint to a $w$-basepoint. (Recall Subsection 2.7.)
Definition 3.3.
Let $L\subset Y$ be a balled link, and let $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ be an auxiliary datum for it. A Heegaard datum represents the auxiliary datum $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ if it satisfies the following.
  • The Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ represents the pointed link $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z})$ inside the pointed manifold $(Y,\boldsymbol{v})$.
  • The coefficient ring is $R=\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{n}^{k_{n}}\rrbracket$, where the $k_{l}$’s and the weight function are as follows: (recall that the baseballs are ordered) if the $l$th baseball $B_{l}$ is a link baseball, then $k_{l}=1/2$ and we assign the weight $U_{l}^{1/2}$ to both link basepoints corresponding to $B_{l}$. If the $l$th baseball $B_{l}$ is a free baseball, then $k_{l}=1$ and we assign the weight $U_{l}$ to the free basepoint corresponding to $B_{l}$.
A Heegaard datum represents $L\subset Y$ if there exists an auxiliary datum for $L\subset Y$ that the Heegaard datum represents.
We define the unoriented link Floer chain complex of a balled link, and hence the unoriented link Floer homology of a balled link. We will show in Proposition 3.7 that the chain complex is well defined up to chain homotopy equivalence, and also show naturality, i.e. that we can choose the homotopy equivalence up to homotopy.
Definition 3.4.
Let $L\subset Y$ be a balled link, and let us choose a weakly admissible Heegaard datum ${\mathcal{}H}$ that represents $L\subset Y$. The unoriented link Floer chain complex $\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)$ of $L\subset Y$ is $\boldsymbol{CF}^{-}({\mathcal{}H})$. Define the infinity version as
$\boldsymbol{CFL}^{\prime}{}^{\infty}(Y,L)=\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)\otimes_{R}R^{\infty}=\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)\otimes_{R}R[U_{1}^{-1},\cdots,U_{n}^{-1}].$
If the marked link baseballs are $B_{i_{1}},\cdots,B_{i_{n}}$ and the marked free baseballs are $B_{j_{1}},\cdots,B_{j_{m}}$ for $n+m\geq 1$, then define the hat version as
$\widehat{CFL^{\prime}}(Y,L)=\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)/(U_{i_{1}}^{1/2},\cdots,U_{i_{n}}^{1/2},U_{j_{1}},\cdots,U_{j_{m}}).$
If none of the baseballs of $L$ are marked, then we do not define the hat version; instead, we consider a new balled link $L^{\sharp}$ given by $L$ together with a marked free baseball (or equivalently, together with a unlinked, marked unknot), and consider $\widehat{CFL^{\prime}}(Y,L^{\sharp})$ instead. Note that $\widehat{HFL^{\prime}}(Y,L^{\sharp})$ is isomorphic to the homology of $\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)/U_{1}$.
Remark 3.5.
We require at least one of the baseballs to be marked to talk about the hat version, although we could have defined it as $\boldsymbol{CFL}^{\prime}{}^{-}(Y,L)/U_{1}$. This is to explicitly keep track of the marked point that corresponds to the variable we quotient out by. Our definitions follow the definitions of $I^{\sharp}$ and $I^{\natural}$ (also compare with unreduced/reduced Khovanov homology), although our definitions are slightly more general as we allow more than one link component to be marked. In this generalized version, $\widehat{HFL}$ of a link is the hat version of the unoriented link Floer homology of a corresponding balled link where every component is marked.
Recall that there has to be exactly one “marked point” on the link to define $I^{\natural}$, and that $I^{\sharp}$ of an unmarked link $L$ is defined as the $I^{\natural}$ of the marked link $L^{\sharp}$, given by adding a unlinked, marked unknot to $L$. For $I^{\sharp}$ and $I^{\natural}$, there are no “basepoints” on unmarked link components, and so Dehn twisting along an unmarked link component induces the identity. Although we use basepoints in unoriented link Floer homology, the last statement is true in unoriented link Floer homology as well, by Proposition 3.10.
As usual, the infinity version is simple.
Proposition 3.6.
Let $L$ be a balled link in $S^{3}$, and let the coefficient ring be $R=\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{n}^{k_{n}}\rrbracket$. Then $\boldsymbol{HFL}^{\prime}{}^{\infty}(S^{3},L)$ does not depend on the underlying link. In other words,
$\boldsymbol{HFL}^{\prime}{}^{\infty}(S^{3},L)\simeq\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{n}^{k_{n}},U\rrbracket[U_{1}^{-1},\cdots,U_{n}^{-1},U^{-1}]/(U_{1}-U,\cdots,U_{n}-U).$
Proof.
This argument is standard. Consider a weakly admissible Heegaard diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ for $(S^{3},L)$. We can also assume that $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{z})$ is also weakly admissible. Let $\boldsymbol{CF}^{\prime}{}^{\infty}$ be the usual, unoriented link Floer chain complex, and let the corresponding weight function be $w^{\prime}$. Let $\boldsymbol{CF}^{\infty}$ be the chain complex whose underlying group is the same as that of $\boldsymbol{CF}^{\prime}{}^{\infty}$, but if the $l$th baseball $B_{l}$ is a link baseball, then assign the weight $U_{l}$ to the $z$-basepoint corresponding to $B_{l}$ and assign the weight $1$ to the $w$-basepoint corresponding to $B_{l}$ If the $l$th baseball $B_{l}$ is a free baseball, then let $k_{l}=1$ and assign the weight $U_{l}$ to the free basepoint corresponding to $B_{l}$. Let this new weight function be $w$.
We claim that $\boldsymbol{CF}^{\prime}{}^{\infty}$ and $\boldsymbol{CF}^{\infty}$ are isomorphic as chain complexes over $R^{\infty}$. Fix an intersection point ${\bf x}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$, and for each ${\bf y}\in\boldsymbol{\alpha}\cap\boldsymbol{\beta}$, choose a domain ${\mathcal{}D}_{{\bf y}}\in D({\bf x},{\bf y})$. Then the $R^{\infty}$-linear map $\boldsymbol{CF}^{\prime}{}^{\infty}\to\boldsymbol{CF}^{\infty}$ given by
${\bf y}\mapsto\frac{w({\mathcal{}D}_{{\bf y}})}{w^{\prime}({\mathcal{}D}_{{\bf y}})}{\bf y}$
is a chain map, since $w^{\prime}({\mathcal{}P})=w({\mathcal{}P})$ for any cornerless two-chain ${\mathcal{}P}$ (indeed, $n_{z_{i}}({\mathcal{}P})=n_{w_{i}}({\mathcal{}P})$ for all $i$). It is an isomorphism since it has an obvious inverse. ∎
The homology groups $\boldsymbol{HFL}^{\prime}{}^{-}(Y,L)$, $\boldsymbol{HFL}^{\prime}{}^{\infty}(Y,L)$, and $\widehat{HFL^{\prime}}(Y,L)$ are well-defined as groups. See [JTZ21], [HM17, Proposition 2.3], [Zem19, Proposition 3.5].
Proposition 3.7 (Naturality).
Let $L$ be a balled link in a three-manifold $Y$. Given two Heegaard data ${\mathcal{}H},{\mathcal{}H}^{\prime}$ that represent $L\subset Y$, there is a chain homotopy equivalence
$\boldsymbol{CF}^{-}({\mathcal{}H})\to\boldsymbol{CF}^{-}({\mathcal{}H}^{\prime})$
over the coefficient ring $R$, which is well-defined and functorial up to homotopy.
Proof.
The statement holds if the auxiliary datum $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ is fixed, by [JTZ21]. We would like to define a chain homotopy equivalence, canonical up to homotopy, for two choices of auxiliary data.
For two choices of link basepoints $w_{i},z_{i}$, the chain homotopy equivalence is given by the basepoint moving map. We move the link basepoints only inside the corresponding link baseball, and so this is well-defined up to homotopy.
For free baseballs, given two choices of the free basepoint and the push-off of the link away from the free basepoint, we claim that the chain homotopy equivalence that is induced by isotoping the link and the basepoint inside the free baseball does not depend on the isotopy, up to chain homotopy. First note that the basepoint moving map corresponding to a meridional loop is homotopic to the identity, since it can be realized as a free-stabilization region going around the corresponding link basepoint. (We thank Ian Zemke for pointing this out: that using [MO10, Proposition 6.2] and/or [Zem15, Proposition 14.20] is unnecessary.) Now, a map given by isotoping the link and the basepoint is homotopic to the composition of the homotopy equivalences (canonical up to homotopy) induced by first pushing the basepoint to the boundary of the free baseball, then isotoping the unknotted strand (intersection of the link and the free baseball) inside the interior of the free baseball, and then moving the basepoint back to where it was. Indeed, this is easy to check if the isotopy is “small,” i.e. when we can move the basepoint to the boundary and back without intersecting the support of the isotopy, but any isotopy is a composition of small isotopies.
Finally, consider both reversing the orientation of a link component of $\overrightarrow{L}$ and swapping the labels of the two corresponding link basepoints at the same time. (Note that only doing one of these does not result in an auxiliary datum for the same balled link.) Given a Heegaard diagram for one of these, swapping the labels of the two relevant link basepoints results in a Heegaard diagram for the other. This induces an isomorphism between the chain complexes, and we can check that this is well-defined, squares to the identity, and commutes with all the above maps. ∎
Remark 3.8.
In fact, since the coefficient ring is of the form $\widetilde{R}\llbracket V_{1},\cdots,V_{l}\rrbracket$ where $V_{1},\cdots,V_{l}$ are the variables corresponding to the unmarked free basepoints, we have that any two choices of the unmarked free basepoints give canonically isomorphic homology groups, since in [Zem15, Proposition 14.20], only the $CF^{-}({\mathcal{}H},\mathfrak{s})\otimes_{\mathbb{F}_{2}}\left\langle\theta^{+}\right\rangle$ part survives in homology. (Compare [Zem15, Corollary 14.19].)
Remark 3.9.
The space ${\rm Spin}^{c}(Y(\overrightarrow{L}))$ of ${\rm Spin}^{c}$-structures for all the different choices of auxiliary datum $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ are canonically identified; and so we can consider ${\rm Spin}^{c}(Y(L))$ and define $\boldsymbol{CFL^{\prime}}^{-}(Y,L;\mathfrak{s})$ for $\mathfrak{s}\in{\rm Spin}^{c}(Y(L))$. These are also canonical in the above sense.
If we only want to consider finitely many holomorphic disks, then we can work with $\mathfrak{s}$-strongly admissible Heegaard diagrams (Definition 2.50). Hence, we can also work over polynomial rings and define $CFL^{\prime}{}^{-}(Y,L;\mathfrak{s})$.
Proposition 3.10.
Let $L\subset Y$ be a balled link. Then, Dehn twisting along an unmarked link component induces the identity on both $\boldsymbol{HFL}^{\prime}{}^{-}(Y,L)$ and $\widehat{HFL^{\prime}}(Y,L)$.
Proof.
We use the explicit formula given in [Sar15], [Zem17, Theorem B]. Although we have free baseballs, we can consider our chain complex as a canonical direct summand of the chain complex given by putting an unknotted unknot for each free baseball so that they intersect, and change its label to a link baseball (compare the discussion at the start of Subsection 3.3). Hence, the formula still holds.
Let $w,z$ be the two basepoints on the link component that we Dehn twist along. For notational convenience, let $U^{1/2}$ be the variable that they correspond to, and let $U_{i}$ be the other variables. The point is that the coefficient ring is $R^{\prime}\llbracket U^{1/2}\rrbracket$ for some $R^{\prime}$. Recall that we define
$\Phi_{z}{\bf x}=U^{-1/2}\sum_{{\bf y}\in\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}}}\sum_{\phi\in D({\bf x},{\bf y}),\ \mu(\phi)=1}n_{z}(\phi)\#\mathcal{M}(\phi)U^{n_{w}(\phi)+n_{z}(\phi)}\prod_{i}U_{i}^{\frac{1}{2}\left(n_{w_{i}}(\phi)+n_{z_{i}}(\phi)\right)}{\bf y}.$
and that $\Phi_{w}$ is defined similarly. The map induced by the Dehn twist is
${\rm Id}+\Phi_{w}\Phi_{z}.$
This is the identity on homology, since $\Phi_{w}+\Phi_{z}$ is 00 on homology by the discussion right before [Zem15, Corollary 14.19] and $\Phi_{z}^{2}$ is also 00 on homology [Zem17, Lemma 9.7]. Indeed, $\Phi_{w}+\Phi_{z}$ is given by differentiating the coefficients of $\partial$ considered as a matrix with entries in $R$, with respect to $U^{1/2}$, as in [Zem15, Remark 14.12]. The argument of [Zem17, Lemma 9.7] shows that
$H{\bf x}=U^{-1}\sum_{{\bf y}\in\mathbb{T}_{\boldsymbol{\alpha}}\cap\mathbb{T}_{\boldsymbol{\beta}}}\sum_{\phi\in D({\bf x},{\bf y}),\ \mu(\phi)=1}\frac{n_{z}(\phi)(n_{z}(\phi)-1)}{2}\#\mathcal{M}(\phi)U^{n_{w}(\phi)+n_{z}(\phi)}\prod_{i}U_{i}^{\frac{1}{2}\left(n_{w_{i}}(\phi)+n_{z_{i}}(\phi)\right)}{\bf y}$
is an $R$-linear chain homotopy $\Phi_{z}^{2}\simeq 0$. ∎
Definition 3.11.
A band on a balled link $L\subset Y$ is a band on the underlying link that does not intersect any of the baseballs. A balled link $L^{\prime}\subset Y$ is obtained by a band move on a balled link $L$ if the underlying link $L^{\prime}$ is obtained by a band move on the underlying link $L$, the orderings of the baseballs are the same, and the labels and the markings of the balls of $L^{\prime}$ and $L$ are the same, except that
  • if the band is a merge band, then exactly one unmarked link baseball of $L$ becomes an unmarked free baseball;
  • if the band is a split band, then exactly one unmarked free baseball of $L$ becomes an unmarked link baseball.
Note that if the band is a merge or a split band, then the label of at least one baseball has to change in order for $L^{\prime}$ to be a balled link. Also note that the baseball that changes label in the above definition is never marked.
3.2. Heegaard triple subordinate to a band
To define band maps, we first define what it means for a Heegaard triple to be subordinate to a band. Since the baseballs’ labels might change depending on the type of the band, we handle each band type separately.
We closely follow [Zem19, Section 6.1]; note that our bands are always $\beta$-bands. Also see Figure 1.5. Also compare [OS06, Section 4.1].
Figure 3.2. The region $H$, the tangle $\overrightarrow{L_{0}}$, and the $\beta$ and $\gamma$ circles for Definitions 3.13 and 3.15. (We have chosen the labels $w_{2}$ and $z_{2}$ arbitrarily.)
3.2.1. Non-orientable bands
Definition 3.12.
Let $B$ be a non-orientable band from a balled link $L$ to $L^{\prime}$. A tuple of auxiliary data $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$ for $L$ and $L^{\prime}$, respectively, is adapted to $B$ if:
  1. (1)
    the basepoints for $L$ and $L^{\prime}$ are the same, i.e. $\boldsymbol{w}=\boldsymbol{w}^{\prime}$, $\boldsymbol{z}=\boldsymbol{z}^{\prime}$, and $\boldsymbol{v}=\boldsymbol{v}^{\prime}$; and
  2. (2)
    $\overrightarrow{L^{\prime}}$ is given by surgering $\overrightarrow{L}$ along $B$.
Define the sutured manifold obtained from $Y\backslash N(\overrightarrow{L}\cup B\cup\boldsymbol{v})$ as follows. On $\partial N(\overrightarrow{L}\cup B)$, we add meridional sutures for each basepoint of $\boldsymbol{w}\sqcup\boldsymbol{z}$, such that $R_{-}$ [^40] Our convention is that $R_{-}$ corresponds to the $\alpha$-handlebody part. corresponds to the parts of $\partial N(\overrightarrow{L}\cup B)$ that live over the connected components of $\overrightarrow{L}\backslash(\boldsymbol{w}\sqcup\boldsymbol{z})$ that are contained in the link baseballs. Also, add one suture to each connected component of $\partial N(\boldsymbol{v})$.
Let $\overrightarrow{L_{0}}$ be the connected component of $\overrightarrow{L}\backslash(\boldsymbol{w}\sqcup\boldsymbol{z})$ that intersects the band $B$, let $H\subset N(\overrightarrow{L}\cup B)$ be the points that live over $\overrightarrow{L_{0}}\cup B$, and let $S=\partial H\cap\partial N(\overrightarrow{L}\cup B)=\partial H\cap R_{+}$. In particular, $H$ is a solid torus, and $S$ is a torus minus two disks. See Figure 3.2 for a schematic.
Definition 3.13.
Let $L,L^{\prime}\subset Y$ be balled links, and let $B$ be a non-orientable band from $L$ to $L^{\prime}$. Let $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$ be auxiliary data for $L$ and $L^{\prime}$ that are adapted to $B$. We say that the Heegaard triple
$(\Sigma,\boldsymbol{\alpha}=\{\alpha^{1},\cdots,\alpha^{n}\},\boldsymbol{\beta}=\{\beta^{1},\cdots,\beta^{n}\},\boldsymbol{\gamma}=\{\gamma^{1},\cdots,\gamma^{n}\},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$
is subordinate to $B$ and the auxiliary data $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$, if:
  1. (1)
    The diagram $(\Sigma,\{\alpha^{1},\cdots,\alpha^{n}\},\{\beta^{2},\cdots,\beta^{n}\},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ corresponds to the sutured manifold $Y\backslash N(\overrightarrow{L}\cup B\cup\boldsymbol{v})$ defined above.
  2. (2)
    The circles $\gamma^{2},\cdots,\gamma^{n}$ are standard translates of $\beta^{2},\cdots,\beta^{n}$.
  3. (3)
    Let $\overrightarrow{L_{0}},H,S$ be as above. Let $\beta\subset S$ be the closed curve that is defined up to isotopy (one can show this, for instance, by an innermost disk/arc argument) by the property that $\beta$ bounds a disk $D_{\beta}$ in $H$, such that $D_{\beta}$ does not intersect $\overrightarrow{L}$ and is non-separating in $H$. Similarly, let $\gamma\subset S$ be the closed curve that bounds a non-separating disk $D_{\gamma}$ in $H$, that does not intersect $\overrightarrow{L^{\prime}}$.
    Then, $\beta^{1}$ is a closed curve that is obtained by projecting $\beta$ onto $\Sigma\backslash(\beta^{2}\cup\cdots\cup\beta^{n})$ and $\gamma^{1}$ is a curve that is obtained by projecting $\gamma$ onto $\Sigma\backslash(\gamma^{2}\cup\cdots\cup\gamma^{n})$.
A Heegaard triple is subordinate to $B$, if there is some choice of auxiliary data for $L$ and $L^{\prime}$ adapted to $B$, such that the Heegaard triple is subordinate to $B$ and those auxiliary data.
3.2.2. Merge and split bands
Let us first consider merge bands.
Definition 3.14.
Let $B$ be a merge band from a balled link $L$ to $L^{\prime}$. Let $B_{z}$ be the link baseball of $L$ that becomes a free baseball in $L^{\prime}$. A tuple of auxiliary data $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$ for $L$ and $L^{\prime}$, respectively, is adapted to $B$ if:
  1. (1)
    if $w\in\boldsymbol{w}$ and $z\in\boldsymbol{z}$ are the two basepoints that correspond to the link baseball $B_{z}$, then $\boldsymbol{w}^{\prime}=\boldsymbol{w}\backslash\{w\}$, $\boldsymbol{z}^{\prime}=\boldsymbol{z}\backslash\{z\}$, and $\boldsymbol{v}^{\prime}=\boldsymbol{v}\cup\{z\}$; and
  2. (2)
    in the complement of the baseball $B_{z}$, $\overrightarrow{L^{\prime}}$ coincides with the link given by surgering $\overrightarrow{L}$ along $B$.
Define the sutured manifold obtained from $Y\backslash N(\overrightarrow{L}\cup B\cup\boldsymbol{v})$ using the exact same definition as in Subsubsection 3.2.1. Define $\overrightarrow{L_{0}},H,S$ in the same way as well. In this case, $H$ is a $3$-ball, and $S$ is a sphere minus four disks. The tangle $\overrightarrow{L_{0}}\subset H$ is a $4$-ended trivial (i.e. rational) tangle. Let $\overrightarrow{L_{0}^{\prime}}\subset H$ be the tangle given by surgering $\overrightarrow{L_{0}}$ along $B$. Note that $\overrightarrow{L_{0}^{\prime}}\neq\overrightarrow{L^{\prime}}\cap H$. See Figure 3.2 for a schematic.
Definition 3.15.
Let $L,L^{\prime}\subset Y$ be balled links, and let $B$ be a merge band from $L$ to $L^{\prime}$. Let $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$ be auxiliary data for $L$ and $L^{\prime}$, respectively, adapted to $B$. We say that the Heegaard triple
$(\Sigma,\boldsymbol{\alpha}=\{\alpha^{1},\cdots,\alpha^{n}\},\boldsymbol{\beta}=\{\beta^{1},\cdots,\beta^{n}\},\boldsymbol{\gamma}=\{\gamma^{1},\cdots,\gamma^{n}\},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$
is subordinate to $B$ and the auxiliary data $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z},\boldsymbol{v})$ and $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime},\boldsymbol{v}^{\prime})$, if:
  1. (1)
    The diagram $(\Sigma,\{\alpha^{1},\cdots,\alpha^{n}\},\{\beta^{2},\cdots,\beta^{n}\},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ corresponds to the sutured manifold $Y\backslash N(\overrightarrow{L}\cup B\cup\boldsymbol{v})$ defined above.
  2. (2)
    There exists a path in $\Sigma\backslash\boldsymbol{\alpha}$ between $w$ and $z$ that intersects the circle $\beta^{2}$ at exactly one point, and is disjoint from all the other $\beta$-circles and all the $\gamma$-circles.
  3. (3)
    The circle $\gamma^{2}$ bounds a disk $D$ in $\Sigma$ such that $D\cap(\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})=\{w,z\}$, where $w,z$ are as in Definition 3.14.
  4. (4)
    The circles $\gamma^{3},\cdots,\gamma^{n}$ are standard translates of $\beta^{3},\cdots,\beta^{n}$.
  5. (5)
    Let $\overrightarrow{L_{0}},H,S$ be as above. Let $\beta\subset S$ be the closed curve (defined up to isotopy) that bounds a disk $D_{\beta}$ in $H$, such that $D_{\beta}$ does not intersect $\overrightarrow{L}$ and separates $H$ into two components, such that each component has one component of $\overrightarrow{L_{0}}$. Similarly, let $\gamma\subset S$ be the closed curve that bounds a disk $D_{\gamma}$ in $H$ that does not intersect $\overrightarrow{L^{\prime}}$ and separates $H$ into two components, such that each component has one component of $\overrightarrow{L_{0}^{\prime}}$.
    Then, $\beta^{1}$ is a closed curve that is obtained by projecting $\beta$ onto $\Sigma\backslash(\beta^{2}\cup\cdots\cup\beta^{n})$ and $\gamma^{1}$ is a curve that is obtained by projecting $\gamma$ onto $\Sigma\backslash(\gamma^{2}\cup\cdots\cup\gamma^{n})$.
A Heegaard triple is subordinate to $B$, if there is some choice of auxiliary data for $L$ and $L^{\prime}$ adapted to $B$, such that the Heegaard triple is subordinate to $B$ and those auxiliary data.
Now, we can define Heegaard triples subordinate to split bands in a simple way.
Definition 3.16.
Let $L,L^{\prime}\subset Y$ be balled links, and let $B$ be a split band from $L$ to $L^{\prime}$. Consider the dual, merge band $B^{\ast}$ from $L^{\prime}$ to $L$. We say that the Heegaard triple $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ is subordinate to $B$ if the Heegaard triple $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\gamma},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ is subordinate to $B^{\ast}$. (We similarly define auxiliary data adapted to $B$ and Heegaard triples subordinate to $B$ and some auxiliary data.)
Remark 3.17.
For each of Definitions 3.13, 3.15, and 3.16, we can check that $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ represents $(\overrightarrow{L},\boldsymbol{w}\sqcup\boldsymbol{z})$ inside $(Y,\boldsymbol{v})$ and that the diagram $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\gamma},\boldsymbol{v}^{\prime}\sqcup\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime})$ represents $(\overrightarrow{L^{\prime}},\boldsymbol{w}^{\prime}\sqcup\boldsymbol{z}^{\prime})$ inside $(Y,\boldsymbol{v}^{\prime})$.
3.2.3. The subdiagram $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ of a Heegaard triple $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ subordinate to a band
Let us study the Heegaard diagram $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ and the chain complex associated to it.
Lemma 3.18.
Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a Heegaard triple subordinate to a band $B$ from $L$ to $L^{\prime}$. Then $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ is a diagram for the connected sum of a pointed link $M\subset S^{1}\times S^{2}$ with a pointed unlink in a pointed #$\#^{n}(S^{1}\times S^{2})$. Each component of $M$ and the unlink has exactly two basepoints.
If the band $B$ is non-orientable, then the link $M$ is given by the left hand side of Figure 3.3, and if the band $B$ is a merge or a split band, then $M$ is given by the right hand side of Figure 3.3 (ignore the labels $D_{i}$ of the elementary two-chains), i.e. $M=S^{1}\times\{a,b\}\subset S^{1}\times S^{2}$ for some $a,b\in S^{2}$.
Proof.
This follows from the same argument as [Zem19, Lemma 6.4]. Let us use the same notations as in Definitions 3.13, 3.15, and 3.16.
If $B$ is non-orientable, then after isotoping and handlesliding the $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ curves, the Heegaard diagram $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ becomes a stabilization [^41] Strictly speaking, we allow renaming the basepoints and the variables. of the left hand side of Figure 3.3, where the basepoints $w$ and $z$ are the basepoints of the component of the link that $B$ intersects.
Similarly, if $B$ is a merge or a split band, then we can arrange the diagram such that it is a stabilization of the left hand side of Figure 3.3. If $B$ is a merge band, then $w_{1}$ and $z_{1}$ correspond to the basepoints $w$ and $z$ in Definition 3.14; and $w_{2},z_{2}$ are the basepoints on the other component of the link $\overrightarrow{L}$ that $B$ intersects. (The roles of $w_{1},z_{1}$ and $w_{2},z_{2}$ can be reversed.) ∎
Figure 3.3. Local Heegaard diagrams for the band maps. Unless otherwise specified, for the diagram on the left, work over $\mathbb{F}\llbracket U^{1/2}\rrbracket$ and assign the weight $U^{1/2}$ to both $w$ and $z$; for the diagram on the right, work over $\mathbb{F}\llbracket U_{1}^{1/2},U_{2}^{1/2}\rrbracket$ and assign the weight $U_{i}^{1/2}$ to $w_{i}$ and $z_{i}$.
Let us study Heegaard diagrams on the right of Figure 3.3.
Lemma 3.19.
Consider the Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\beta},\boldsymbol{\gamma},\{w_{1},z_{1},w_{2},z_{2}\})$ given by the right hand side of Figure 3.3. We work over the ring $R=\mathbb{F}[W_{1},Z_{1},W_{2},Z_{2}]$, where we assign weights $W_{1},Z_{1},W_{2},Z_{2}$ to $w_{1},z_{1},w_{2},z_{2}$, respectively. Then $CF^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ is one of the following, depending on the almost complex structure.
${ax}$${ay}$${ax}$${ay}$${bx}$${by}$${bx}$${by}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{z_{1}+z_{2}}$$\scriptstyle{z_{1}+z_{2}}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{z_{1}+z_{2}}$$\scriptstyle{z_{1}+z_{2}}$$\scriptstyle{z_{1}+z_{2}}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{w_{1}+w_{2}}$$\scriptstyle{z_{1}+z_{2}}$
Proof.
Note that the Heegaard diagrams $(\mathbb{T}^{2},\boldsymbol{\beta},\boldsymbol{\gamma},\{z_{1},z_{2}\})$ and $(\mathbb{T}^{2},\boldsymbol{\beta},\boldsymbol{\gamma},\{w_{1},w_{2}\})$ are weakly admissible, and the “fully blocked” chain complex $\widetilde{CF}(\boldsymbol{\beta},\boldsymbol{\gamma})$ (i.e. the chain complex is freely generated over $\mathbb{F}$ and we only count differentials without any basepoints) of either of these Heegaard diagrams have no differential, since there are four generators and $\widetilde{HF}$ has rank four.
The following are the two-chains of all the Maslov index $1$ domains:
  • $ax\to bx,ay\to by$: $D_{2}$, $D_{3}+D_{6}$
  • $bx\to ax,by\to ay$: $D_{4}+D_{5}$, $D_{1}+D_{3}+D_{4}$, $D_{1}+D_{7}$
  • $ax\to ay,bx\to by$: $D_{3}+D_{4}$, $D_{7}$
  • $ay\to ax,by\to bx$: $D_{1}+D_{2}$, $D_{5}+D_{6}$, $D_{1}+D_{3}+D_{6}$
Among these, the domains with two-chain $D_{2}$, $D_{3}+D_{6}$, $D_{4}+D_{5}$, $D_{3}+D_{4}$, $D_{7}$, or $D_{5}+D_{6}$ always have exactly one holomorphic representative regardless of the almost complex structure.
However, $D_{1}+D_{7}\in D(bx,ax)$, for instance, depends on the almost complex structure. Note that exactly one of $D_{1}+D_{3}+D_{4}$ and $D_{1}+D_{7}\in D(bx,ax)$ has an odd number of holomorphic representatives, since $\widetilde{CF}$ of $(\mathbb{T}^{2},\boldsymbol{\beta},\boldsymbol{\gamma},\{z_{1},z_{2}\})$ has no differentials. Similarly, exactly one of these two domains, viewed as domains in $D(by,ay)$, contributes to the differential, and the corresponding statement holds for $D_{1}+D_{2}$ and $D_{1}+D_{3}+D_{6}$.
Now, the lemma follows from that $\left(\partial^{-}\right)^{2}=0$ for the chain complex $CF^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ over $R$. Indeed, we get two cases depending on whether $D_{1}+D_{3}+D_{4}$ or $D_{1}+D_{7}$ contributes to the differential. ∎
Hence, we can compute the homology groups for the two Heegaard data of Figure 3.3. First, $H_{1}^{orb}$ of the corresponding $L\subset Y=S^{1}\times S^{2}$ is (non-canonically) isomorphic to $\mathbb{Z}\oplus\mathbb{Z}/2$ where the meridian represents $(0,1)$. They are both homologically $\mathbb{Z}$-gradable and Alexander $\mathbb{Z}/2$-splittable; the $\mathbb{Z}/2$-splitting corresponds to the collapsed ${\rm Spin}^{c}$-splitting under a projection $\mathbb{Z}\oplus\mathbb{Z}/2\to\mathbb{Z}/2$.
In both cases, the homology $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ is $f\mathbb{F}\llbracket U^{1/2}\rrbracket\oplus g\mathbb{F}\llbracket U^{1/2}\rrbracket$, where $(f,g)=(\tau,\sigma)$ for the diagram on the left, and $f=ay+bx$ and $g=by$ or $ax+by$ for the diagram on the right. Both $f$ and $g$ are homogeneous with respect to both the $\mathbb{Z}$-grading and the ${\rm Spin}^{c}$-splitting (and hence also the Alexander $\mathbb{Z}/2$-grading), and they lie in the same $\mathbb{Z}$-grading, but their ${\rm Spin}^{c}$-structures and Alexander $\mathbb{Z}/2$-gradings differ. [^42] In fact, $f$ and $g$ have different relative homological $\mathbb{Z}/2$-gradings ${\rm gr}_{\boldsymbol{w}}$ and ${\rm gr}_{\boldsymbol{z}}$ (defined such that $U_{i}^{1/2}$ has grading 00), although we will never use them: it is possible to use these $\mathbb{Z}/2$-gradings to distinguish the two top $\mathbb{Z}$-grading generators, but we have chosen to use the Alexander $\mathbb{Z}/2$-gradings..
3.3. Band maps
We define the band maps for the minus version, on the chain level. These maps induce maps on the hat and infinity versions.
3.3.1. Coefficient rings and weight functions
We can define the band maps as a triangle counting map, $\mu_{2}$, where the Heegaard datum has nontrivial local systems. As hinted in Remark 2.16, we can avoid using nontrivial local systems by composing a $\mu_{2}$ with some purely algebraic maps. Remark 3.22 explains how the former way works.
In this subsubsection, we fix the notations that we use in Subsubsections 3.3.2 and 3.3.3.
Let us first consider the simple case, where $B$ is a non-orientable band from a balled link $L$ to $L^{\prime}$. Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a Heegaard triple subordinate to it. Then, we consider the coefficient ring $R$ and the weight function such that they, together with $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$, form a Heegaard datum that represents $L\subset Y$. Note that the Heegaard datum obtained by replacing $\boldsymbol{\beta}$ with $\boldsymbol{\gamma}$ represents $L^{\prime}\subset Y$. We also use the same coefficient ring and weight function for the Heegaard triple.
Now, let us focus on the case where $B$ is a merge or a split band from $L$ to $L^{\prime}$. Let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a Heegaard triple subordinate to it. The complication is that we cannot use the same coefficient ring (and even the same basepoints) for a Heegaard datum that represents $L$ and a Heegaard datum that represents $L^{\prime}$.
Indeed, say $B$ is a merge band. Let $B_{z}$ be the baseball that changes label, and let $w,z\in B_{z}$ be the two basepoints corresponding to $B_{z}$. Let $R=\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{n}^{k_{n}}\rrbracket$ be the coefficient ring and $w_{L}:\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\to R$ be the weight function such that they, together with $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$, form a Heegaard datum that represents $L\subset Y$. If $B_{z}$ is the $l$th baseball, then $k_{l}=1/2$. Let
$S=\mathbb{F}\llbracket U_{1}^{k_{1}},\cdots,U_{l-1}^{k_{l-1}},U_{l},U_{l+1}^{k_{l+1}}\cdots,U_{n}^{k_{n}}\rrbracket.$
Then, the Heegaard datum that consists of $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\backslash\{w\})$, coefficient ring $S$, and weight function
$w_{L^{\prime}}:(\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})\backslash\{w\}\to S:s\mapsto\begin{cases}U_{l}&{\rm if\ }s=z\\ w_{L}(s)&{\rm otherwise}\end{cases}$
represents the balled link $L^{\prime}\subset Y$.
Since $w$ and $z$ belong to the same connected component of $\Sigma\backslash(\boldsymbol{\alpha}\cup\boldsymbol{\gamma})$, we can either view $w$ and $z$ as two basepoints with weight $U_{l}^{1/2}$ or one basepoint with weight $U_{l}$. Formally speaking, if we define $\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$ using $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ and $w_{L}$, and define $\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$ using $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\backslash\{w\})$ and $w_{L^{\prime}}$, then the chain complex $\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$ naturally splits into the direct sum of two chain complexes $\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$ and $U_{l}^{1/2}\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$.
We use the coefficient ring $R$ and weight function $w_{L}$ for the Heegaard triple.
If $B$ is a split band, we can define the coefficient rings $R$ and $S$, and the weight functions using the dual, merge band $B^{\ast}$, as in Definition 3.16. For the Heegaard triple, we use the coefficient ring $R$ and the weight function for the link with more components.
3.3.2. A first definition
To define the band maps, we restrict our attention to those Heegaard triples $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ subordinate to a band $B$ such that Heegaard subdatum with attaching curves $\boldsymbol{\beta},\boldsymbol{\gamma}$ is a stabilization of one of the data in Figure 3.3. We will show that the band map is well-defined in Subsubsection 3.3.3; we will then also see that we could work with any Heegaard triple subordinate to a band.
Definition 3.20.
Let a Heegaard datum with underlying Heegaard diagram $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a stabilization of one of the two Heegaard data of Figure 3.3. Define the cycle $\Theta_{B}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ as follows.
  • If it is a stabilization of the left hand side of Figure 3.3, then let $\Theta_{B}=S(\tau)$ be the top grading generator that contains the intersection point $\tau$.
  • If it is a stabilization of the right hand side of Figure 3.3, then let $\Theta_{B}=S(ay+bx)$ be the sum of the top grading generators that contain the intersection points $a,y$ and $b,x$.
The elements $\Theta_{B}$ defined above are indeed cycles: first, they are cycles for the Heegaard data given in Figure 3.3, by Lemma 3.19. In the stabilized data, they are invariant under changing the almost complex structure, as there are no relevant nonconstant Maslov index 00 domains. By Proposition 2.39, they are cycles in a pinched almost complex structure, and so they are cycles in any almost complex structure. By abuse of notation, we denote the homology class that $\Theta_{B}$ represents as $\Theta_{B}$.
Definition 3.21.
Let $B$ be a band from $L$ to $L^{\prime}$, and let $(\Sigma,\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a weakly admissible Heegaard triple subordinate to it. Further assume that the Heegaard subdatum with attaching curves $\boldsymbol{\beta},\boldsymbol{\gamma}$ is a stabilization of one of the Heegaard data in Figure 3.3. Define the band maps as follows.
  • If $B$ is non-orientable, the band map is the triangle counting map
    $\mu_{2}(-,\Theta_{B}):\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\to\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}).$
  • If $B$ is a merge band, then the band map is the composite
    $\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\xrightarrow{\mu_{2}(-,\Theta_{B})}\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})\to U_{l}^{1/2}\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})\xrightarrow{\cdot U_{l}^{-1/2}}\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}),$
    where the second map is projection onto the summand.
  • If $B$ is a split band, then the band map is the composite
    $\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\to\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\xrightarrow{\mu_{2}(-,\Theta_{B})}\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}).$
    where the first map is inclusion into the summand.
If $B$ is non-orientable, then the band map is $R$-linear; if $B$ is a merge or split band, then it is $S$-linear.
Remark 3.22.
We can describe the merge and split band maps using local systems. We consider the merge band map first. Assume that Heegaard subdatum with attaching curves $\boldsymbol{\beta},\boldsymbol{\gamma}$ is a stabilization of the right side of Figure 3.3. Work over the coefficient ring $S$. Instead of considering both basepoints $w$ and $z$, we only consider one of them, say $z$. Let $G$ be an oriented arc from $z$ to $w$, which is disjoint from $\boldsymbol{\alpha}$ and $\boldsymbol{\gamma}$, and intersects $\boldsymbol{\beta}$ at exactly one point. Consider the rank $2$ local system $E$ on $\boldsymbol{\beta}$ described in Subsection 2.2. Now, $\boldsymbol{CF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\simeq\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}^{E})$, and under this canonical identification, the merge band map is
$\mu_{2}(-,e_{1}^{\ast}\Theta_{B}):\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}^{E})\to\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}).$
The split band map is similar: in this case, $\boldsymbol{\gamma}$ is the attaching curve that intersects $G$ and has a nontrivial local system, and the split band map is
$\mu_{2}(-,e_{0}\Theta_{B}):\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\to\boldsymbol{CF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}^{E}).$
3.3.3. Well-definedness
As in [OS06, Proposition 4.6] and [Zem19, Lemma 6.5], one key step of proving that the band maps are well-defined is showing that $\Theta_{B}$ is “invariant under handleslides,” which is usually proven using that $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ is very simple. Our case is slightly more complicated; we exploit the ${\rm Spin}^{c}$-splitting from Subsubsection 2.8.1.
Lemma 3.23.
Let a Heegaard datum with underlying Heegaard diagram $(\Sigma,\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ be a stabilization of one of the two Heegaard data of Figure 3.3. Then the homology class $\Theta_{B}\in\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ lies in a torsion ${\rm Spin}^{c}$-structure, and hence $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma};\mathfrak{s}(\Theta_{B}))$ has a relative $\mathbb{Z}$-grading. The homology class $\Theta_{B}$ is the unique non-trivial homology class in the top grading part of $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma};\mathfrak{s}(\Theta_{B}))$.
Proof.
Recall the computation of $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ for the destabilized datum from right after Lemma 3.19. Also, one can compute $c_{1}(\mathfrak{s}(\Theta_{B}))$ using Equation 2.3; we discussed the left one in Example 2.48. ∎
We can consider this property of $\Theta_{B}$ for any weakly admissible $(\Sigma,\widetilde{\boldsymbol{\beta}},\widetilde{\boldsymbol{\gamma}},\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$ such that $\widetilde{\boldsymbol{\beta}}$ and $\widetilde{\boldsymbol{\gamma}}$ be strongly equivalent to $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ [^43] This in particular holds if $\widetilde{\boldsymbol{\beta}}$ and $\widetilde{\boldsymbol{\gamma}}$ come from a Heegaard triple diagram subordinate to a band., respectively (i.e. they are related by handleslides and isotopies in $\Sigma\backslash(\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z})$), and hence define $\Theta_{B}\in\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\beta}},\widetilde{\boldsymbol{\gamma}})$, as hinted in Subsubsection 3.3.2. Also the $\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\beta}},\widetilde{\boldsymbol{\gamma}})$’s for such $\widetilde{\boldsymbol{\beta}}$ and $\widetilde{\boldsymbol{\gamma}}$ are canonically isomorphic; $\Theta_{B}$ is invariant under this isomorphism since the isomorphism preserves the above properties.
Remark 3.24.
In practice, it is easy to compute the relative homological grading, but computing the ${\rm Spin}^{c}$-structure of generators is not as easy. Here is a more practical way to handle this: the following condition, together with that $\Theta_{B}$ has the top grading in $\bigoplus_{c_{1}(\mathfrak{s})=0}\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma};\mathfrak{s})$ and is homogeneous with respect to the ${\rm Spin}^{c}$-splitting [^44] The Alexander $\mathbb{Z}/2$-splitting is sufficient in this case. uniquely characterizes $\Theta_{B}$. Consider the chain complex ${\mathcal{}CF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ given by working over the coefficient ring ${\mathcal{}R}=\mathbb{F}\llbracket\{V_{j}\},\{W_{i}\},\{Z_{i}\}\rrbracket$ and assigning the weights $V_{j},W_{i},Z_{i}$ to $v_{j},w_{i},z_{i}$, respectively. Consider $R$ as an ${\mathcal{}R}$-module by identifying the weights of $v_{j},w_{i},z_{i}$ (in $R$) with $V_{j},W_{i},Z_{i}$, respectively.
Then, we have a map
${\mathcal{}HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})\otimes_{{\mathcal{}R}}R\to\boldsymbol{HF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma}).$
If the Heegaard datum is a stabilization of the left diagram of Figure 3.3, then $\Theta_{B}$ is in the image of this map; if it is a stabilization of the right diagram, then $\Theta_{B}$ is not in the image of this map.
Now, we can show that the band maps are well-defined.
Proposition 3.25.
On homology, the band maps do not depend on the Heegaard triple subordinate to the band.
Proof.
Let $B$ be a band from a balled link $L$ to $L^{\prime}$. We will in fact show that $\mu_{2}(-,\Theta_{B}):\boldsymbol{CF}_{R}^{-}\to\boldsymbol{CF}_{R}^{-}$ (instead of the band map which may involve composing $\mu_{2}$ with purely algebraic maps) does not depend, up to homotopy, on the Heegaard triple subordinate to the band. We omit the subscript $R$ from now on.
Let us first fix an arbitrary choice of auxiliary data adapted to $B$ and show that $\mu_{2}(-,\Theta_{B})$ does not depend, up to homotopy, on the Heegaard triple subordinate to $B$ and the auxiliary data. The proof is basically the same as [OS06, Proposition 4.6] and [Zem19, Lemma 6.5]. One first shows that two Heegaard triples subordinate to the band can be related by some moves as in [OS06, Lemma 4.5] and [Zem19, Lemma 6.3]. The maps are invariant under stabilization because of the same reasons (see Proposition 2.40), and that they are invariant under various handleslides and isotopies follows from the discussion after Lemma 3.23.
Now, let us show that the band maps do not depend on the Heegaard triple. For any two choices of auxiliary data adapted to $B$, there exists a finite sequence of auxiliary data adapted to $B$ such that for each consecutive pair, there exists a Heegaard triple that is subordinate to $B$ for both auxiliary data, if we allow moving the basepoints inside each elementary two-chain of the Heegaard triple diagram and allow swapping the labels (whether they are $w$ or $z$) of two link basepoints in the same link component. Hence the band maps are well-defined. ∎
3.3.4. ${\rm Spin}^{c}$-structures and gradings
If $B$ is a band from a balled link $L$ to $L^{\prime}$, then we will not relate ${\rm Spin}^{c}(Y(L))$ and ${\rm Spin}^{c}(Y(L^{\prime}))$. However, we have $c_{1}(\mathfrak{s}(\Theta_{B}))=0$ by Lemma 3.23, and so Proposition 2.65 implies that for each $c\in H^{2}(Y)$, the band maps restrict to
$\boldsymbol{HFL^{\prime}}^{-}(Y,L;c)\to\boldsymbol{HFL^{\prime}}^{-}(Y,L^{\prime};c),$
and that it preserves the relative homological grading.
We can work with $c$-strongly admissible triples (Definition 2.66) and define $HFL^{\prime}{}^{-}(Y,L;c)\to HFL^{\prime}{}^{-}(Y,L^{\prime};c)$ as well.
4. An unoriented skein exact triangle
We can now state Theorem 1.16 precisely. Recall that for merge and split bands between balled links, the label of exactly one baseball had to change.
Definition 4.1.
Three balled links $L_{a},L_{b},L_{c}\subset Y$ form an unoriented skein triple if there exists a ball $B^{3}\subset Y$ that is disjoint from all the baseballs, such that the underlying links are identical outside $B^{3}$, and they differ as in Figure 1.1 inside $B^{3}$. All the baseballs, the orderings of the baseballs, and the markings are identical, except that there exists exactly one baseball whose label can change.
Theorem 4.2.
If three balled links $L_{a},L_{b},L_{c}\subset Y$ form an unoriented skein triple, then the band maps given by the bands described in Figure 1.2 induce exact triangles
$\cdots\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{b})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{c})\to\boldsymbol{HFL^{\prime}}^{-}(Y,L_{a})\to\cdots$
$\cdots\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,L_{a})\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,L_{b})\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,L_{c})\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,L_{a})\to\cdots$
$\cdots\to\widehat{HFL^{\prime}}(Y,L_{a})\to\widehat{HFL^{\prime}}(Y,L_{b})\to\widehat{HFL^{\prime}}(Y,L_{c})\to\widehat{HFL^{\prime}}(Y,L_{a})\to\cdots.$
The proof of Theorem 4.2 boils down to the following local computation on the torus.
Theorem 4.3.
Consider the multi Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c},\{v,w,z\},G)$ given by Figure 4.1. Work over the ring $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$, and assign the weight $U_{1}$ to $v$, and $U_{2}^{1/2}$ to $w$ and $z$. The attaching curve $\boldsymbol{\beta}_{c}$ has a nontrivial, rank $2$ local system given by the oriented arc $G$ as in Subsection 2.2. Let
$\boldsymbol{\rho}=e_{0}\rho_{1}+e_{0}\rho_{2}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c}^{E}),\ \boldsymbol{\sigma}=e_{1}^{\ast}\sigma_{1}+e_{1}^{\ast}\sigma_{2}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\boldsymbol{\beta}_{a}).$
[…]
${{\boldsymbol{\beta}_{a}}}$${{\boldsymbol{\beta}_{b}}}$${{\boldsymbol{\beta}_{c}^{E}}}$${{\boldsymbol{\beta}_{c}^{E}}}$${{\boldsymbol{\beta}_{a}}}$${{\boldsymbol{\beta}_{b}}}$$\scriptstyle{\tau}$$\scriptstyle{\boldsymbol{\rho}}$$\scriptstyle{\boldsymbol{\sigma}}$$\scriptstyle{\tau}$
Figure 4.1. The genus $1$ diagram for Theorem 4.3
Proof of Theorem 4.2 assuming Theorem 4.3.
Without loss of generality, assume that $L_{c}$ has one more link component than $L_{a}$ and $L_{b}$. This is equivalent to that the two endpoints on the left hand side of each tangle in Figure 1.2 are connected to each other, and the two endpoints on the right hand side are connected to each other. Then, one can find a weakly admissible Heegaard datum with Heegaard diagram $(\Sigma,\boldsymbol{\alpha},S(\boldsymbol{\beta}_{a}),S(\boldsymbol{\beta}_{b}),S(\boldsymbol{\beta}_{c}),\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z},G)$ such that $(\Sigma,S(\boldsymbol{\beta}_{a}),S(\boldsymbol{\beta}_{b}),S(\boldsymbol{\beta}_{c}),\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z},G)$ is a stabilization of Figure 4.1 and the three triple diagrams $(\Sigma,\boldsymbol{\alpha},S(\boldsymbol{\beta}_{a}),S(\boldsymbol{\beta}_{b}),\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\sqcup\{\partial_{+}G\})$, $(\Sigma,\boldsymbol{\alpha},S(\boldsymbol{\beta}_{b}),S(\boldsymbol{\beta}_{c}),\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\sqcup\{\partial_{+}G\})$, and $(\Sigma,\boldsymbol{\alpha},S(\boldsymbol{\beta}_{c}),S(\boldsymbol{\beta}_{a}),\boldsymbol{v}\sqcup\boldsymbol{w}\sqcup\boldsymbol{z}\sqcup\{\partial_{+}G\})$ are subordinate to the corresponding bands.
Consider the cycles $S(\boldsymbol{\rho})\in\boldsymbol{CF}^{-}(S(\boldsymbol{\beta}_{b}),S(\boldsymbol{\beta}_{c}^{E}))$, $S(\boldsymbol{\sigma})\in\boldsymbol{CF}^{-}(S(\boldsymbol{\beta}_{c}^{E}),S(\boldsymbol{\beta}_{a}))$, and $S(\tau)\in\boldsymbol{CF}^{-}(S(\boldsymbol{\beta}_{a}),S(\boldsymbol{\beta}_{b}))$, that are used to define the band maps. Let $\underline{S(\boldsymbol{\beta}_{ab})}:=S(\boldsymbol{\beta}_{a})\xrightarrow{S(\tau)}S(\boldsymbol{\beta}_{b})$ and let us consider the maps $\underline{S(\boldsymbol{\rho})}:\underline{S(\boldsymbol{\beta}_{ab})}\to S(\boldsymbol{\beta}_{c}^{E})$ and $\underline{S(\boldsymbol{\sigma})}:S(\boldsymbol{\beta}_{c}^{E})\to\underline{S(\boldsymbol{\beta}_{ab})}$ given by $S(\boldsymbol{\rho})$ and $S(\boldsymbol{\sigma})$, respectively. By Proposition 2.39, $\underline{S(\boldsymbol{\rho})}$ and $\underline{S(\boldsymbol{\sigma})}$ are cycles and $\mu_{2}(\underline{S(\boldsymbol{\rho})},\underline{S(\boldsymbol{\sigma})^{\prime}})$ and $\mu_{2}(\underline{S(\boldsymbol{\sigma})},\underline{S(\boldsymbol{\rho})^{\prime}})$ are $\Theta^{+}$ in a pinched almost complex structure. Consider the images $\underline{f}$ and $\underline{g}$ of $\underline{S(\boldsymbol{\rho})}$ and $\underline{S(\boldsymbol{\sigma})}$, respectively, under the $A_{\infty}$-functor given by changing the almost complex structures. Note that the $S(\boldsymbol{\beta}_{b})\to S(\boldsymbol{\beta}_{c}^{E})$ component of $\underline{f}$ is still $S(\boldsymbol{\rho})$ [^45] This is all we need, but actually there is no $S(\boldsymbol{\beta}_{a})\to S(\boldsymbol{\beta}_{c}^{E})$ component because of grading reasons. and that the $S(\boldsymbol{\beta}_{c}^{E})\to S(\boldsymbol{\beta}_{a})$ component of $\underline{g}$ is still $S(\boldsymbol{\sigma})$.
The following maps are quasi-isomorphisms by Lemma 2.34:
$\mu_{2}(-,\mu_{2}(\underline{f},\underline{g^{\prime}})):\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{S(\boldsymbol{\beta}_{ab})})\to\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{S(\boldsymbol{\beta}_{ab})^{\prime}})$
$\mu_{2}(-,\mu_{2}(\underline{g^{\prime}},\underline{f})):\boldsymbol{CF}^{-}(\boldsymbol{\alpha},S(\boldsymbol{\beta}_{c}^{E})^{\prime})\to\boldsymbol{CF}^{-}(\boldsymbol{\alpha},S(\boldsymbol{\beta}_{c}^{E})).$
Hence,
$\mu_{2}(-,\underline{f}):\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{S(\boldsymbol{\beta}_{ab})})\to\boldsymbol{CF}^{-}(\boldsymbol{\alpha},S(\boldsymbol{\beta}_{c}^{E}))$
is a quasi-isomorphism. Similarly, $\mu_{2}(-,\underline{g})$ is also a quasi-isomorphism. Hence, the triangles in Theorem 4.2 are exact at $L_{b}$ and $L_{a}$ by Lemma 2.30.
That the triangles are exact at $L_{c}$ follows similarly, using the following: since $\mu_{2}(\underline{\boldsymbol{\rho}},\underline{\boldsymbol{\sigma}^{\prime}})$ and $\mu_{2}(\underline{\boldsymbol{\sigma}},\underline{\boldsymbol{\rho}^{\prime}})$ are $\Theta^{+}$, if we consider the twisted complex $\underline{\boldsymbol{\beta}_{bc}}:=\boldsymbol{\beta}_{b}\xrightarrow{\boldsymbol{\rho}}\boldsymbol{\beta}_{c}^{E}$ and consider the maps $\underline{\tau}:\boldsymbol{\beta}_{a}\to\underline{\boldsymbol{\beta}_{bc}}$ and $\underline{\boldsymbol{\sigma}}:\underline{\boldsymbol{\beta}_{bc}}\to\boldsymbol{\beta}_{a}$ given by $\tau$ and $\boldsymbol{\sigma}$, respectively, then they are cycles and $\mu_{2}(\underline{\tau},\underline{\boldsymbol{\sigma}^{\prime}})$ and $\mu_{2}(\underline{\boldsymbol{\sigma}},\underline{\tau^{\prime}})$ are also $\Theta^{+}$. ∎
Remark 4.4.
Recall from Subsubsection 3.3.4 that we can define $HFL^{\prime}{}^{-}(Y,L;c)$ and band maps for $c\in H^{2}(Y)$. The above proof works for $HFL^{\prime}{}^{-}$ as well, if we work with $c$-strongly admissible diagrams (Subsubsection 2.10.1).
[…]
These intermediate objects have nice topological interpretations. Indeed, we get the following theorem as a corollary.
Theorem 4.5.
Let $K\cup L$ be a balled link in a closed, oriented three-manifold $Y$, where $K$ is an unmarked knot. Given any framing $\lambda$ of $K$, we have exact triangles
$\cdots\to\boldsymbol{HFL^{\prime}}^{-}(Y,K\cup L)\to\boldsymbol{HFL^{\prime}}^{-}(Y_{\lambda}(K),L)\to\boldsymbol{HFL^{\prime}}^{-}(Y_{\lambda+2\mu}(K),L)\to\boldsymbol{HFL^{\prime}}^{-}(Y,K\cup L)\to\cdots$
$\cdots\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,K\cup L)\to\boldsymbol{HFL^{\prime}}^{\infty}(Y_{\lambda}(K),L)\to\boldsymbol{HFL^{\prime}}^{\infty}(Y_{\lambda+2\mu}(K),L)\to\boldsymbol{HFL^{\prime}}^{\infty}(Y,K\cup L)\to\cdots$
$\cdots\to\widehat{HFL^{\prime}}(Y,K\cup L)\to\widehat{HFL^{\prime}}(Y_{\lambda}(K),L)\to\widehat{HFL^{\prime}}(Y_{\lambda+2\mu}(K),L)\to\widehat{HFL^{\prime}}(Y,K\cup L)\to\cdots,$
where $\mu$ is the meridian of $K$. These maps come from Heegaard triple diagrams, and are $\mathbb{F}\llbracket U_{1}^{1/2},\cdots,U_{n}^{1/2}\rrbracket$-linear where $U_{1}^{1/2},\cdots,U_{n}^{1/2}$ are the variables that correspond to the components of $L$.
Proof.
This reduces to the local computation carried out in Section 7, using an argument similar to the above proof of Theorem 4.2 assuming Theorem 4.3. ∎
We recover Theorem 1.3 by letting $L$ be the empty link with one marked, free baseball. In this case, we could have defined $\widehat{CFL^{\prime}}$ of $K$ as $\boldsymbol{CFL}^{\prime}{}^{-}/U$, as in Remark 3.5.
Remark 4.6.
Theorem 4.5 does not work for $HFL^{\prime}{}^{-}$ [^46] Indeed, we are not in the setting of Subsection 2.10., just like Ozsváth and Szabó’s surgery exact triangles [OS04b, Section 9]. This is because, morally speaking, the maps correspond to cobordism maps where we sum over infinitely many different ${\rm Spin}^{c}$-structures.
Remark 4.7.
It is possible to interpret the above maps as the composition of a link cobordism map ([Zem19], [JZ24, Theorem 9.3]) and an algebraic map as in Definition 3.21. The underlying four-manifolds are the same as those for Ozsváth and Szabó’s $2$-surgery exact triangle: they are given by attaching a $2$-handle to $Y_{\lambda+2\mu}(K)\times I$, $Y\times I$, or to the four-manifold obtained by attaching a one-handle to $(Y_{\lambda}(K)\sqcup\mathbb{RP}^{3})\times I$ that connects the two components. For the last case, we have to choose an element of $\boldsymbol{HF}^{-}(\mathbb{RP}^{3})$; the generator we choose lies in a different ${\rm Spin}^{c}$-structure from that for Ozsváth and Szabó’s $2$-surgery exact triangle.
5. The plan for proving Theorem 4.3
We will first establish three pairs of maps which are morally quasi-inverses of each other, and we will then use this to prove Theorem 4.3 in Section 8. We consider the giant weakly admissible, multi-Heegaard diagram given by the three diagrams in Figure 5.1, where the underlying pointed surface is $(\mathbb{T}^{2},\{v,w,z\})$ and the attaching circles are $\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}},\boldsymbol{\beta}_{c}$, and their standard translates. Let the coefficient ring be $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$ and assign the weight $U_{1}$ to $v$, and $U_{2}^{1/2}$ to $w$ and $z$. Note that $\boldsymbol{\beta}_{\infty}=\boldsymbol{\beta}_{c}$. The attaching curve $\boldsymbol{\beta}_{\infty}$ has a nontrivial $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$-local system $E$, given by the oriented arc $G$. Also note that it is not necessary to wind any of the curves to put all of the figures into one weakly admissible diagram.
Figure 5.1. The genus $1$ diagrams for Claims 5.1, 5.2, and 5.3.
Claim 5.1.
[…]
Proof.
We show this in Section 6: we define maps $\underline{\zeta}$, $\underline{\xi}$ in a destabilized diagram; $\underline{S(\zeta)}$ and $\underline{S(\xi)}$ are the wanted morally quasi-inverses. ∎
Claim 5.2.
[…]
${{\boldsymbol{\beta}_{0}}}$${{\boldsymbol{\beta}_{2}}}$${{\widetilde{\boldsymbol{\beta}_{0}}}}$${{\widetilde{\boldsymbol{\beta}_{2}}}}$${{\widetilde{\boldsymbol{\beta}_{0}}}}$${{\widetilde{\boldsymbol{\beta}_{2}}}}$${{\boldsymbol{\beta}_{0}}}$${{\boldsymbol{\beta}_{2}}}$$\scriptstyle{\theta}$$\scriptstyle{\Theta_{0}^{+}}$$\scriptstyle{\Theta_{2}^{+}}$$\scriptstyle{\widetilde{\theta}}$$\scriptstyle{\Theta_{0}^{+}}$$\scriptstyle{\Theta_{2}^{+}}$$\scriptstyle{\widetilde{\theta}}$$\scriptstyle{\theta}$
Proof.
Let us show that the dashed maps are cycles. First, the maps $\theta,\widetilde{\theta},\Theta_{0}^{+},\Theta_{2}^{+}$ are cycles, and note that the $\boldsymbol{\beta}_{i}$ and $\widetilde{\boldsymbol{\beta}_{i}}$ are related by handleslides and $\Theta_{i}^{+}$ is the top grading generator. Since $\mu_{2}(\Theta_{0}^{+},\widetilde{\theta})=\mu_{2}(\theta,\Theta_{2}^{+})$ on homology, and they must be equal on the chain level since they are the in the top grading ($(\boldsymbol{\beta}_{0},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}})$ and $(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\beta}_{2}})$ are $\mathbb{Z}$-gradable). Similarly, the dashed map on the right is a cycle. Their compositions satisfy the conditions of Lemma 2.34, with respect to the filtrations $\boldsymbol{\beta}_{2}\subset\{\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2}\}$ and $\widetilde{\boldsymbol{\beta}_{2}}\subset\{\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}}\}$. ∎
Claim 5.3.
The mapping cones $\left(\widetilde{\boldsymbol{\beta}_{0}}\xrightarrow{\widetilde{\theta}}\widetilde{\boldsymbol{\beta}_{2}}\right)$ and $\boldsymbol{\beta}_{\infty}^{E}$ are morally quasi-isomorphic over $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$, if the almost complex structure is pinched along the free stabilization region.
Proof.
We show this in Section 7: we define maps $e_{0}\underline{\xi}$, $e_{1}^{\ast}\underline{\zeta}$ in a destabilized diagram (the names are unfortunately the same as the maps for Claim 5.1); $\underline{S(e_{0}\xi)}$ and $\underline{S(e_{1}^{\ast}\zeta)}$ are the wanted morally quasi-inverses. ∎
Remark 5.4.
Claims 5.2 and 5.3 hold over the ring $\mathbb{F}\llbracket V,W,Z\rrbracket$, where we assign weights $V,W,Z$ to $v,w,z$, respectively. However, we need to assign the same weights to $w$ and $z$ for Claim 5.1.
Remark 5.5.
For Theorem 4.5, we only need to additionally check that $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)=0$ in addition to Claim 5.3.
Remark 5.6.
In fact, the above three pairs of twisted complexes are morally homotopy equivalent. For Claim 5.1, $u_{2}^{-1}\underline{S(\zeta)}$ and $\underline{S(\xi)}$, where $u_{2}=\sum_{n=1}^{\infty}U_{2}^{n^{2}-n}$, are morally homotopy inverses. For Claim 5.2, the given maps are morally homotopy inverses: one way to show this is to check that $(\mathbb{T}^{2},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}},\boldsymbol{\beta}_{0}^{\prime},\boldsymbol{\beta}_{2}^{\prime},\{v,w,z\})$ is $\mathbb{Z}$-gradable. For Claim 5.3, $\underline{S(e_{0}\xi)}$ and $u_{1}^{-1}\underline{S(e_{1}^{\ast}\zeta)}$, where $u_{1}=\sum_{n=1}^{\infty}U_{1}^{n^{2}-n}$, are morally homotopy inverses.
We believe that finishing off the proof from Claims 5.1, 5.2, and 5.3 is standard; we spell out one way to prove Theorem 4.3 in Section 8.
6. Local computation for Claim 5.1
Figure 6.1. Heegaard diagrams $(\mathbb{T}^{2},\beta_{a},\beta_{b},\beta_{0},\beta_{2},\beta_{a}^{\prime},\beta_{b}^{\prime},\{w,z\})$ and $(\mathbb{T}^{2},\beta_{0},\beta_{2},\beta_{a},\beta_{b},\beta_{0}^{\prime},\beta_{2}^{\prime},\{w,z\})$
[…]
${{\beta_{a}}}$${{\beta_{b}}}$${{\beta_{0}}}$${{\beta_{2}}}$${{\beta_{0}}}$${{\beta_{2}}}$${{\beta_{a}}}$${{\beta_{b}}}$$\scriptstyle{\tau}$$\scriptstyle{\zeta}$$\scriptstyle{\theta}$$\scriptstyle{\xi}$$\scriptstyle{\theta}$$\scriptstyle{\tau}$
We claim that they are cycles and that $\mu_{2}(\underline{\zeta},\underline{\xi^{\prime}})$ and $\mu_{2}(\underline{\xi},\underline{\zeta^{\prime}})$ are $\Theta^{+}$ modulo $U^{1/2}$. Recall that for instance, for $\mu_{2}(\underline{\zeta},\underline{\xi^{\prime}})$, this means that the composition of the vertical maps on the left hand side is the right hand side modulo $U^{1/2}$.
${{\beta_{a}}}$${{\beta_{b}}}$${{\beta_{a}}}$${{\beta_{b}}}$${{\beta_{0}}}$${{\beta_{2}}}$${{\beta_{a}^{\prime}}}$${{\beta_{b}^{\prime}}}$${{\beta_{a}^{\prime}}}$${{\beta_{b}^{\prime}}}$$\scriptstyle{\tau}$$\scriptstyle{\zeta}$$\scriptstyle{\tau}$$\scriptstyle{\Theta_{a}^{+}}$$\scriptstyle{\Theta_{b}^{+}}$$\scriptstyle{\theta}$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{\tau^{\prime}}$$\scriptstyle{\tau^{\prime}}$
Showing the above is equivalent to the following:
[…]
  • $\mu_{1}(\underline{\zeta})=0$: $\mu_{1}(\zeta)=0$, $\mu_{2}(\tau,\zeta)=0$, $\mu_{2}(\zeta,\theta)=0$, $\mu_{3}(\tau,\zeta,\theta)=0$
  • $\mu_{1}(\underline{\xi})=0$: $\mu_{1}(\xi)=0$, $\mu_{2}(\theta,\xi)=0$, $\mu_{2}(\xi,\tau)=0$, $\mu_{3}(\theta,\xi,\tau)=0$
  • $\mu_{2}(\underline{\zeta},\underline{\xi^{\prime}})$: $\mu_{3}(\zeta,\theta,\xi^{\prime})=0$, $\mu_{4}(\tau,\zeta,\theta,\xi^{\prime})=\Theta_{a}^{+}$, $\mu_{4}(\zeta,\theta,\xi^{\prime},\tau^{\prime})=\Theta_{b}^{+}$, $\mu_{5}(\tau,\zeta,\theta,\xi^{\prime},\tau^{\prime})=0$ modulo $U^{1/2}$
  • $\mu_{2}(\underline{\xi},\underline{\zeta^{\prime}})$: $\mu_{3}(\xi,\tau,\zeta^{\prime})=0$, $\mu_{4}(\theta,\xi,\tau,\zeta^{\prime})=\Theta_{0}^{+}$, $\mu_{4}(\xi,\tau,\zeta^{\prime},\theta^{\prime})=\Theta_{2}^{+}$, $\mu_{5}(\theta,\xi,\tau,\zeta^{\prime},\theta^{\prime})=0$ modulo $U^{1/2}$
Remark 6.1.
In fact, we can explicitly compute $\mu_{2}(\underline{\zeta},\underline{\xi^{\prime}})$ and $\mu_{2}(\underline{\xi},\underline{\zeta^{\prime}})$. All the $\mu_{3}$’s and $\mu_{5}$’s vanish, and the $\mu_{4}$’s are $u\Theta^{+}$ for the unit $u=\sum_{n=1}^{\infty}U^{n^{2}-n}$.
Some of these are easy to verify. Indeed, the $\mu_{1}$’s, $\mu_{2}(\tau,\zeta)$, and $\mu_{2}(\xi,\tau)$ vanish because there are no relevant bigons and triangles. The hexagon count $\mu_{5}$ vanishes because there are no relevant hexagons that do not contain any basepoints.
To verify the rest, we work in the universal cover of $\mathbb{T}^{2}$. It will be useful to have names for some right triangles.
Definition 6.2.
Let $\beta_{\infty}$ be the straight, vertical line that is in the middle of $\beta_{a}$ and $\beta_{b}$. We abuse notation and label the intersection points as follows: $\beta_{\infty}\cap\beta_{0}=\{\zeta\}$, $\beta_{2}\cap\beta_{\infty}=\{\xi\}$. Also, let $\beta_{0}\cap\beta_{2}=\{\theta_{0},\theta_{1}\}$, where $\theta_{1}=\theta$. See Figure 6.2.
  • Let triangles $T_{n}^{+}$ and $S_{n}^{+}$, respectively, for $n\in\mathbb{Z}_{\geq 1}$ be the $n$th right-side up triangle that has vertices $\theta_{1},\xi,\zeta$ and $\theta_{0},\xi,\zeta$, respectively.
  • Let triangles $T_{n}^{-}$ and $S_{n}^{-}$, respectively, for $n\in\mathbb{Z}_{\geq 1}$ be the $n$th upside down triangle that has vertices $\theta_{1},\xi,\zeta$ and $\theta_{0},\xi,\zeta$, respectively.
Figure 6.2. Right triangles and their names
6.1. The $\mu_{2}$’s vanish
For $\mu_{2}(\zeta,\theta)$ and $\mu_{2}(\theta,\xi)$, there are two families of triangles that cancel each other. These are related by “rotation by $\pi$.” In fact, all of the quadrilaterals (and hexagons) that cancel in pairs are also related by “rotation by $\pi$” in some sense.
We consider $\mu_{2}(\zeta,\theta)$. The two smallest triangles are drawn in the left hand side of Figure 6.3. These have one basepoint, and hence have weight $U^{1/2}$. The next smallest triangles are drawn in Figure 6.5, and they have weight $U^{5/2}$, as they contain five basepoints.
We view these triangles as small perturbations of $T_{n}^{\pm}$. (We think of $\beta_{a}$ and $\beta_{b}$ as being close to each other: the region between them is “small.”) Such triangle $T_{n}^{+}$ or $T_{n}^{-}$ uniquely determines the corresponding triangle that contributes to $\mu_{2}(\zeta,\theta)$. For each $n$, there are two of them, one from $T_{n}^{+}$ and one from $T_{n}^{-}$, and they have weight $U^{n^{2}-n+\frac{1}{2}}$. Hence they all cancel.
Figure 6.3. Left: two small triangles with weight $U^{1/2}$; Right: Two small quadrilaterals with weight $1$
6.2. The $\mu_{3}$’s vanish
We will write $\zeta$ instead of $\zeta^{\prime}$ and $\beta_{0}$ instead of $\beta_{0}^{\prime}$, etc. since distinguishing them in this subsection is meaningless.
Let us show that $\mu_{3}(\tau,\zeta,\theta)$, $\mu_{3}(\zeta,\theta,\xi)$, and $\mu_{3}(\theta,\xi,\tau)$ vanish. Any quadrilateral that contributes in these quadrilateral count has vertices $\xi,\tau,\zeta,\theta$. There are two families of quadrilaterals with vertices $\xi,\tau,\zeta,\theta$ that cancel each other. The two smallest quadrilaterals are drawn in the right hand side of Figure 6.3. These do not have any basepoints, and hence have weight $1$. The next smallest are drawn in Figure 6.6, and they have weight $U^{2}$.
We view these quadrilaterals as small perturbations of $T_{n}^{\pm}$’s as in Subsection 6.1. Given a triangle $T_{n}^{\pm}$, specifying such a quadrilateral is equivalent to choosing a lift of $\tau$ as vertices of the quadrilateral. For each of $T_{n}^{+}$ and $T_{n}^{-}$, there is $12n-$ 1 such choices, and there are exactly $2n^{2}-2n$ basepoints for each such quadrilateral. Hence they all cancel.
Computing $\mu_{3}(\xi,\tau,\zeta)$ is slightly more complicated since the quadrilaterals do not need to have $\theta$ as a vertex, but if we work modulo $U^{1/2}$, then they all have $\theta$ as a vertex, and there are exactly two of them, which cancel.
6.3. The $\mu_{4}$’s are the identity modulo $U^{1/2}$
For the $\mu_{4}$’s, in all the cases, the Maslov index $-2$ pentagons are small perturbations of the above quadrilaterals (not every quadrilateral has a small perturbation that contributes to $\mu_{4}$). In each case, there is exactly one pentagon with vertices $\tau,\zeta,\theta,\xi$ and $\Theta_{i}^{+}$ ($i=a,b,0,2$) that does not contain any basepoints. This pentagon is a small perturbation of the top quadrilateral in the right hand side of Figure 6.3. Two of these are drawn in Figure 6.4.
Figure 6.4. Relevant pentagons for $\mu_{4}(\tau,\zeta,\theta,\xi^{\prime})$ and $\mu_{4}(\theta,\xi,\tau,\zeta^{\prime})$ modulo $U^{1/2}$. These are small perturbations of the top quadrilateral of the right hand side of Figure 6.3.
Figure 6.5. Two triangles with weight $U^{5/2}$
Figure 6.6. Six quadrilaterals with weight $U^{2}$. The top quadrilaterals are small perturbations of $T_{2}^{+}$ and the bottom ones are small perturbations of $T_{2}^{-}.$
7. Local computation for Claim 5.3
Figure 7.1. The genus $1$ Heegaard diagram $(\mathbb{T}^{2},\beta_{0},\beta_{2},\beta_{\infty},\beta_{0}^{\prime},\beta_{2}^{\prime},\beta_{\infty}^{\prime},v)$
Consider the Heegaard diagram Figure 7.1, and work over the ring $\mathbb{F}\llbracket U\rrbracket$ and assign the weight $U$ to $v$. The circle $\beta_{\infty}$ has a nontrivial local system $E=e_{0}\mathbb{F}\llbracket U\rrbracket\oplus e_{1}\mathbb{F}\llbracket U\rrbracket$ whose monodromy is given by the oriented arc $G$.
[…]
${{\beta_{0}}}$${{\beta_{2}}}$${{\beta_{\infty}^{E}}}$${{\beta_{\infty}^{E}}}$${{\beta_{0}}}$${{\beta_{2}}}$$\scriptstyle{\theta}$$\scriptstyle{e_{0}\xi}$$\scriptstyle{e_{1}^{\ast}\zeta}$$\scriptstyle{\theta}$
As in Section 6, we will show that they are cycles and that $\mu_{2}(e_{0}\underline{\xi},e_{1}^{\ast}\underline{\zeta^{\prime}})$ and $\mu_{2}(e_{1}^{\ast}\underline{\zeta},e_{0}\underline{\xi^{\prime}})$ are $\Theta^{+}$ modulo $U$.
As before, this is equivalent to showing the following.
  • $\mu_{1}(\theta)=0$
  • $\mu_{1}(e_{0}\xi)=0$, $\mu_{2}(\theta,e_{0}\xi)=0$
  • $\mu_{1}(e_{1}^{\ast}\zeta)=0$, $\mu_{2}(e_{1}^{\ast}\zeta,\theta)=0$
  • $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)=0$, $\mu_{3}(\theta,e_{0}\xi,e_{1}^{\ast}\zeta^{\prime})=\Theta_{0}^{+}$, $\mu_{3}(e_{0}\xi,e_{1}^{\ast}\zeta^{\prime},\theta^{\prime})=\Theta_{2}^{+}$, $\mu_{4}(\theta,e_{0}\xi,e_{1}^{\ast}\zeta^{\prime},\theta^{\prime})=0$ modulo $U$
  • $\mu_{3}(e_{1}^{\ast}\zeta,\theta,e_{0}\xi^{\prime})={\rm Id}_{E}\Theta_{\infty}^{+}$ modulo $U$
We also need to check $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)=0$ for Theorem 4.5.
Remark 7.1.
We can explicitly compute $\mu_{2}(e_{0}\underline{\xi},e_{1}^{\ast}\underline{\zeta^{\prime}})$ and $\mu_{2}(e_{1}^{\ast}\underline{\zeta},e_{0}\underline{\xi^{\prime}})$. Indeed, we get $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)=\mu_{4}(\theta,e_{0}\xi,e_{1}^{\ast}\zeta^{\prime},\theta^{\prime})=0$ and the $\mu_{3}$’s are $u\Theta^{+}$ where the $u$ is the same as in Section 6: $u=\sum_{n=1}^{\infty}U^{n^{2}-n}$.
Similarly, the $\mu_{1}$’s vanish because there are no bigons, and the $\mu_{4}$ vanish because there are no pentagons with Maslov index $-2$.
7.1. The $\mu_{2}$’s vanish
We evaluate the triangle counting maps. Recall Definition 6.2.
We compute $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)$, which is the most complicated case, since it is not a priori obvious that the triangles $S_{n}^{+}$ and $S_{n}^{-}$ do not contribute. Recall that $\phi=e_{1}e_{0}^{\ast}+Ue_{0}e_{1}^{\ast}$.
  • The triangle $T_{n}^{+}$ contributes
    $U^{(n-1)^{2}}e_{1}^{\ast}\phi^{2n-1}e_{0}\theta=U^{n^{2}-n}\theta;$
  • The triangle $T_{n}^{-}$ contributes
    $U^{(n-1)^{2}}e_{1}^{\ast}\left(U\phi^{-1}\right)^{2n-1}e_{0}\theta=U^{n^{2}-n}\theta;$
    (note that $U\phi^{-1}=\phi$)
  • The triangle $S_{n}^{\pm}$ contributes some multiple of
    $e_{1}^{\ast}\phi^{2k}e_{0}\theta_{0}=0;$
and so they all cancel and we get $\mu_{2}(e_{0}\xi,e_{1}^{\ast}\zeta)=0$.
The rest vanish as well by direct computation.
7.2. The $\mu_{3}$’s are the identity
Figure 7.2. The two quadrilaterals highlighted in yellow are the two quadrilaterals that contribute modulo $U$, and they add up to ${\rm Id}_{E}\Theta_{\infty}^{+}$.
We compute $\mu_{3}(e_{1}^{\ast}\zeta,\theta,e_{0}\xi^{\prime})$, which is the most complicated case. If we work modulo $U$, then the two small quadrilaterals highlighted in yellow in Figure 7.2 are the only quadrilaterals that contribute. The top one outputs $e_{0}e_{1}^{\ast}\phi\Theta_{\infty}^{+}=e_{0}e_{0}^{\ast}\Theta_{\infty}^{+}$ and the bottom one outputs $U\phi^{-1}e_{0}e_{1}^{\ast}\Theta_{\infty}^{+}=e_{1}e_{1}^{\ast}\Theta_{\infty}^{+}$. Hence, they add up to ${\rm Id}_{E}\Theta_{\infty}^{+}.$
For the other $\mu_{3}$’s, there is exactly one quadrilateral that contributes, which gives $\Theta_{0}^{+}$ or $\Theta_{2}^{+}$.
8. Completing the proof of Theorem 4.3
8.1. A homological $\mathbb{Z}$-grading
A useful observation that we will repeatedly use is that the following Heegaard datum is $\mathbb{Z}$-gradable. Compare [MO08, Section 3] and [BL12, Section 5.2].
Proposition 8.1.
Consider the multi-Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c},\{v,w,z\},G)$ given in Figure 4.1, and add any number of standard translates of $\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c}$. Work over $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$ and assign the weight $U_{1}$ to $v$, and $U_{2}^{1/2}$ to $w$ and $z$. Then, this Heegaard datum is $\mathbb{Z}$-gradable.
Proof.
This is a corollary of Proposition 2.65: the $\boldsymbol{\beta}_{i}$’s are homologous, and the intersection points are all in torsion ${\rm Spin}^{c}$-structures. ∎
We let the top $\mathbb{Z}$-grading elements of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})$ and $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c}^{E})$ have $\mathbb{Z}$-grading 00. Then, the top $\mathbb{Z}$-grading elements of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{c}^{E})$, $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\boldsymbol{\beta}_{a}^{\prime})$, $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\boldsymbol{\beta}_{b}^{\prime})$, $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{a}^{\prime})$ have $\mathbb{Z}$-grading 00, $-1$, $-1$, $-1$, respectively.
[…]
8.2. An Alexander $\mathbb{Z}/2$-splitting
It turns out that the following Heegaard datum with attaching curves $\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}},\boldsymbol{\beta}_{c}$ and their standard translates is $\mathbb{Z}/2$-splittable.
Figure 8.1. A cornerless two-chain ${\mathcal{}D}$ whose boundary has exactly one circle with slope $2$.
Proposition 8.2.
Consider the multi-Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}},\boldsymbol{\beta}_{c}^{E},\{v,w,z\},G)$ defined in Section 5, and add any number of standard translates of these attaching curves. Work over $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$ and assign the weight $U_{1}$ to $v$, and $U_{2}^{1/2}$ to $w$ and $z$. Then, this Heegaard datum is $\mathbb{Z}/2$-splittable.
Proof.
We check that $P({\mathcal{}D})$ is even for any cornerless two-chain ${\mathcal{}D}$. Lemma 2.62 and Corollary 2.63 are not directly applicable, but we can reduce the number of cases to check using similar ideas.
Perhaps the most nontrivial special case to check is for the cornerless two-chain in Figure 8.1, whose boundary has exactly one circle with slope $2$. We have $P=6$, which is even. Given any cornerless two-chain ${\mathcal{}D}$, using the above cornerless two-chain and small cornerless two-chains between two circles which are standard translates of each other, we can reduce it to the case where $\partial{\mathcal{}D}$ has no components with slope $2$. Hence, it also does not have any components with slope 00, and so its boundary only consists of vertical circles and homologically trivial circles. This case follows from an argument similar to that of Lemma 2.62: the two-chain is a sum of cornerless two-chains with two vertices, and we can check these directly. ∎
We let all the generators that we have considered in Section 5 have Alexander $\mathbb{Z}/2$-grading 00. This is indeed possible: there are two things to check. One is that we do not get a contradiction from the maps from Claims 5.2, and the other is that the $\Theta^{+}$’s have Alexander $\mathbb{Z}/2$-grading 00. Consider the intersection point $s\in\boldsymbol{\beta}_{2}\cap\widetilde{\boldsymbol{\beta}_{0}}$ that corresponds to $\theta$. Then, the former follows from that ${\rm gr}_{A}^{\mathbb{Z}/2}(\Theta_{0}^{+})={\rm gr}_{A}^{\mathbb{Z}/2}(\theta)+{\rm gr}_{A}^{\mathbb{Z}/2}(s)$ and ${\rm gr}_{A}^{\mathbb{Z}/2}(\widetilde{\theta})={\rm gr}_{A}^{\mathbb{Z}/2}(\Theta_{2}^{+})+{\rm gr}_{A}^{\mathbb{Z}/2}(s)$. The latter also follows from direct computations.
8.3. Another computation
Let us briefly study the top $\mathbb{Z}$-grading component of $\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\underline{\boldsymbol{\beta}_{ab}^{\prime}})$. The top $\mathbb{Z}$-grading (grading 00) generators are $\Theta_{a}^{+}$, $\Theta_{b}^{+}$, $x$, and $y$ in Figure 8.2. (The intersection points in the free-stabilizing region are not drawn.) The Alexander $\mathbb{Z}/2$-grading of $x$ is $1$, and the rest are 00. One can check that, among linear combinations of $\Theta_{a}^{+}$, $\Theta_{b}^{+}$, and $y$, the sum $\Theta_{a}^{+}+\Theta_{b}^{+}$ is the only nontrivial cycle, regardless of the almost complex structure.
Figure 8.2. A Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{a}^{\prime},\boldsymbol{\beta}_{b}^{\prime},z,w,v)$
8.4. Independence under variation of the almost complex structure
We show that the statement of Theorem 4.3 is insensitive of the almost complex structure. Let us first check that the maps are invariant. In all the cases, this will follow from that there are no relevant domains with the right Maslov index, which is easy to check, thanks to Proposition 8.1.
The generators $\tau\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b})$, ${\rm Id}_{E}\Theta^{+}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\boldsymbol{\beta}_{c}^{\prime}{}^{E})$, $\boldsymbol{\sigma}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c}^{E})$, and $\boldsymbol{\rho}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\boldsymbol{\beta}_{a})$ are invariant under changing the almost complex structure, since the chain map induced by changing the almost complex structure preserves both the $\mathbb{Z}$-grading and the Alexander $\mathbb{Z}/2$-grading, and the above generators are the unique nontrivial cycle in the corresponding $\mathbb{Z}\oplus\mathbb{Z}/2$-grading. Hence, in particular, $\underline{\boldsymbol{\beta}_{ab}}$ is preserved under changing the almost complex structure.
Since the top $\mathbb{Z}$-grading of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a}[-1],\boldsymbol{\beta}_{c}^{E})$ is $-1$ and $\underline{\boldsymbol{\sigma}}\in\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\boldsymbol{\beta}_{c}^{E})$ has $\mathbb{Z}$-grading 00, the elements $\underline{\boldsymbol{\sigma}}\in\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\boldsymbol{\beta}_{c}^{E})$ and $\underline{\boldsymbol{\rho}}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c}^{E},\underline{\boldsymbol{\beta}_{ab}})$ are preserved under changing the almost complex structure.
The maps $\Theta_{a}^{+}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a}[-1],\boldsymbol{\beta}_{a}^{\prime}[-1])$ and $\Theta_{b}^{+}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{b}^{\prime})$ are invariant as well. Since the top $\mathbb{Z}$-grading of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a}[-1],\boldsymbol{\beta}_{b})$ is $-1$, they are invariant as elements of $\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\underline{\boldsymbol{\beta}_{ab}^{\prime}})$. Hence, $\Theta^{+}=\underline{\Theta_{a}^{+}}+\underline{\Theta_{b}^{+}}$ is invariant.
Finally, if we know that $\mu_{2}(\underline{\boldsymbol{\sigma}},\underline{\boldsymbol{\rho}^{\prime}})=\Theta^{+}$ and $\mu_{2}(\underline{\boldsymbol{\rho}},\underline{\boldsymbol{\sigma}^{\prime}})=\Theta^{+}$ for some choice of almost complex structure, then in any other almost complex structure, we know that they are equal in homology, but since they are in the top $\mathbb{Z}$-grading, they must be equal.
8.5. Finishing off the proof
[…]
We claim that the top $\mathbb{Z}$-grading components of $\mu_{2}(\underline{\widetilde{\boldsymbol{\rho}}},\underline{\widetilde{\boldsymbol{\sigma}}^{\prime}}):\underline{\boldsymbol{\beta}_{ab}}\to\underline{\boldsymbol{\beta}_{ab}^{\prime}}$ and $\mu_{2}(\underline{\widetilde{\boldsymbol{\sigma}}},\underline{\widetilde{\boldsymbol{\rho}}^{\prime}}):\boldsymbol{\beta}_{c}^{E}\to\boldsymbol{\beta}_{c}^{\prime}{}^{E}$ are $\Theta^{+}$. Let $\boldsymbol{\beta}=\boldsymbol{\beta}_{a}^{\prime\prime}$ be another standard translate of $\boldsymbol{\beta}_{a}$. Then, $(\mathbb{T}^{2},\boldsymbol{\beta},\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{a}^{\prime},\boldsymbol{\beta}_{b}^{\prime},v,w,z)$ is weakly admissible, and both homologically $\mathbb{Z}$-gradable and Alexander $\mathbb{Z}/2$-splittable. Since
$\mu_{2}(-,\mu_{2}(\underline{\widetilde{\boldsymbol{\rho}}},\underline{\widetilde{\boldsymbol{\sigma}}^{\prime}})):\boldsymbol{CF}^{-}(\boldsymbol{\beta},\underline{\boldsymbol{\beta}_{ab}})\to\boldsymbol{CF}^{-}(\boldsymbol{\beta},\underline{\boldsymbol{\beta}_{ab}^{\prime}})$
is a quasi-isomorphism and $\boldsymbol{HF}^{-}(\boldsymbol{\beta},\underline{\boldsymbol{\beta}_{ab}})\neq 0$, it cannot strictly lower the $\mathbb{Z}$-grading. Since the top $\mathbb{Z}$-grading of $\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\underline{\boldsymbol{\beta}_{ab}^{\prime}})$ is 00, the $\mathbb{Z}$-grading 00 part of $\mu_{2}(\underline{\widetilde{\boldsymbol{\rho}}},\underline{\widetilde{\boldsymbol{\sigma}}^{\prime}})$ is nontrivial, and so it must be $\Theta^{+}=\underline{\Theta_{a}^{+}}+\underline{\Theta_{b}^{+}}$, as it is the unique cycle in $\mathbb{Z}$-grading 00 and Alexander $\mathbb{Z}/2$-grading 00. We similarly get that $\mu_{2}(\underline{\widetilde{\boldsymbol{\sigma}}},\underline{\widetilde{\boldsymbol{\rho}}^{\prime}})={\rm Id}_{E}\Theta_{c}^{+}$.
Now, we claim that $\underline{\boldsymbol{\rho}}$ and $\underline{\boldsymbol{\sigma}}$ are the top $\mathbb{Z}$-grading components of $\underline{\widetilde{\boldsymbol{\rho}}}$ and $\underline{\widetilde{\boldsymbol{\sigma}}}$, respectively. Since the top $\mathbb{Z}$-grading of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c}^{E})$ is 00 and the top $\mathbb{Z}$-grading of $\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{a}[-1],\boldsymbol{\beta}_{c}^{E})$ is $-1$, the top $\mathbb{Z}$-grading and Alexander $\mathbb{Z}/2$-grading 00 component of $\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\boldsymbol{\beta}_{c}^{E})$ is generated by $e_{0}\underline{bx}$ and $e_{0}\underline{ay}$, and $e_{0}\underline{bx}+e_{0}\underline{ay}=\underline{\boldsymbol{\rho}}$ is the unique nontrivial cycle. Since the top grading component of $\underline{\widetilde{\boldsymbol{\rho}}}$ is a cycle, it must be $\underline{\boldsymbol{\rho}}$. Similarly, $\underline{\boldsymbol{\sigma}}$ is the top grading component of $\underline{\widetilde{\boldsymbol{\sigma}}}$.
Since $\underline{\boldsymbol{\rho}}$ and $\underline{\boldsymbol{\sigma}}$ have $\mathbb{Z}$-grading 00 and the $\mathbb{Z}$-grading 00 part of $\mu_{2}(\underline{\widetilde{\boldsymbol{\rho}}},\underline{\widetilde{\boldsymbol{\sigma}}^{\prime}})$ and $\mu_{2}(\underline{\widetilde{\boldsymbol{\sigma}}},\underline{\widetilde{\boldsymbol{\rho}}^{\prime}})$ are both $\Theta^{+}$, we get that $\mu_{2}(\underline{\boldsymbol{\rho}},\underline{\boldsymbol{\sigma}^{\prime}})=\Theta^{+}\in\boldsymbol{CF}^{-}(\underline{\boldsymbol{\beta}_{ab}},\underline{\boldsymbol{\beta}_{ab}^{\prime}})$ and $\mu_{2}(\underline{\boldsymbol{\sigma}},\underline{\boldsymbol{\rho}})=\Theta^{+}\in\boldsymbol{CF}^{-}(\boldsymbol{\beta}_{c},\boldsymbol{\beta}_{c}^{\prime}{}^{E})$.
9. Computation for planar links
In this section, we compute the band maps for planar links. First, let us compute the unoriented link Floer homology of unlinks.
Proposition 9.1.
Let $L$ be a balled link in $S^{3}$, where the underlying link is an unlink. Let the link baseballs correspond to variables $U_{1},\cdots,U_{k}$. Then, the unoriented link Floer homology of the balled link $L$ is canonically isomorphic to
$\mathbb{F}\llbracket U_{1}^{1/2},\cdots,U_{k}^{1/2},U\rrbracket/(U_{1}-U,\cdots,U_{k}-U)$
over the coefficient ring, hence in particular over $\mathbb{F}\llbracket U_{1}^{1/2},\cdots,U_{k}^{1/2}\rrbracket$.
Proof.
Iterate Proposition 2.41. ∎
Note that as an $\mathbb{F}\llbracket U\rrbracket$-module, the homology is freely generated and has rank $2^{k}$.
Although this is tangential to the rest of this section, we compute non-orientable band maps between unlinks: they agree with Khovanov homology.
Theorem 9.2.
Non-orientable band maps between unlinks in $S^{3}$ are zero.
Proof.
[…]
$\boldsymbol{HFL}^{\prime}{}^{\infty}(S^{3},UL_{k})\to\boldsymbol{HFL}^{\prime}{}^{\infty}(S^{3},UL_{k})$
is zero. Since the map $\boldsymbol{HFL}^{\prime}{}^{-}(S^{3},UL_{k})\to\boldsymbol{HFL}^{\prime}{}^{\infty}(S^{3},UL_{k})$ is injective, the map on the minus version is also zero. ∎
The band maps for planar links in unoriented link Floer homology coincide with that in equivariant Khovanov homology given by the Lee deformation.
Theorem 9.3.
Let $B$ be a merge band from $L$ to $L^{\prime}$ where $L$ and $L^{\prime}$ are balled unlinks in $S^{3}$. Assume that the band $B$ and the underlying links of $L$ and $L^{\prime}$ (ignoring the baseballs) all lie in $S^{2}\subset S^{3}$. Without loss of generality, assume that the coefficient ring for $L$ is $\mathbb{F}\llbracket U_{1}^{1/2},U_{2}^{1/2},U_{3}^{k_{3}},\cdots,U_{n}^{k_{n}}\rrbracket$, that the band intersects the link components with link baseballs that correspond to $U_{1}$ and $U_{2}$, and that the baseball that corresponds to $U_{2}$ becomes a free baseball for the link $L^{\prime}$. Let $S=\mathbb{F}\llbracket U_{1},U_{2},U_{3}^{k_{3}},\cdots,U_{n}^{k_{n}},U\rrbracket/(U_{1}-U,\cdots,U_{k}-U)$, and consider $\boldsymbol{HFL^{\prime}}^{-}(S^{3},L)$ and $\boldsymbol{HFL^{\prime}}^{-}(S^{3},L^{\prime})$ as modules over $S$, freely generated by $1,U_{1}^{1/2},U_{2}^{1/2},U_{1}^{1/2}U_{2}^{1/2}$ and $1,U_{1}^{1/2}$, respectively.
  • The merge band map for $B:L\to L^{\prime}$ is the $S$-linear map given by
    $1\mapsto 1,\ U_{1}^{1/2}\mapsto U^{1/2},\ U_{2}^{1/2}\mapsto U^{1/2},\ U_{1}^{1/2}U_{2}^{1/2}\mapsto U.$
  • The split band map for the dual band $B^{\ast}:L^{\prime}\to L$ is the $S$-linear map given by
    $1\mapsto U_{1}^{1/2}+U_{2}^{1/2},\ U^{1/2}\mapsto U+U_{1}^{1/2}U_{2}^{1/2}.$
The corresponding statements for the hat and infinity versions follow immediately. The proof of Theorem 9.3 occupies the remainder of this section.
A Heegaard datum for split and merge band maps between planar links with $k$ and $k+1$ components can be obtained by successively free link stabilizing such a Heegaard datum for $k=1$. Hence, Theorem 9.3 follows from the computation for $k=1$, together with Propositions 2.39 and 2.41.
Consider the multi-Heegaard diagram in Figure 9.1, and recall the maps $\tau,\boldsymbol{\rho},\boldsymbol{\sigma}$ from Theorem 4.3. Work over $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$ and assign the weight $U_{1}$ to $v$, and $U_{2}^{1/2}$ to $w$ and $z$. Then, $\mu_{2}(-,\boldsymbol{\rho}):\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{c}^{E})$ and $\mu_{2}(-,\boldsymbol{\sigma}):\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{c}^{E})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{a})$ are the split and merge band maps, respectively, between planar links with one and two components.
Figure 9.1. Multi-Heegaard diagram for (merge and split) band maps between planar links
We will not use $\boldsymbol{\rho}$ and $\boldsymbol{\sigma}$ directly to compute the band maps. Proposition 2.65 also shows that the Heegaard datum with Heegaard diagram
$(\mathbb{T}^{2},\text{$\boldsymbol{\alpha}$,}\boldsymbol{\beta}_{a},\boldsymbol{\beta}_{b},\boldsymbol{\beta}_{c},\boldsymbol{\beta}_{a}^{\prime},\boldsymbol{\beta}_{b}^{\prime},\boldsymbol{\beta}_{c}^{\prime},\{v,w,z\},G)$
is $\mathbb{Z}$-gradable. Hence, the merge and split band maps are the top $\mathbb{Z}$-grading components of the compositions of
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{a}),$
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E}),$
respectively, which are induced by the three maps discussed in Section 5 (see Remark 5.6). Indeed, the band maps must be the top grading components since the merge and split band maps cannot be zero, because of the exact triangle (it is also possible to compute the gradings directly).
Remark 9.4.
As the band maps are $\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$-linear and homogeneous, the skein exact triangle narrows the maps down to two possibilities:
${\rm merge}:1\mapsto U_{1}^{1/2},\ {\rm split}:1\mapsto 1,\ U_{1}^{1/2}\mapsto 0;{\rm or}$
${\rm merge}:1\mapsto U_{1}^{1/2}+U_{2}^{1/2},\ {\rm split}:1\mapsto 1,\ U_{1}^{1/2}\mapsto U_{2}^{1/2}$
we will show that it is indeed the latter.
9.1. A local computation
Figure 9.2 is key to computing the band maps.
Figure 9.2. The key multi-Heegaard diagram for computing the band maps for planar links
Proposition 9.5.
[…]
$\boldsymbol{HF}^{-}(\alpha,\underline{\beta})=c^{-}\mathbb{F}\llbracket U^{1/2}\rrbracket\oplus d_{1}\mathbb{F}\llbracket U^{1/2}\rrbracket$
after identifying $c^{-}$ and $d_{1}$ with their homology classes. Furthermore, the maps $\underline{\xi}:\underline{\beta_{02}}\to\beta_{\infty}$ and $\underline{\zeta}:\beta_{\infty}\to\underline{\beta_{02}}$ induced by $\xi$ and $\zeta$, respectively, are cycles and we have
  • $\mu_{2}(-,\underline{\xi}):\boldsymbol{HF}^{-}(\alpha,\underline{\beta_{02}})\to\boldsymbol{HF}^{-}(\alpha,\beta_{\infty}):c^{-}\mapsto uU^{1/2}x,\ d_{1}\mapsto x$;
  • $\mu_{2}(-,\underline{\zeta}):\boldsymbol{HF}^{-}(\alpha,\beta_{\infty})\to\boldsymbol{HF}^{-}(\alpha,\underline{\beta_{02}}):x\mapsto c^{-}+uU^{1/2}d_{1},$
where $u=\sum_{n=1}^{\infty}U^{n^{2}-n}$ is the unit from Sections 6 and 7.
Proof.
This computation is similar in spirit to the computations carried out in Sections 6 and 7. ∎
Remark 9.6.
Similarly to Sections 6 and 7, we do not need the full computation of the $\mu_{2}$’s in Proposition 9.5, as we only care about the top $\mathbb{Z}$-grading components of the compositions $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{c}^{E})$ and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{c}^{E})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{a})$. For instance, computing them modulo $U$ is sufficient for this paper.
9.2. Split bands
Figure 9.3. Handleslide $\boldsymbol{\alpha}$ to $\widetilde{\boldsymbol{\alpha}}$ to compute in the pinched almost complex structure
To compute $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E})$, we consider another curve $\widetilde{\boldsymbol{\alpha}}$ as in Figure 9.3 and compute the maps paired with $\widetilde{\boldsymbol{\alpha}}$ instead. Then, we use the following commutative diagram
[…]
We can compute every chain complex that arises as vertices in the above diagram, and we can compute all the maps except $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E})$.
9.2.1. The homology groups
Let us first study the relevant homology groups. We remark that the $A_{\infty}$-category that consists of $\boldsymbol{\alpha},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\alpha}},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}}$ and their standard translates (in fact, $\boldsymbol{\alpha}$, $\widetilde{\boldsymbol{\alpha}}$ is a standard translate of $\boldsymbol{\beta}_{0}$, $\widetilde{\boldsymbol{\beta}_{0}}$, respectively) is both $\mathbb{Z}$-gradable and $\mathbb{Z}/2$-splittable.
The Heegaard diagram $(\mathbb{T}^{2},\boldsymbol{\alpha},\boldsymbol{\beta}_{0},\{v,w,z\})$ is a free-stabilization of the Heegaard diagram $(\mathbb{T}^{2},\alpha,\beta_{0},\{w,z\})$ from Figure 9.2. Hence we identify (as $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$-modules, where $U_{1}$ acts as $U_{2}$, and so equivalently as $\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$-modules)
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})\simeq c^{+}\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket\oplus c^{-}\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$
where by $c^{+}$ and $c^{-}$, we really mean their free-stabilizations $S(c^{+})$ and $S(c^{-})$ from Subsection 2.6. In fact, these are canonical since $c^{+}$ is the top $\mathbb{Z}$-grading generator and $c^{-}=A_{\gamma}(c^{+})$ where $\gamma$ is the homology class of the vertical circle in Figure 9.3 and $A_{\gamma}$ is the $H_{1}$-action from [OS04c, Subsection 4.2.5] (see [Zem15, Subsection 5.3]). Since $\widetilde{\boldsymbol{\alpha}}$ and $\widetilde{\boldsymbol{\beta}_{0}}$ are obtained by handsliding $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}_{0}$, respectively, there are canonical isomorphisms from $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})$ to $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\widetilde{\boldsymbol{\beta}_{0}})$ and $\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\widetilde{\boldsymbol{\beta}_{0}})$, and we use $c^{+},c^{-}$ to denote the image of these generators under the canonical isomorphisms.
Similarly, identify
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2})\simeq d_{0}\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket\oplus d_{1}\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket.$
This is also canonical since $d_{0}$ and $d_{1}$ are the top grading generators in their respective Alexander $\mathbb{Z}/2$ grading or $\{v,z\}$-${\rm Spin}^{c}$-structure. We similarly write $d_{0}$ and $d_{1}$ for the image of these elements under the canonical isomorphisms from $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2})$ to groups canonically isomorphic to it. We also identify
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E})\simeq e_{0}x\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket\oplus e_{1}x\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket\simeq\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\boldsymbol{\beta}_{\infty}^{E}).$
Also, let $b$ be the top grading generator of $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\simeq\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$.
We discuss $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})$ and related groups in the next subsubsection.
9.2.2. The maps $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$ and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$
Although the homology groups $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})$ and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\widetilde{\boldsymbol{\beta}_{0}})$; and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2})$ and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\widetilde{\boldsymbol{\beta}_{2}})$ are canonically isomorphic via $\mu_{2}(-,\Theta_{0}^{+})$ and $\mu_{2}(-,\Theta_{2}^{+})$, respectively, it is not a priori clear whether $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})$ and $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$ are. Indeed, the isomorphism $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$ is given by the following dashed map on homology.
${{\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})}}$${{\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2})}}$${{\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\widetilde{\boldsymbol{\beta}_{0}})}}$${{\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\widetilde{\boldsymbol{\beta}_{2}})}}$$\scriptstyle{\mu_{2}(-,\theta)}$$\scriptstyle{\mu_{2}(-,\Theta_{0}^{+})}$$\scriptstyle{\mu_{3}(-,\theta,\Theta_{2}^{+})+\mu_{3}(-,\Theta_{0}^{+},\widetilde{\theta})}$$\scriptstyle{\mu_{2}(-,\Theta_{2}^{+})}$$\scriptstyle{\mu_{2}(-,\widetilde{\theta})}$
In other words, this is the case where we know that the outer square commutes up to homotopy, but we really care which homotopy we chose.
Fortunately, the chain complex has a nice splitting into two sub chain complexes:
$\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})=\left(\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})\oplus\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2};\mathfrak{s}_{0})\right)\oplus\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2};\mathfrak{s}_{1}),$
where $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2};\mathfrak{s}_{0})$ is generated, as an $\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$-module, by the intersection points whose $\{v,z\}$-${\rm Spin}^{c}$-structure is the same as $d_{0}$ and $\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2};\mathfrak{s}_{1})$ is generated by the rest. Let us call this the ${\rm Spin}^{c}$-splitting for now [^47] This is basically the $\mathbb{Z}/2$-splitting, but $U_{2}^{1/2}$ has grading 00 in this case..
Since we only consider $\boldsymbol{\alpha},\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{2},\widetilde{\boldsymbol{\alpha}},\widetilde{\boldsymbol{\beta}_{0}},\widetilde{\boldsymbol{\beta}_{2}}$, there is also a $\mathbb{Z}$-grading. Shift the $\mathbb{Z}$-grading of $\boldsymbol{\beta}_{0}$ or $\boldsymbol{\beta}_{2}$ appropriately such that $\underline{\boldsymbol{\beta}_{02}}$ is a twisted complex in the graded sense. Then, there is a canonical splitting, i.e. an $\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$-linear isomorphism
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\simeq x\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket\oplus y\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$
where $x$ and $y$ are homogeneous with respect to both the $\mathbb{Z}$-grading and the ${\rm Spin}^{c}$-splitting, and $x$ and $y$ lie in different ${\rm Spin}^{c}$-summands. In particular, with respect to the splitting (as modules)
$\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})=\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{0})\oplus\boldsymbol{CF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{2})$
we can assume that $x=(c^{-},\ast)$ for some $\ast$, and $y=(0,d_{1})$. By abuse of notation, we denote $x$ and $y$ as $c^{-}$ and $d_{1}$, respectively.
Use the same notation for $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$. This is now justified, since the isomorphism $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$ preserves both the $\mathbb{Z}$-grading and the ${\rm Spin}^{c}$-splitting, and is $\mathbb{F}\llbracket U_{2}^{1/2}\rrbracket$-linear. Hence, the map sends $c^{-}$ to $c^{-}$ and $d_{1}$ to $d_{1}$. Similarly, the isomorphism $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\underline{\widetilde{\boldsymbol{\beta}_{02}}})$ sends $c^{-}$ to $c^{-}$ and $d_{1}$ to $d_{1}$.
Remark 9.7.
The chain complex $\boldsymbol{CF}^{-}(\alpha,\underline{\beta_{02}})$ is simpler if the curve $\alpha$ in Figure 9.2 had slope $1$ instead of 00. However, it would be harder to compute the maps considered in this subsection.
9.2.3. The maps $\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}})$ and $\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\boldsymbol{\beta}_{\infty}^{E})$
We pinch the neck and use Proposition 9.5, which gives the following:
  • $\mu_{2}(-,\zeta):\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\underline{\boldsymbol{\beta}_{02}}):b\mapsto c^{-}+uU_{2}^{1/2}d_{1}$;
  • $\mu_{2}(-,e_{0}\xi):\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\underline{\widetilde{\boldsymbol{\beta}_{02}}})\to\boldsymbol{HF}^{-}(\widetilde{\boldsymbol{\alpha}},\boldsymbol{\beta}_{\infty}^{E}):c^{-}\mapsto ue_{1}x,\ d_{1}\mapsto e_{0}x$.
9.2.4. The split band map
Putting all these together, we get that the composite is
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E}):b\mapsto ue_{1}x+uU_{2}^{1/2}e_{0}x.$
and hence the split band map is
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E}):b\mapsto e_{1}x+U_{2}^{1/2}e_{0}x.$
Remark 9.8.
We could have worked with the maps discussed in Remark 5.6 that are morally homotopy inverses of each other. In this case, the composite is indeed just
$\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{b})\to\boldsymbol{HF}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta}_{\infty}^{E}):b\mapsto e_{1}x+U_{2}^{1/2}e_{0}x.$
9.3. Merge bands
One way to compute the merge band map is to argue as in Subsection 9.2. There is a simpler way: the merge band map for planar links can be computed using the same triple Heegaard diagram as the split band map, by composing with different algebraic maps as in Definition 3.21.
Figure 9.4. Multi-Heegaard diagram for both the merge and split band maps for planar links
Indeed, consider Figure 9.4. Let $R=\mathbb{F}\llbracket U_{1}^{1/2},U_{2}^{1/2}\rrbracket$ and $S=\mathbb{F}\llbracket U_{1},U_{2}^{1/2}\rrbracket$. Assign the weight $U_{i}^{1/2}$ to both $w_{i}$ and $z_{i}$, and let $\Theta_{B}=ay+bx\in\boldsymbol{CF}^{-}(\boldsymbol{\beta},\boldsymbol{\gamma})$ over any ring. We have computed the split band map, which is
$\boldsymbol{HF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\to\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\xrightarrow{\mu_{2}(-,\Theta_{B})}\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}).$
Identifying
$\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\simeq\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})\simeq R/(U_{1}-U_{2}),\ \boldsymbol{HF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\simeq S/(U_{1}-U_{2}),$
the split band map is $S$-linear, and is given by $1\mapsto U_{1}^{1/2}+U_{2}^{1/2}$. Hence, as the map
$\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\xrightarrow{\mu_{2}(-,\Theta_{B})}\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})$
is $R$-linear, it is determined by $1\mapsto U_{1}^{1/2}+U_{2}^{1/2}$. Now we can compute the merge band map: it is
$\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\beta})\xrightarrow{\mu_{2}(-,\Theta_{B})}\boldsymbol{HF}_{R}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})\to U_{1}^{1/2}\boldsymbol{HF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma})\xrightarrow{\cdot U_{1}^{-1/2}}\boldsymbol{HF}_{S}^{-}(\boldsymbol{\alpha},\boldsymbol{\gamma}),$
which is the $S$-linear map given by
$1\mapsto 1,\ U_{1}^{1/2}\mapsto U_{2}^{1/2},$
which is what we wanted.